### v1

# Generalized Recurrence Criteria for Classes of Open Quantum Walks

Abstract
In this paper, we study the recurrence of open quantum walks (OQWs) induced by finite-dimensional coins $(L,B,R)$. The focus is on homogeneous OQWs with a set of vertices $\mathbb{Z}$, the set of integers. We present three distinct recurrence criteria, each adapted to different types of coins. The first criterion was developed for a class of Lazy OQWs in any finite dimension, where the presented criterion is associated with an auxiliary map and its only invariant state, resulting in the first recurrence criterion for Lazy OQWs. The second one is restricted to Lazy OQWs of dimension 2, where we provide a complete characterization of the recurrence for this lower dimension. Finally, we present a general criterion for finite-dimensional coins in the non-lazy case $(B=0)$, which generalizes many of the previously known results. This new criterion holds for irreducible and reducible OQWs through a decomposition of the Hilbert space where our quantum states act.
keywords:
Open Quantum Walks, Recurrence, Lazy walks, Coin, General Criterion
1 Introduction
Open Quantum Walks (OQWs) represent a natural extension of classical Markov chains through open quantum systems. Originally proposed by [2], OQWs are models that capture the behavior of open quantum systems, in which the measurement of the particle position after each time step leads to a classical Markov chain when it is joined with the space of particle states.
In this context, we consider homogeneous OQWs defined on the set of integers, that is, the graph of this type of random walk is the set of vertices $V=\mathbb{Z}$, where the linear operators $R$ and $L$ play the roles of transitions between the vertices of the quantum walk, which are constant and each jump only occurs from one vertex to its nearest neighbors. Homogeneous OQWs in the Lattice are explored and detailed in [6].
Recurrence is an important statistic of any random walk, emerging as a central question of interest in the quantum context, and can be directly connected with the classical recurrence criteria for random walks in $\mathbb{Z}$. For example, for homogeneous classical random walks on $\mathbb{Z}$, it is well known that a walk is recurrent if and only if the transition parameters $r$ (rate of jumping to right) and $l$ (rate of jumping to left) are equal (or $r=1/2=l$ in the case $b=0$) [11]. This classical random walk has its graph illustrated in Figure 1.
[…]
Figure 1: Classical Lazy Random Walk on $\mathbb{Z}$.
With this description, it is natural to seek criteria to determine under what conditions a homogeneous OQW is recurrent given $R$ and $L$. This problem has been addressed in the literature with the development of specific criteria, which are essential to understanding the dynamics and long-term behavior of these systems.
A Criterion for PQ Matrices can be found in [10]. Also, in that work the authors analyze the positive recurrence of this type of walk. From the analysis of the spectrum of $L^{*}L$ and $R^{*}R,$ it was showed in [7] that $R$ and $L$ having only singular values equal to $1/\sqrt{2}$ is a sufficient condition to guarantee that those OQWs are recurrent. More recently, it was obtained general criteria for recurrence of non-Lazy OQWs in dimension 2 and for irreducible finite-dimensional OQWs [9].
In this article, we present a general criterion for the recurrence of homogeneous OQWs, based on the transition operators $R$ and $L$ when acting in finite-dimensional vector spaces, that is, we allow the OQW to be reducible in any finite dimension. Furthermore, we present the first recurrence criterion for lazy OQWs, using results obtained from recent studies on Central Limit Theorems for this kind of OQWs.
A first Central Limit Theorem for OQWs was presented by [1]. Here we will use more recent results. These works presented a Central Limit Theorem for Lazy OQWs [8] and one for any homogeneous OQW in Lattice when the degrees of freedom are finite [5].
This work is structured as follows. In Section 2 we define the open quantum walks induced by a coin and their transition probabilities. Section 3 is devoted to presenting a recurrence criterion for a class of Lazy OQWs, and one restricted to Lazy OQWs of dimension 2. In Section 4 we focus on the case where the coin is non-lazy and we present a general recurrence criterion for finite-dimensional coins. Section 5 illustrates the results with comprehensive examples.
2 General Settings
Let us consider a set of vertices $V$, which might be finite or countable infinite. Let $\mathcal{K}=\mathbb{C}^{V}$ be the state space of a quantum system with as many degrees of freedom as the number of vertices. In the case where $V$ is infinite countable, we assume $\mathcal{K}$ to be any separable Hilbert space with an orthonormal basis indexed by $V$, denote by $\left\{\lvert i\rangle\right\}_{i\in V}$. To emphasize the position of the particle of our quantum walk, we consider a quantum system described by the separable complex Hilbert space
$\mathcal{H}=\bigoplus_{i\in V}\mathfrak{h}_{i},$
where the $\mathfrak{h}_{i}$ are separable Hilbert spaces. The label $i\in V$ represents the position of the particle and, when the particle is located at the vertex $i\in V$ , its internal state is encoded in the space $\mathfrak{h}_{i},$ as it will be detailed below.
We consider the set of density operators over $\mathcal{H},$
$\mathcal{D}(\mathcal{H})=\{\rho\in\mathcal{H}:\rho\geq 0,\mathrm{Tr}(\rho)=1\},$
and introduce
$\mathcal{S}(\mathcal{H}\otimes\mathcal{K})=\left\{\tau\in\mathcal{I}(\mathcal{H}\otimes\mathcal{K}):\textstyle\tau=\displaystyle\sum_{j\in\mathbb{Z}}\rho(j)\otimes\lvert j\rangle\langle j\rvert\right\},$
the set of density operators acting on $\mathcal{H}\otimes\mathcal{K}.$ Fixed a set of vertices $V,$ an open quantum random walk (OQW) on $V$ is a completely positive map
$\Phi:\mathcal{H}\otimes\mathcal{K}\mapsto\mathcal{H}\otimes\mathcal{K},$
given by
$\Phi(\tau)=\sum_{i\in V}\Bigg{(}\sum_{j\in V}B_{i}^{j}\rho_{j}B_{i}^{j^{*}}\Bigg{)}\otimes|i\rangle\langle i|,\;\;\;\tau\in\mathcal{S}(\mathcal{H}\otimes\mathcal{K}),$
where $B_{i}^{j}$ are bounded operators on $\mathcal{H}$ corresponding to the transition effect from site $j$ to $i$, and
$\sum_{i\in V}B_{i}^{j*}B_{i}^{j}=I,\;\;\;j\in V,$
which guarantees that $\mathrm{Tr}(\Phi(\tau))=\mathrm{Tr}(\tau)$, for every $\tau\in\mathcal{D}(\mathcal{H})$, meaning that $\Phi$ is trace-preserving, thus $\Phi$ preserves the set $\mathcal{S}(\mathcal{H}\otimes\mathcal{K}),$ i.e., $\Phi(\tau)\in\mathcal{S}(\mathcal{H}\otimes\mathcal{K})\;\forall\tau\in\mathcal{S}(\mathcal{H}\otimes\mathcal{K}).$
In this work, we focus on the OQW model presented in [2], where the set of vertices is $V=\mathbb{Z}.$ For simplicity, we will say integer line, or simply the line, to refer to the set $\mathbb{Z}$ of integer numbers. An OQW on $\mathbb{Z}$ concerns the evolution of density operators of the form
$\tau=\sum_{j\in\mathbb{Z}}\rho(j)\otimes\lvert j\rangle\langle j\rvert\;\in\;\mathcal{H}\otimes\mathbb{C}^{\mathbb{Z}},$
where the $V=\mathbb{Z}$ represents the position of the walker and $\mathfrak{h}_{i}$ is the Hilbert space describing the internal degrees of freedom for site $\lvert i\rangle$, which we call the coin space and will be denoted by $\mathfrak{h}_{i}=\mathbb{C}^{k}$. Finally, we consider homogeneous lazy open quantum walks: at each step either the particle on site $\lvert j\rangle$ moves exactly one position to some of its two nearest-neighbours, in the left ($\lvert j-1\rangle$) or in the right ($\lvert j+1\rangle$), or it remains at same site ($\lvert j\rangle$), and the coin is constant, i.e., the same transition rule is applied at each site. Therefore, the linear operator $\Phi$ can be written by
$\Phi(\tau)=\sum_{i\in\mathbb{Z}}\sum_{j=i-1}^{i+1}M_{i}^{j}\tau(M_{i}^{j})^{*},$
where
$M_{i}^{i-1}=L\otimes\lvert i-1\rangle\langle i\rvert,\quad M_{i}^{i}=B\otimes\lvert i\rangle\langle i\rvert,\quad M_{i}^{i+1}=R\otimes\lvert i+1\rangle\langle i\rvert.$
The linear operators $L,B,R\in\mathbb{M}_{k}({\mathbb{C}})$ are matrices such that $L^{*}L+B^{*}B+R^{*}R=I_{k},$ and simplified by
$\Phi(\tau)=\sum_{i\in\mathbb{Z}}\Big{(}L\rho(i+1)L^{*}+B\rho(i)B^{*}+R\rho(i-1)R^{*}\Big{)}\otimes|i\rangle\langle i|,\;\;\;\tau\in\mathcal{S}(\mathcal{H}\otimes\mathcal{K}).$ (1)
We say that the triple of matrices $(L,B,R)$ is a coin if eq. (1) holds. Therefore, we can assume that $\mathfrak{h}_{i}=\mathfrak{h}\;\forall i\in\mathbb{Z},$ and if $\mathfrak{h}=\mathbb{C}^{d}$ for some $d$, the dimension of the coin is said to be $d$. In this article, we consider only the case $d<\infty.$
If $B\neq 0,$ the quantum walk exhibits a lazy behavior, meaning that in the context of quantum walks there is a non-zero probability for the walker to stay in the same vertex after a step. We refer to the OQW above as an OQW induced by the coin $(L,B,R)$ in which the walk is inherently lazy. This lazy behavior introduces distinct properties compared to its non-lazy counterpart. Notably, the presence of the matrix $B$ influences the walk’s propensity to remain in its current vertex, thereby affecting the overall dynamics. If $B=0,$ then $\Phi$ will be called an OQW induced by a coin $(L,R)$ or an non-lazy OQW induced by a coin $(L,R).$
The dynamic described above is represented by the graph in Figure 2.
[…]
Figure 2: Lazy Quantum Walk on $\mathbb{Z}$.
The OQW induced by a coin $(L,B,R)$ is the homogeneous, nearest-neighbor OQW on the integer line with the left and right transitions given by $L$ and $R$, respectively, and $B$ represents the rate of no jump at some instant. In this sense, the quantum trajectory of this kind of OQW on the line, induced by a coin $(L,B,R)$ and starting from a state $\tau$ of the form $\tau=\sum_{i\in\mathbb{Z}}\rho(i)\otimes|i\rangle\langle i|,$ is any path generated by the Markov chain $(X_{n},\tau_{n})_{n\geq 0}$, where $X_{n}$ denotes the position of the particle at time $n$ and $\tau_{n}$ its internal degree. The transition probabilities are given by
$\begin{split}\mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left(i-1,\frac{L\sigma L^{*}}{\mathrm{Tr}(L\sigma L^{*})}\right)\;\Bigg{|}\;(x_{n},\rho_{n})=(i,\sigma)\,\right)&=\mathrm{Tr}(L\sigma L^{*}),\\ \mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left(i,\frac{B\sigma B^{*}}{\mathrm{Tr}(B\sigma B^{*})}\right)\;\Bigg{|}\;(x_{n},\rho_{n})=(i,\sigma)\,\right)&=\mathrm{Tr}(B\sigma B^{*}),\\ \mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left(i+1,\frac{R\sigma R^{*}}{\mathrm{Tr}(R\sigma R^{*})}\right)\;\Bigg{|}\;(x_{n},\rho_{n})=(i,\sigma)\,\right)&=\mathrm{Tr}(R\sigma R^{*}).\end{split}$
for every $i\in\mathbb{Z}$, $\sigma\in\mathcal{S}(\mathfrak{h}_{i})$, and initial law
$\mathbb{P}\left(\,(X_{0},\rho_{0})=\left(i,\frac{\rho(i)}{\mathrm{Tr}\,\rho(i)}\right)\;\right)=\mathrm{Tr}(\rho(i)),$
and all other transition probabilities are null.
3 Recurrence criteria for Lazy Open Quantum Walks
3.1 Faithfulness of Density Operators
Density operators play a fundamental role in open quantum random walks, providing a statistical description of this type of quantum system. Faithful and non-faithful density operators are two categories that arise in the study of quantum states, being characterized by their positive definiteness and positive-semidefiniteness, respectively.
Faithful density operators are defined as being positive definite. This means that the eigenvalues of a faithful density operator are strictly positive and it is invertible. In essence, faithfulness ensures a faithful representation of the quantum system, capturing the full richness of its physical properties.
On the other hand, non-faithful density operators are positive-semidefinite, thus they have $\lambda=0$ as one of their eigenvalues. Thereby, that there may be ambiguity in the quantum state description, as multiple states can correspond to the same non-faithful density operator. The positivity-semidefiniteness criterion relaxes the strictness of positive definiteness, permitting a broader range of quantum state representations.
An important subset of the set of non-faithful density operators is composed of the pure density operators. A pure density is a density $\rho$ that is a projection of rank 1, i.e., there exists some vector $\lvert u\rangle\in\mathbb{C}^{k}$ such that $\rho=\lvert u\rangle\langle u\rvert$.
Let us consider an OQW starting the walk on vertex $\lvert i\rangle$ with initial density operator $\rho.$ For evolution on time $n=0,1,2,\ldots,$ we define
$p_{n}(j)=\mathbb{P}(\mbox{the quantum particle, at time }n,\mbox{ is in site}\;\lvert j\rangle)=\mathrm{Tr}(\rho_{n}(j)),$
where
$\Phi^{n}(\rho\otimes\lvert i\rangle\langle i\rvert)=\sum_{s\in V}\rho_{n}(s)\otimes\lvert s\rangle\langle s\rvert.$
With this notation,
$p_{ji;\rho}(n)=p_{n}(\rho\otimes|i\rangle\to|j\rangle)=\mathrm{\mathrm{Tr}}(\rho_{n}(j)\otimes\lvert j\rangle\langle j\rvert)=\mathrm{\mathrm{Tr}}\left(\Phi^{n}(\rho\otimes\lvert i\rangle\langle i\rvert)(I\otimes\lvert j\rangle\langle j\rvert)\right)$
denote the probability of being at site $j$ at step $n,$ given that we started at site $i,$ with initial density $\rho$ concentrated at $i.$ Therefore, the dynamics starts with a density operator $\rho$ concentrated at some vertex $\lvert i\rangle,$ takes the evolution up to time ($n$) through the linear operator $\Phi^{n},$ producing a new density operator
$\rho_{n}=\sum_{s}\rho_{n}(s)\otimes\lvert s\rangle\langle s\rvert=\Phi^{n}(\rho\otimes\lvert i\rangle\langle i\rvert),\quad\mathrm{Tr}\left(\sum_{s}\rho_{n}(s)\right)=1,$
and then we project $\rho_{n}$ onto the subspace generated by vertex $\lvert j\rangle.$ Note that
$\Phi^{n}(\rho\otimes\lvert i\rangle\langle i\rvert)(I\otimes\lvert j\rangle\langle j\rvert)=\rho_{n}(j)\otimes\lvert j\rangle\langle j\rvert,$
which represents the data concentrated at vertex $\lvert j\rangle$ at time $n.$
Fixed some $i\in V,$ we denote the set of density operators acting on $\mathfrak{h}_{i}$ by
$\mathcal{S}(\mathfrak{h}_{i})=\left\{\rho:\mathfrak{h}_{i}\rightarrow\mathfrak{h}_{i}:\;\mathrm{Tr}(\rho)=1,\;\rho=\rho^{*};\;\rho\geq 0\right\}.$
Consider an OQW on some set of vertices $V$, let $\lvert i\rangle\in V$ and $\rho\in\mathcal{S}(\mathfrak{h}_{i})$. We say that a vertex $\lvert i\rangle$ is
  1. 1.
    $\rho$-recurrent if
    $\sum_{n=0}^{\infty}p_{ii;\rho}(n)=\infty.$
    Otherwise, $i$ is said to be $\rho$-transient;
  2. 2.
    recurrent, if $i$ is $\rho-$recurrent for all $\rho\in\mathcal{S}(\mathfrak{h}_{i});$
  3. 3.
    transient, if $i$ is $\rho-$transient for all $\rho\in\rho\in\mathcal{S}(\mathfrak{h}_{i}).$
Concerning all vertices, an OQW is called: recurrent if every vertex is recurrent; transient if every vertex is transient.
The following Proposition describes $\rho$-recurrence in a concentrated vertex $\lvert i\rangle\in V$ when assumed that $\lvert i\rangle$ is recurrent for some density and $\mathfrak{h}_{i}$ has a finite internal degree of freedom. An immediate consequence will also be given for the case where $\lvert i\rangle$ is $\sigma$-transient for some non-faithful density.
Theorem 3.1.
Consider an OQW on some set of vertices $V,$ let $i\in V$ and assume that $\mathfrak{h}_{i}$ is finite-dimensional. If $\lvert i\rangle$ is $\tau$-recurrent for some $\tau\in\mathcal{S}(\mathfrak{h}_{i}),$ then
  1. 1.
    vertex $\lvert i\rangle$ is $\rho$-recurrent for all faithful $\rho\in\mathcal{S}(\mathfrak{h}_{i})$;
  2. 2.
    there exists a non-faithful density operator $\tilde{\rho}$ such that $\lvert i\rangle$ is $\tilde{\rho}$-recurrent;
  3. 3.
    there exists a pure density operator $\lvert w\rangle\langle w\rvert$ such that the OQW is $\lvert w\rangle\langle w\rvert$-recurrent, where $\lvert w\rangle$ is an unitary vector of $\mathfrak{h}_{i}.$
Proof.
1. Let $\rho\in\mathcal{S}(\mathfrak{h}_{i})$ be faithful. By the compactness of $\mathcal{S}(\mathfrak{h}_{i}),$ there exists $c>0$ such that $\rho>c\tau,$ thus
$\begin{split}\sum_{n=0}^{\infty}p_{ii;\rho}(n)=&\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\rho\otimes\lvert i\rangle\langle i\rvert\right)\right(I\otimes\lvert i\rangle\langle i\rvert)]\\ >&c\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\tau\otimes\lvert i\rangle\langle i\rvert\right)(I\otimes\lvert i\rangle\langle i\rvert)\right]\\ =&c\sum_{n=0}^{\infty}p_{ii;\tau}(n)\\ =&\infty.\end{split}$
2. Let $\rho\in\mathcal{S}(\mathfrak{h}_{i})$ be non-faithful. By the Spectral Theorem, $\rho$ can be written as
$\rho=\sum_{x=1}^{n}\lambda_{x}\lvert x\rangle\langle x\rvert,$ (2)
where the vectors $\lvert x\rangle$ are the eigenvectors of $\rho$ with corresponding eigenvalues $\lambda_{x}$. Since $\rho$ is non-faithful, there is at least one null eigenvalue and the remainder of eigenvalues are non-negative summing $1$ Thus, (2) can be rewritten as
$\rho=\sum_{x\in S}\lambda_{x}\lvert x\rangle\langle x\rvert,\quad S\varsubsetneq\{1,\ldots,n\}.$
Take a sequence of positive numbers $(\alpha_{r})_{r\in R},$ where $R:=\{1,\ldots,n\}/S\neq\emptyset,$ whose sum is $1$
Defining
$\rho_{X}=\sum_{x\in S}\frac{\lambda_{x}}{2}\lvert x\rangle\langle x\rvert+\sum_{x\in R}\frac{\alpha_{x}}{2}\lvert x\rangle\langle x\rvert=\sum_{x=1}^{n}\frac{\tilde{\alpha}_{x}}{2}\lvert x\rangle\langle x\rvert,\quad\tilde{\alpha}_{x}=\begin{cases}\lambda_{x},&\mbox{if }x\in S\\ \alpha_{x},&\mbox{if }x\in R\end{cases},$
we get by the first item that $\lvert i\rangle$ is $\rho_{X}$-recurrent, since $\rho_{X}$ is faithful.
Now, define
$\rho_{Y}=\sum_{x\in R}\alpha_{x}\lvert x\rangle\langle x\rvert,$
which is a non-faithful density operator. Since $2\rho_{X}=\rho+\rho_{Y},$ we obtain by linearity the equality
$\sum_{n=0}^{\infty}p_{ii;\rho}(n)+\sum_{n=0}^{\infty}p_{ii;\rho_{Y}}(n)=\sum_{n=0}^{\infty}\left(p_{ii;\rho}(n)+p_{ii;\rho_{Y}}(n)\right)=2\sum_{n=0}^{\infty}p_{ii;\rho_{X}}(n).$
The sum on the right hand diverges once $\lvert i\rangle$ is $\rho_{X}$-recurrent, implying that at least one of the sums on the left-hand side diverges. Therefore, $\lvert i\rangle$ is $\rho$-recurrent or $\rho_{Y}$-recurrent, both being non-faithful.
3. Let $\tilde{\rho}$ be one non-faithful density given by item (2), then we have
$\tilde{\rho}=\sum_{j}\lambda_{j}\lvert e_{j}\rangle\langle e_{j}\rvert,$
where each $\lambda_{j}\geq 0$ is an eigenvalue of $\tilde{\rho}$ with corresponding eigenvector $\lvert e_{j}\rangle.$ If $\tilde{\rho}$ has only one eigenvector, then we are done. If $\tilde{\rho}$ has at least two eigenvectors, then
$\infty=\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\rho\otimes\lvert i\rangle\langle i\rvert\right)\right(I\otimes\lvert i\rangle\langle i\rvert)]=\sum_{j,\lambda_{j}\neq 0}\left(\lambda_{j}\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\lvert e_{j}\rangle\langle e_{j}\rvert\otimes\lvert i\rangle\langle i\rvert\right)\right(I\otimes\lvert i\rangle\langle i\rvert)]\right),$
where at least one of the members of the sum must be infinity, and some $\lambda_{j}$ is positive. This completes the proof.
$\square$
By a contra-position argument, we obtain the following direct consequence of Theorem 3.1.
Corollary 3.2.
Let $\Phi$ be an OQW on some set of vertices $V$ and $i\in V,$ and $\dim(\mathfrak{h}_{i})<\infty.$ If $\lvert i\rangle$ is $\tau$-transient for some faithful $\tau\in\mathcal{S}(\mathfrak{h}_{i}),$ then $\lvert i\rangle$ is transient.
Another direct consequence of Theorem 3.1 is related to the homogeneous case.
Corollary 3.3.
Let $\Phi$ be a homogeneous OQW on $\mathbb{Z}$ with $\dim(\mathfrak{h})=d<\infty.$ If $\Phi$ is $\tau$-recurrent for some $\tau\in\mathcal{S}(\mathfrak{h}),$ then
  1. 1.
    $\Phi$ is $\rho$-recurrent for all faithful $\rho\in\mathcal{S}(\mathfrak{h})$;
  2. 2.
    there exists a non-faithful density operator $\tilde{\rho}$ such that $\Phi$ is $\tilde{\rho}$-recurrent;
  3. 3.
    there exists a pure density operator $\lvert w\rangle\langle w\rvert$ such that $\Phi$ is $\lvert w\rangle\langle w\rvert$-recurrent, where $\lvert w\rangle$ is an unitary vector of $\mathbb{C}^{d}.$
3.2 Enclosures and Invariant States
Firstly, a subspace $\mathcal{Y}$ of $\mathfrak{h}$ is called an enclosure for a quantum channel $\mathcal{L}$ if for every positive $\sigma\in L^{1}(\mathfrak{h}),$ we have
$\mathrm{supp}(\sigma)\subset\mathcal{Y}\Rightarrow\mathrm{supp}\left(\mathcal{L}(\sigma)\right)\subset\mathcal{Y}.$
If an enclosure contains no other non-trivial enclosures, then it is called a minimal enclosure.
A minimal enclosure with finite dimension is the support of a unique invariant state of the channel, that is, a density operator $\sigma$ such that $\mathcal{L}(\sigma)=\sigma$. For more on enclosures of OQWs, we refer the reader to [5, 6, 3].
Following this property and using the homogeneity of OQWs induced by a coin, we split $\mathcal{S}(\mathfrak{h})$ into two subsets $\mathcal{S}(\mathcal{Y})$ and $\mathcal{S}(\mathcal{X})$ in such a way that the walk is transient for $\mathcal{S}(\mathcal{Y})$ and recurrent for $\mathcal{S}(\mathcal{X}),$ where $\mathcal{Y}$ is a enclosure of $\mathcal{L}:$
Proposition 3.4.
Consider an OQW induced by a coin $(L,B,R)$ of dimension $d<\infty$. We are in one (and only one) of the following situations:
  • $(L,B,R)$ is recurrent;
  • $(L,B,R)$ is transient;
  • $(L,B,R)$ is recurrent with respect to all densities except for a convex set $\mathcal{S}(\mathcal{Y}).$ In this case, $\mathcal{Y}$ is an enclosure of $\mathcal{L},$
    $\mathfrak{h}=\mathcal{Y}\oplus\mathcal{X}$
    for some subspace $\mathcal{X}$ of $\mathfrak{h},$ with $\mathcal{Y},\mathcal{X}\neq\{0\}.$
Proof.
Let us suppose that $(L,B,R)$ is neither recurrent nor transient. By Proposition 3.1, there exist unitary vectors $\lvert u\rangle,\lvert v\rangle\in\mathfrak{h}$ such that the walk is $\lvert u\rangle\langle u\rvert$-recurrent and $\lvert v\rangle\langle v\rvert$-transient. Let $\mathcal{Y}$ be the subset of $\mathfrak{h}$ such that $\mathcal{S}(\mathcal{Y})$ is the set of densities of $\mathcal{B}(\mathfrak{h})$ in which $(L,B,R)$ is transient, thus we pick $\rho\in\mathcal{S}(\mathcal{Y})$ and we can affirm that $\rho_{n}\in\mathcal{S}(\mathcal{Y}),$ where $(X_{n},\rho_{n})_{n}$ are the quantum trajectories of the OQW. Indeed, if we had $\rho_{n}\in\mathcal{S}(\mathfrak{h})\setminus\mathcal{S}(\mathcal{Y})$ for some $n,$ then we could refresh the walk and the vertex $\lvert X_{n}\rangle$ would be $\rho_{n}$-recurrent, contradicting the hypothesis of homogeneity, thus we must have that $\mathcal{Y}$ is an enclosure of $\mathcal{L}.$ Therefore, take $\mathcal{X}=\mathcal{Y}^{\perp}$ and the trichotomy holds.
$\square$
In our subsequent discussion, we consider a coin $(L,B,R)$ and its auxiliary map $\mathcal{L}:\mathcal{B}(\mathfrak{h})\rightarrow\mathcal{B}(\mathfrak{h})$ given by
$\mathcal{L}(\rho)=L\rho L^{*}+B\rho B^{*}+R\rho R^{*},\quad\rho\in\mathcal{B}(\mathfrak{h}).$
Note that $\mathcal{L}$ is a quantum channel, thus it has at least one invariant state, that is, there exists some density operator $\rho_{\infty}\in\mathcal{B}(\mathfrak{h})$ satisfying $\mathcal{L}(\rho_{\infty})=\rho_{\infty}.$ When the invariant state of $\mathcal{L}$ is unique, we will say that the map $\mathcal{L}$ is ergodic.
We will use the following result obtained straightforwardly from [8].
Theorem 3.5.
Consider a Lazy OQW induced by a coin $(L,B,R)$ of dimension $d<\infty.$ If the auxiliary map $\mathcal{L}$ admits only one invariant state $\rho_{\infty},$ then
$\frac{X_{n}}{n}\stackrel{{\scriptstyle n\rightarrow\infty}}{{\longrightarrow}}m,\qquad\text{where }\qquad m=\mathrm{Tr}\left(L\rho_{\infty}L^{*}\right)-\mathrm{Tr}\left(R\rho_{\infty}R^{*}\right),$
in probability.
Proof.
This result can be obtained by using a Doob decomposition as described in [8, Equation 27] and [1, Theorem 5.2] for a Martingale $M_{n}$ satisfying
$X_{n}-n.m=X_{0}+M_{n}+\beta_{n},$
where $M_{n}/n\rightarrow 0$ and $\beta_{n}\rightarrow 0$ when $n\rightarrow 0.$
$\square$
Theorem 3.6 (Chung-Fuchs Theorem for OQWs with ergodic auxiliary map).
Consider an OQW induced by a coin $(L,B,R)$ such that $\mathcal{Y}$ is an enclosure Suppose that the initial density operator is $\rho\in\mathcal{S}(\mathcal{Y})$ with quantum trajectories $(X_{n},\rho_{n})_{n\geq 0}$ and $X_{n}/n\rightarrow 0$ in probability, then $(L,B,R)$ is $\sigma$-recurrent for some $\sigma\in\mathcal{S}(\mathcal{Y}).$
Proof.
The proof is analogous to [9, Theorem 7] and we will proceed by remarking on how it can be generalized. Firstly, the maximum appearing in [9, Lemma 6] can be taken in $\mathcal{D}=\mathcal{S}(\mathcal{Y}).$ Indeed, if the initial state is $\rho\in\mathcal{S}(\mathcal{Y}),$ then $\rho_{n}\in\mathcal{S}(\mathcal{Y})$ for every $n,$ thereby the maximum taken in the proof of [9, Lemma 6] can be reduced to $\mathcal{S}(\mathcal{Y}).$ Therefore, for the same reason, the density operator considered in [9, Theorem 7] belongs to $\mathcal{S}(\mathcal{Y}).$
$\square$
Theorem 3.7 (Recurrence Criteria).
Let us consider an OQW on $\mathbb{Z}$ induced by a coin $(L,B,R)$ of dimension $d$ such that the auxiliary map $\mathcal{L}$ is ergodic. Then $\mathcal{L}$ has a unique invariant state $\rho_{\infty}$ and
  • $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)=\textmd{Tr}\left(R\rho_{\infty}R^{*}\right)\Rightarrow$ the walk is recurrent;
  • $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)\neq\textmd{Tr}\left(R\rho_{\infty}R^{*}\right)\Rightarrow$ the walk is transient.
Proof.
We suppose that the walk is neither recurrent nor transient. In this case, the walk is recurrent for some density and transient for all densities in some $\mathcal{S}(\mathcal{Y}),$ where $\mathcal{Y}$ is an enclosure (Proposition 3.4). Moreover, the unique invariant state $\rho_{\infty}$ of $\mathcal{L}$ must belong to $\mathcal{S}(\mathcal{Y}).$
If $\;\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)=\textmd{Tr}\left(R\rho_{\infty}R^{*}\right),$ then $X_{n}/n\rightarrow 0$ by Theorem 3.5, thus $(L,B,R)$ is $\rho$-recurrent for some $\rho\in\mathcal{S}(\mathcal{Y})$ by Theorem 3.6, which is a contradiction.
[…]
Therefore, if $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)=\textmd{Tr}\left(R\rho_{\infty}R^{*}\right),$ then the walk is $\rho$-recurrent for some density and thus it is recurrent. If $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)\neq\textmd{Tr}\left(R\rho_{\infty}R^{*}\right),$ then the walks is transient for every density.
$\square$
We remark that the operator $B$ does influence in the recurrence of $(L,B,R).$ Indeed, even it does not appear on the bullet points that describe the recurrence criterion, the invariant state $\rho_{\infty}$ may change when we pick distinct $B$’s. See example 5.3 below.
3.3 General Recurrence Criteria for OQWs induced by a coin $(L,B,R)$ of dimension 2
The following Proposition is a generalisation of [9, Proposition 12, item (3)]. We can prove it in an analogous demonstration given by [9], however, we would like to do it in terms of the non-ergodicity of the auxiliary map induced by this coin.
Proposition 3.8.
Consider an OQW induced by a coin $(L,B,R)$ of dimension 2 such that $L,B,R$ share two unitary linearly independent eigenvectors. Then for any $\lvert u\rangle$ orthogonal to one of these two eigenvectors, we have that $\lvert u\rangle$ is also a common eigenvector of $L,B,R.$
Proof.
Suppose $L,B,R$ share two unitary and linearly independent unitary eigenvectors $\lvert v\rangle,\lvert w\rangle,$ then $\tau_{1}=\lvert v\rangle\langle v\rvert$ and $\tau_{2}=\lvert w\rangle\langle w\rvert$ are distinct invariant states of the auxiliary map $\mathcal{L},$ thus $\mathcal{L}$ is not ergodic. This means that there are two minimal enclosures $\mathcal{Y}_{1},\mathcal{Y}_{2},$ and thus $\mathcal{Y}_{1}\perp\mathcal{Y}_{2}.$ The dimension of the coin gives that $\mathcal{Y}_{1}$ and $\mathcal{Y}_{2}$ are one-dimensional and therefore there exist $\lvert u_{1}\rangle,\lvert u_{2}\rangle\in\mathbb{C}^{2}$ such that $\mathcal{Y}_{1}=\mbox{span}(\lvert u_{1}\rangle)$ and $\mathcal{Y}_{2}=\mbox{span}(\lvert u_{2}\rangle).$ Thereby $\lvert u_{1}\rangle$ and $\lvert u_{2}\rangle$ are common orthonormal eigenvectors of $L,B,R.$
$\square$
Theorem 3.9.
Consider an OQW on $\mathbb{Z}$ induced by a coin $(L,B,R)$ of dimension 2.
(1) If $L,B,R$ have at most one common eigenvector, let $\rho_{\infty}$ be the unique invariant density of the auxiliary map. Then, we have
$(1.1)\quad\textmd{Tr}(L\rho_{\infty}L^{*})\neq\textmd{Tr}(R\rho_{\infty}R^{*})\Rightarrow(L,B,R)\mbox{ is transient},$
$(1.2)\quad\textmd{Tr}(L\rho_{\infty}L^{*})=\textmd{Tr}(R\rho_{\infty}R^{*})\Rightarrow(L,B,R)\mbox{ is recurrent}.$
(2) If $L,B,R$ have two linearly independent eigenvectors in common, $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ unitaries, then they satisfy $\lvert u_{1}\rangle\bot\lvert u_{2}\rangle,$ we can put
$L\lvert u_{j}\rangle=l_{j}\lvert u_{j}\rangle,\qquad B\lvert u_{j}\rangle=b_{j}\lvert u_{j}\rangle,\qquad R\lvert u_{j}\rangle=r_{j}\lvert u_{j}\rangle,\qquad j=1,2,$
and
$(2.1)\quad|l_{1}|=|r_{1}|\mbox{ and }|l_{2}|=|r_{2}|$$\Rightarrow$ $(L,B,R)\mbox{ is recurrent,}$,
$(2.2)\quad|l_{1}|\neq|r_{1}|\mbox{ and }|l_{2}|\neq|r_{2}|$$\Rightarrow$ $(L,B,R)\mbox{ is transient,}$,
$(2.3)\quad|l_{i}|=|r_{i}|\mbox{ and }|l_{j}|\neq|r_{j}|$$\Rightarrow$ $(L,B,R)\mbox{ is transient with respect to }\sigma_{j}=\lvert u_{j}\rangle\langle u_{j}\rvert\mbox{ and it is}$
recurrent with respect to all densities but $\mbox{recurrent with respect to all densities but }\sigma_{j}.$
Proof.
If the elements of the coin do not share at least two orthogonal eigenvectors, then the auxiliary map is ergodic, thus item (1) is a particular case of Theorem 3.7.
Now we suppose that $L,B,R$ share two linearly independent eigenvectors, then we can apply Proposition 3.8 to obtain $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ common orthonormal eigenvectors of $L,B,R.$ We assume that $|l_{1}|=|r_{1}|$ and $|l_{2}|=|r_{2}|,$ then for $\rho=\lvert u\rangle\langle u\rvert,$ where $\lvert u\rangle$ is some common eigenvector of $L,B,R,$ we have $p_{j,j+1;\rho}(n)=p_{j,j-1;\rho}(n)=w$ for all $j\in\mathbb{Z}$ and all $n\geq 0$ for some constant $w$ and therefore, in this case, we have a classical homogeneous walk which is recurrent. Since it is valid for any common eigenvector, the walk is recurrent by Proposition 3.1.
Analogously, if $|l_{1}|\neq|r_{1}|$ and $|l_{2}|\neq|r_{2}|,$ then for $\rho_{1}=\lvert u_{1}\rangle\langle u_{1}\rvert,\;\rho_{2}=\lvert u_{2}\rangle\langle u_{2}\rvert$ where $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ are orthonormal common eigenvectors of $L,B,R,$ we have $p_{j,j+1;\rho}(n)<p_{j,j-1;\rho}(n)$ (or $>$) for all $j\in\mathbb{Z}$ and all $n\geq 0$ and therefore in this case we have a classical homogeneous walk which is transient for both $\rho_{1},\rho_{2}$. Since the walk is transient for two distinct eigenvectors, the walk is transient by Proposition 3.1.
For the last case, $|l_{i}|=|r_{i}|$ and $|l_{j}|\neq|r_{j}|,$ $(L,B,R)$ is $\rho_{i}$-recurrent and $\rho_{j}$-transient, where $\rho_{i}=\lvert u_{i}\rangle\langle u_{i}\rvert$ and $\rho_{j}=\lvert u_{j}\rangle\langle u_{j}\rvert.$ This situation is a particular case of the last item in Proposition 3.4.
$\square$
4 A General Recurrence Criteria for Open Quantum Walks
Consider a coin $(L,R)$ of dimension $n<\infty.$ By [5], we can obtain a decomposition of $\mathfrak{h}$ in the following way:
$\mathfrak{h}=\mathfrak{R}\oplus\mathfrak{T},$
where
$\mathfrak{R}=\sup\{\mbox{supp}(\omega)\;|\;\omega\mbox{ is an invariant state for }\mathcal{L}\}.$
Since $\mathfrak{h}$ is finite-dimensional, any minimal enclosure is included in $\mathfrak{R},$ and is the support of a unique extremal invariant state. Moreover, we have a unique decomposition of $\mathfrak{R}$ of the form
$\mathfrak{R}=\bigoplus_{\alpha\in A}\mathfrak{X}_{\alpha},$
where $\{\mathfrak{X}_{\alpha}\}_{\alpha}\in A$ is a finite set of mutually orthogonal enclosures and every $\mathfrak{X}_{\alpha}$ is minimal in the set of enclosures satisfying
[…]
Further, we can write
$\mathfrak{R}=\bigoplus_{\alpha\in A}\mathfrak{X}_{\alpha}=\bigoplus_{\alpha\in A}\bigoplus_{\beta\in I_{\alpha}}\mathfrak{V}_{\alpha,\beta},$
where every $\mathfrak{X}_{\alpha}$ either is a minimal enclosure or assumes a decomposition (not unique) as the sum of mutually orthogonal isomorphic minimal enclosures $\{\mathfrak{V}_{\alpha,\beta},\;\beta\in I_{\alpha}\}$. Given a minimal enclosure $\mathfrak{V}_{\alpha},$ let $\tau_{\alpha}$ be the unique invariant state with support on $\mathfrak{V}_{\alpha}.$ Introducing the parameter
$m_{\alpha}=\mathrm{Tr}\left(L\tau_{\alpha}L^{*}\right)-\mathrm{Tr}\left(R\tau_{\alpha}R^{*}\right),$
we have by [5, Lema 4.2] that $m_{\alpha}$ is independent of the particular minimal enclosure $\mathfrak{V}_{\alpha}$ in $\mathfrak{X}_{\alpha}.$ Therefore, we can combine Proposition 3.4, Theorem 3.7 and the discussion above to state the property
Proposition 4.1.
Consider a coin $(L,R).$ We have a decomposition
$\mathfrak{h}=\left(\bigoplus_{\alpha\in A}\mathcal{Y}_{\alpha}\right)\oplus\mathcal{X},$ (3)
where $A$ is a finite set, $\mathcal{Y}_{\alpha}$ is a subspace of $\mathfrak{h}$ being the support of exactly one invariant state $\tau_{\alpha}$ for $\mathcal{L},$ $\mathcal{X}$ contains no enclosure, and
  • $\textmd{Tr}\left(L\tau_{\alpha}L^{*}\right)=\dfrac{1}{2}\Rightarrow$ the walk is $\tau_{\alpha}$-recurrent for all $\tau\in\mathcal{S}(\mathcal{Y}_{\alpha})$;
  • $\textmd{Tr}\left(L\tau_{\alpha}L^{*}\right)\neq\dfrac{1}{2}\Rightarrow$ the walk is $\tau_{\alpha}$-transient for all $\tau\in\mathcal{S}(\mathcal{Y}_{\alpha}).$
Once the decomposition on Equation (3) is obtained, we can say that $\mathcal{Y}_{\alpha}$ is either recurrent or transient.
The following result is due to [5] and will be crucial to our criteria. To state it, consider a coin $(L,R),$ and let $\mathcal{Y}$ be an enclosure for $\mathcal{L}.$ Denoting $p_{\mathcal{Y}}$ the orthogonal projection onto $\mathcal{Y},$ we can define the associated absorption operator (see [4]):
$A(\mathcal{Y}):=\lim_{n\rightarrow\infty}\mathcal{L}^{*n}(p_{\mathcal{Y}}).$
Corollary 4.2.
Consider a coin $(L,R)$ with the decomposition of $\mathfrak{h}$ given on Equation (3), and let $a_{\alpha}(\rho)=\mathbb{E}_{\rho}\left[\mathrm{Tr}\left(A(\mathcal{Y}_{\alpha})\rho_{0}\right)\right].$ Then
$\frac{X_{n}-X_{0}}{n}\stackrel{{\scriptstyle n\rightarrow\infty}}{{\longrightarrow}}\sum_{\alpha\in A}a_{\alpha}(\rho)\delta_{m_{\alpha}},$ (4)
where $\delta_{m_{\alpha}}$ is the Dirac measure concentrated in $m_{\alpha}.$
Let us start an OQW induced by a finite-dimensional coin $(L,R)$ and we assume that, in the $n$-th step, the density operator is $\rho_{n}.$ Note that each $\mathcal{Y}_{\alpha}$ on the Proposition is an enclosure, and any $\rho_{n}\in\mathcal{S}(\mathfrak{h})$ can be written in the form
$\rho_{n}=\sum_{\alpha\in A}a_{\alpha}^{(n)}\tau_{\alpha}^{(n)}+\tau^{(n)},$
where $\tau_{\alpha}^{(n)}$ is a positive operator with support on $\mathcal{Y}_{\alpha}$ and $\tau^{(n)}$ is either null or a positive operator with support on $\mathcal{X}$ as in Equation (3). There are three possibilities concerning recurrence when $\tau^{(0)}$ is non-null and its support is not disjoint from $\mathcal{X}.$
1. $(L,R)$ is recurrent for $\mathcal{Y}_{\alpha}$ for some $\alpha$ and supp$\left(\mathcal{L}(\rho)\right)\cap\mathcal{Y}_{\alpha}\neq\emptyset$: in this case, the walk is $\rho$-recurrent. Indeed, concerning the quantum trajectories $(\rho_{n})_{n}$, we could refresh the walk on some instant $n$ at some vertex $\lvert X_{n}\rangle,$ which is $\rho_{n}$-recurrent, since supp$\left(\rho_{n}\right)\cap\mathcal{Y}_{\alpha}\neq\emptyset$ for a sufficient large $n$. If there was some density $\sigma$ with supp$(\sigma)\cap\mathcal{X}$ and $\sigma$-transient, there would be some enclosure contained on $\mathcal{X},$ which is a contradiction by Proposition 3.4. The conclusion is held by the homogeneity of the coin.
2. $(L,R)$ is transient for all $\mathcal{Y}_{\alpha}$ satisfying supp$\left(\mathcal{L}(\rho)\right)\cap\mathcal{Y}_{\alpha}\neq\emptyset$: the walk is $\sigma$-transient for all $\sigma\in\mathcal{S}(\mathcal{X})$. Indeed, in this case, starting the walk on vertex $X_{0}=0,$ $X_{n}/n$ converges (when $n\rightarrow\infty$) to
$m=\sum_{\alpha\in A}a_{\alpha}(\rho)\delta_{m_{\alpha}}=\sum_{\alpha\in A}a_{\alpha}(\rho),$
as given by Equation (4), where the second equality follows from $m_{\alpha}\neq 0,$ since $(L,R)$ is $\mathcal{Y}_{\alpha}$ transient. Thus, $X_{n}$ converges either to $\infty$- ∞ or $\infty$,∞ , in probability, imposing that the walk is $\rho$-transient and thus $(L,R)$ is transient on $\mathcal{S}(\mathcal{X}).$
Therefore, we have shown the following recurrence criterion:
Theorem 4.3 (A Generalized Criteria).
Consider a coin $(L,R)$ of dimension $d<\infty.$ Consider the decomposition of $\mathfrak{h}$ as given in Equation (3), and $\tau_{1},\ldots,\tau_{m}$ the respectively invariant states of $\mathcal{L}$ in $\mathcal{Y}_{\alpha},\alpha=1,\ldots,m.$
$(1)\quad\mathrm{Tr}(L\tau_{j}L^{*})\neq\dfrac{1}{2}\;\forall j\in\{1,\ldots m\}\Rightarrow(L,R)\mbox{ is transient;}$;
$(2)\quad\mathrm{Tr}(L\tau_{j_{k}}L^{*})=\dfrac{1}{2}\;$ for a set of indices $j_{k}\in J\subset\{1,\ldots,m\}$ and $\mathrm{Tr}(L\tau_{j_{l}}L^{*})\neq\dfrac{1}{2}\;$ for the remaining indices $j_{l}\in J^{C},$ and $\sigma$ is a density such that supp$(\sigma)\cap\mathcal{X}\neq\emptyset.$
(2.1) supp$\left(\mathcal{L}(\sigma)\right)\cap\mathcal{Y}_{\alpha}\neq\emptyset$ for some $\alpha\in J,$ then (L,R) is
$\begin{cases}\rho\text{-transient for }\rho\in\mathcal{D}_{1}:=\mathcal{S}(\oplus_{k\in J^{C}}\mathcal{Y}_{k})\\ \rho\text{-recurrent for }\rho\in\mathcal{D}_{1}^{C}.\end{cases}$
(2.2) supp$\left(\mathcal{L}(\rho)\right)\cap\mathcal{Y}_{\alpha}=\emptyset$ for all $\alpha\in J,$ then (L,R) is
$\begin{cases}\rho\text{-transient for }\rho\in\mathcal{D}_{2}:=\mathcal{S}(\oplus_{k\in J^{C}}\mathcal{Y}_{k}\oplus\mathcal{X})\\ \rho\text{-recurrent for }\mathcal{D}_{2}^{C}\end{cases}$
$(3)\quad\mathrm{Tr}(L\tau_{\alpha}L^{*})=\dfrac{1}{2}\;$ for every $\alpha\Rightarrow$ the walk is recurrent.
5 Examples
Example 5.1.
Our first examples will consider lazy OQWs of dimension 2.
  1. 1.
    Let $m,n>0$ and assume that $b:=1-m^{2}-n^{2}\in(0,1).$ Then the coin $(L,B,R)$ induced by
    $L=\begin{bmatrix}0&0\\ 0&-n\end{bmatrix},\qquad B=\begin{bmatrix}-\sqrt{b}&m\\ m&\sqrt{b}\end{bmatrix},\qquad R=\begin{bmatrix}n&0\\ 0&0\end{bmatrix}$
    is well-defined.
    The only invariant state for the auxiliary map is $\rho_{\infty}=I/2.$ Therefore,
    $\mathrm{Tr}(L\rho_{\infty}L^{*})=\frac{n^{2}}{2}=\mathrm{Tr}(R\rho_{\infty}R^{*})\Rightarrow\;(L,B,R)\mbox{ is recurrent,}$,
    by Item (1) of Theorem 3.9.
    It is worth noting that if we assume $m=0,$ then the walk is transient by Item (2) of Theorem 3.9. More explicitly, in this case, for the item (2.2), $|l_{1}|=0,\;|l_{2}|=n,\;|r_{1}|=n,\;|r_{2}|=0.$
    $\Diamond$
  2. 2.
    Let $x_{1},x_{2}\in\mathbb{C},$
    $|x_{1}|\leq\sqrt{\frac{2}{3}},\quad|x_{2}|\leq\sqrt{\frac{3}{4}},$
    then consider the coin $(L,B,R),$ where
    $L=\begin{bmatrix}\frac{1}{\sqrt{3}}&0\\ 0&\frac{1}{2}\end{bmatrix},\qquad B=\begin{bmatrix}\sqrt{\frac{2}{3}-x_{1}^{2}}&0\\ 0&\sqrt{\frac{3}{4}-x_{2}^{2}}\end{bmatrix},\qquad R=\begin{bmatrix}x_{1}&0\\ 0&x_{2}\end{bmatrix}.$
    We are in Item (2) of Theorem 3.9, thus we conclude that
    $(1)\quad|x_{1}|=\frac{1}{\sqrt{3}}\mbox{ and }|x_{2}|=\frac{1}{2}$$\Rightarrow$ $(L,B,R)\mbox{ is recurrent,}$,
    $(2)\quad|x_{1}|\neq\frac{1}{\sqrt{3}}\mbox{ and }|x_{2}|\neq\frac{1}{2}$$\Rightarrow$ $(L,B,R)\mbox{ is transient,}$,
    $(3)\quad|x_{1}|=\frac{1}{\sqrt{3}}\mbox{ and }|x_{2}|\neq\frac{1}{2}$$\Rightarrow$ $(L,B,R)\mbox{ is transient with respect to }\sigma=\begin{bmatrix}0&0\\ 0&1\end{bmatrix}\mbox{ and it is}$
    recurrent with respect to all densities but $\mbox{recurrent with respect to all densities but }\sigma,$
    $(4)\quad|x_{1}|\neq\frac{1}{\sqrt{3}}\mbox{ and }|x_{2}|=\frac{1}{2}$$\Rightarrow$ $(L,B,R)\mbox{ is transient with respect to }\sigma=\begin{bmatrix}1&0\\ 0&0\end{bmatrix}\mbox{ and it is}$
    recurrent with respect to all densities but $\mbox{recurrent with respect to all densities but }\sigma.$
$\Diamond$
Example 5.2.
Consider the Lazy OQW induced by the following coin of dimension 3: $(L,B,R),$ where
$L=\frac{1}{8}\begin{bmatrix}2(1+\sqrt{2})&0&2(1-\sqrt{2})\\ 0&\sqrt{31}&0\\ 2(1-\sqrt{2})&0&2(1+\sqrt{2})\end{bmatrix},\quad B=\frac{1}{16}\begin{bmatrix}\sqrt{30}&2i&\sqrt{30}\\ 2&0&2\\ \sqrt{30}&-2i&\sqrt{30}\end{bmatrix},\quad R=\frac{1}{8}\begin{bmatrix}2(1-\sqrt{2})&0&2(1+\sqrt{2})\\ 0&\sqrt{31}&0\\ 2(1+\sqrt{2})&0&2(1-\sqrt{2})\end{bmatrix}.$
There is a unique invariant state $\rho_{\infty}$ given by
$\rho_{\infty}=\frac{1}{2}\begin{bmatrix}1&0&-1\\ 0&0&0\\ -1&0&1\end{bmatrix},$
and $\mathrm{Tr}\left(L\rho_{\infty}L^{*}\right)=1/2=\mathrm{Tr}\left(R\rho_{\infty}R^{*}\right).$ Therefore, this coin is recurrent by Theorem 3.7.
$\Diamond$
Since the invariant state also depends on the choice of $B,$ it is natural to ask if for a given transient coin $(L,B,R),$ there is another $B^{\prime}$ that makes $(L,B^{\prime},R)$ to be recurrent.
Example 5.3.
Firstly, let
$L=\frac{1}{7}\begin{bmatrix}1&0\\ 0&2\end{bmatrix},\qquad B=\frac{1}{7}\begin{bmatrix}-1&1\\ 2&0\end{bmatrix},\qquad R=\frac{1}{7}\begin{bmatrix}1&1\\ 0&1\end{bmatrix}.$
Then
$\rho_{\infty}=\frac{1}{3}\begin{bmatrix}1&0\\ 0&2\end{bmatrix},$
$\mathrm{Tr}\left(L\rho_{\infty}L^{*}\right)=3/7$ and $\mathrm{Tr}\left(R\rho_{\infty}R^{*}\right)=5/21.$ The coin $(L,B,R)$ is transient. However, if we take
$B^{\prime}=\frac{1}{\sqrt{35}}\begin{bmatrix}5&-1\\ 0&2\end{bmatrix},$
then the invariant state is
$\rho^{\prime}_{\infty}=\begin{bmatrix}1&0\\ 0&0\end{bmatrix},$
and $\mathrm{Tr}\left(L\rho_{\infty}L^{*}\right)=1/\sqrt{7}=\mathrm{Tr}\left(R\rho_{\infty}R^{*}\right).$ Thus the coin $(L,B^{\prime},R)$ is recurrent.
$\Diamond$
We remark that the choices of the coins above were taken in such a way that $L,B^{\prime}$ and $R$ had a common invariant eigenvector, namely $[1,0]^{T},$ which was not an eigenvector of $B.$ This is due to the hard job of, given non-trivial $R$ and $L,$ to find $B^{\prime}$ in such a way that $(L,B^{\prime},R)$ is recurrent. Attempt were conduct using mathematic software with many $L$’s and $R$’s having simple entries, but even though the entries of $B$’s that made $(L,B,R)$ to be transient were very long to write in this paper.
Example 5.4.
Let us consider the coin given in [5, Section 6.2]:
$R=\begin{bmatrix}\sqrt{\frac{3}{8}}&0&0&0\\ -\sqrt{\frac{p_{1}}{2}}&\frac{1}{\sqrt{2}}&0&0\\ -\sqrt{\frac{p_{2}}{2}}&0&\frac{1}{\sqrt{2}}&0\\ \sqrt{\frac{2p_{3}}{3}}&0&0&\frac{1}{\sqrt{3}}\end{bmatrix},\qquad L=\begin{bmatrix}\frac{1}{2\sqrt{2}}&0&0&0\\ \sqrt{\frac{p_{1}}{2}}&\frac{1}{\sqrt{2}}&0&0\\ \sqrt{\frac{p_{2}}{2}}&0&\frac{1}{\sqrt{2}}&0\\ -\sqrt{\frac{p_{3}}{3}}&0&0&\frac{\sqrt{2}}{\sqrt{3}}\end{bmatrix},$
where $p_{1},\;p_{2},p_{3}\geq 0$ and $p_{1}/2+p_{2}/2+p_{3}/2=1/2.$
The invariant states of the auxiliary map are of the form $\rho_{\infty}=\sigma_{1}+\sigma_{2},$ where
$\sigma_{1}=\omega_{1}\begin{bmatrix}0&0&0&0\\ 0&a&b&0\\ 0&b^{*}&1-a&0\\ 0&0&0&0\end{bmatrix},\quad\sigma_{2}=\omega_{2}\begin{bmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&1\end{bmatrix},\quad\omega_{i},\omega_{j}\geq 0,\;\;\omega_{1}+\omega_{2}=1.$
We have
$m_{1}=\mathrm{Tr}(L\sigma_{1}L^{*})=\frac{1}{2},\qquad m_{2}=\mathrm{Tr}(L\sigma_{2}L^{*})=\frac{2}{3}\neq\frac{1}{2}.$
By item (2) of Theorem 4.3, the walk is transient for
$\rho=\begin{bmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&1\end{bmatrix}$
and recurrent for all the remaining density operators on $\mathbb{C}^{4}$, for any appropriate choices of $p_{k},\;k=1,2,3.$
Example 5.5.
Consider an OQW induced by the coin $(L,R),$ where
$L=\begin{bmatrix}\frac{\sqrt{5}}{5}&0&-\frac{\sqrt{5}}{5}\\ 0&\frac{2\sqrt{5}}{5}&\frac{\sqrt{5}}{10}\\ 0&0&\frac{1}{2}\end{bmatrix},\qquad R=\begin{bmatrix}\frac{2\sqrt{5}}{5}&0&\frac{\sqrt{5}}{10}\\ 0&\frac{\sqrt{5}}{5}&-\frac{\sqrt{5}}{5}\\ 0&0&\frac{1}{2}\end{bmatrix}.$
The invariant states of the auxiliary map are of the form $\rho_{\infty}=\sigma_{1}+\sigma_{2},$ where
$\sigma_{1}=\omega_{1}\begin{bmatrix}1&0&0\\ 0&0&0\\ 0&0&0\end{bmatrix},\quad\sigma_{2}=\omega_{2}\begin{bmatrix}0&0&0\\ 0&1&0\\ 0&0&0\end{bmatrix},\quad\omega_{i},\omega_{j}\geq 0,\;\;\omega_{1}+\omega_{2}=1.$
We have
$m_{1}=\mathrm{Tr}(L\sigma_{1}L^{*})=\frac{\sqrt{5}}{5}\neq\frac{1}{2},\qquad m_{2}=\mathrm{Tr}(L\sigma_{2}L^{*})=\frac{2\sqrt{5}}{5}\neq\frac{1}{2}.$
By item (1) of Theorem 4.3, the walk is transient.
Example 5.6.
Let us consider the coin $(L,R)$ of dimension 4 given by
$R=\begin{bmatrix}\frac{\sqrt{2}}{2}&-\frac{\sqrt{5}}{4}&0&\frac{1}{4}\\ 0&\frac{\sqrt{2}}{4}&0&0\\ 0&0&\frac{\sqrt{2}}{2}&0\\ 0&0&0&\frac{\sqrt{6}}{4}\end{bmatrix},\qquad L=\begin{bmatrix}\frac{\sqrt{2}}{2}&\frac{\sqrt{5}}{4}&0&-\frac{1}{4}\\ 0&\frac{1}{2}&0&\frac{\sqrt{5}}{4}\\ 0&0&-\frac{\sqrt{2}}{2}&0\\ 0&0&0&\frac{\sqrt{3}}{4}\end{bmatrix},$
The invariant states of the auxiliary map are of the form $\rho_{\infty}=\sigma_{1}+\sigma_{2},$ where
$\sigma_{1}=\omega_{1}\begin{bmatrix}1&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{bmatrix},\quad\sigma_{2}=\omega_{2}\begin{bmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&1&0\\ 0&0&0&0\end{bmatrix},\quad\omega_{i},\omega_{j}\geq 0,\;\;\omega_{1}+\omega_{2}=1.$
We have
$m_{1}=\mathrm{Tr}(L\sigma_{1}L^{*})=\frac{1}{2},\qquad m_{2}=\mathrm{Tr}(L\sigma_{2}L^{*})=\frac{1}{2}.$
By item (3) of Theorem 4.3, the walk is recurrent.
6 Conclusion
We present new general recurrence criteria for classes of homogeneous OQWs, whose dynamics are governed by Kraus operators that are induced by a finite-dimensional coin.
In the context of these OQWs, we presented three distinct criteria for analyzing recurrence. Firstly, Theorem 3.7 introduces a criterion for ergodic Lazy coins, where the recurrence is determined by the operators $L$ and $R$ and by the only invariant state $\rho_{\infty}$ associated with the corresponding auxiliary map.
Next, Theorem 3.9 focuses on Lazy OQWs of dimension 2, offering a complete and easy-to-apply criterion. The simplicity of this criterion comes from the fact that it depends on the knowledge of the eigenvectors of the operators $L$, $B$, and $R$, allowing a practical analysis of the recurrence of these processes that are not necessarily ergodic.
Finally, Theorem 4.3 extends the analysis to non-lazy coins, using a decomposition of the vector space where the operators $L$, $B$, and $R$ act. By checking recurrence in vector subspaces that contain unique invariant states, this general criterion is extended to any initial density, significantly expanding the scope of recurrence analysis for OQWs in more complex configurations, since it is valid for any finite dimensional coin.
[…]
Acknowledgements This work was supported by Universidade Federal do Pampa (UNIPAMPA).
[…]
Conflict of interest The author declare that there is no conflict of interest.
[…]
Data Availability The datasets generated during the current study are available from the corresponding author on reasonable request.

### v2

@@
-Generalized Recurrence Criteria for Classes of Open Quantum Walks
+Recurrence Criteria for Reducible Homogeneous Open Quantum Walks on the Line
 Abstract
-In this paper, we study the recurrence of open quantum walks (OQWs) induced by finite-dimensional coins $(L,B,R)$.
-The focus is on homogeneous OQWs with a set of vertices $\mathbb{Z}$, the set of integers.
+In this paper, we study the recurrence of homogeneous Open Quantum Walks (HOQWs) induced by finite-dimensional coins $(L,B,R)$.
 We present three distinct recurrence criteria, each adapted to different types of coins.
-The first criterion was developed for a class of Lazy OQWs in any finite dimension, where the presented criterion is associated with an auxiliary map and its only invariant state, resulting in the first recurrence criterion for Lazy OQWs.
-The second one is restricted to Lazy OQWs of dimension 2, where we provide a complete characterization of the recurrence for this lower dimension.
+The first criterion applies to coins whose auxiliary map has a unique invariant state, resulting in the first recurrence criterion for Lazy HOQWs.
+The second one applies to Lazy HOQWs of dimension 2, where we provide a complete characterization of the recurrence for this low-dimensional case.
 Finally, we present a general criterion for finite-dimensional coins in the non-lazy case $(B=0)$, which generalizes many of the previously known results.
-This new criterion holds for irreducible and reducible OQWs through a decomposition of the Hilbert space where our quantum states act.
+This new criterion holds for irreducible and reducible HOQWs.
 keywords:
-Open Quantum Walks, Recurrence, Lazy walks, Coin, General Criterion
+Homogeneous Open Quantum Walks, Recurrence, Lazy Walks, Coin, Reducibility
 1 Introduction
-Open Quantum Walks (OQWs) represent a natural extension of classical Markov chains through open quantum systems.
-Originally proposed by [2], OQWs are models that capture the behavior of open quantum systems, in which the measurement of the particle position after each time step leads to a classical Markov chain when it is joined with the space of particle states.
-In this context, we consider homogeneous OQWs defined on the set of integers, that is, the graph of this type of random walk is the set of vertices $V=\mathbb{Z}$, where the linear operators $R$ and $L$ play the roles of transitions between the vertices of the quantum walk, which are constant and each jump only occurs from one vertex to its nearest neighbors.
-Homogeneous OQWs in the Lattice are explored and detailed in [6].
-Recurrence is an important statistic of any random walk, emerging as a central question of interest in the quantum context, and can be directly connected with the classical recurrence criteria for random walks in $\mathbb{Z}$.
-For example, for homogeneous classical random walks on $\mathbb{Z}$, it is well known that a walk is recurrent if and only if the transition parameters $r$ (rate of jumping to right) and $l$ (rate of jumping to left) are equal (or $r=1/2=l$ in the case $b=0$) [11].
-This classical random walk has its graph illustrated in Figure 1.
+Open Quantum Walks (OQWs) naturally extend classical Markov chains within the framework of open quantum systems.
+Originally proposed by [2], OQWs describe open quantum systems in which the repeated measurement of the particle’s position, together with the space of particle states, gives rise to a Markov chain.
+In this context, we consider homogeneous OQWs (HOQWs) defined on the set of integers, that is, the graph of this type of random walk is the set of vertices $V=\mathbb{Z}$, where 3 linear operators $L,B,R$ play the roles of transitions between the vertices of the quantum walk, which are constant and each jump only occurs from one vertex to its nearest neighbors, or to itself.
+HOQWs on more general lattices are explored and detailed in [7].
+Recurrence is an important probabilistic property of any random walk, often characterized by the expected number of returns to the initial vertex, emerging as a central question of interest in the quantum context, and can be directly connected with the classical recurrence criteria for random walks in $\mathbb{Z}$.
+For example, consider the classical homogeneous random walk on $\mathbb{Z}$ with graph given by Figure 1.
+It is well known that this lazy random walk with $r,l\geq 0,\;b>0,\;l+b+r=1,$ is recurrent if and only if the transition parameters $r$ (rate of jumping to right) and $l$ (rate of jumping to left) are equal.
+Also, if $b=0,$ then the walk is non-lazy and it is recurrent if and only if $r=1/2=l$ [14].
 […]
-Figure 1: Classical Lazy Random Walk on $\mathbb{Z}$.
-With this description, it is natural to seek criteria to determine under what conditions a homogeneous OQW is recurrent given $R$ and $L$.
-This problem has been addressed in the literature with the development of specific criteria, which are essential to understanding the dynamics and long-term behavior of these systems.
-A Criterion for PQ Matrices can be found in [10].
-Also, in that work the authors analyze the positive recurrence of this type of walk.
-From the analysis of the spectrum of $L^{*}L$ and $R^{*}R,$ it was showed in [7] that $R$ and $L$ having only singular values equal to $1/\sqrt{2}$ is a sufficient condition to guarantee that those OQWs are recurrent.
-More recently, it was obtained general criteria for recurrence of non-Lazy OQWs in dimension 2 and for irreducible finite-dimensional OQWs [9].
-In this article, we present a general criterion for the recurrence of homogeneous OQWs, based on the transition operators $R$ and $L$ when acting in finite-dimensional vector spaces, that is, we allow the OQW to be reducible in any finite dimension.
-Furthermore, we present the first recurrence criterion for lazy OQWs, using results obtained from recent studies on Central Limit Theorems for this kind of OQWs.
-A first Central Limit Theorem for OQWs was presented by [1].
-Here we will use more recent results.
-These works presented a Central Limit Theorem for Lazy OQWs [8] and one for any homogeneous OQW in Lattice when the degrees of freedom are finite [5].
+Figure 1: Classical Homogeneous Lazy Random Walk on $\mathbb{Z}$.
+In the context of discrete-time Markov chains, a state is called recurrent if, once the chain enters that state, the probability of returning to it is 1.
+This is equivalent to saying that the average number of returns to that state is infinite, indicating that the chain will revisit the state indefinitely [14].
+With this description, it is natural to seek criteria to determine under what conditions a HOQW is recurrent given $L,B,R.$ This problem has been addressed in the literature with the development of specific criteria, which are essential to understanding the dynamics and long-term behavior of these systems.
+See for instance [8, 13, 12].
+These works define recurrence in terms of probability of return, where a state is said to be recurrent if the probability of the walk returning to it is equal to 1.
+However, it has been shown that for OQWs, the two definitions of recurrence based on returning with probability 1 is not equivalent to the one based on the average number of returns being infinite [3].
+This property highlights a distinct characteristic of OQWs compared to classical Markov chains.
+Once this difference is understood, we discuss the recurrence criteria for OQWs, adopting the definition proposed in [3].
+Specifically, we analyze the recurrence of OQWs based on the average number of returns.
+Furthermore, recurrence criteria concerning the mean number of returns for higher-order HOQWs have been presented in [11].
+The authors of that work extend the analysis of recurrence to a more general class of open quantum walks, such as irreducible coins.
+In this article, we present more general criteria for the recurrence of lazy and non-lazy HOQWs, based on the transition operators $L,B,R$ when acting on finite-dimensional vector spaces, that is, we allow the HOQW to be reducible.
+Furthermore, we present the first recurrence criterion for lazy HOQWs, using results obtained from recent studies on Central Limit Theorems for this kind of HOQWs.
+A first Central Limit Theorem (CLT) for Open Quantum Walks (OQWs) was established in [1].
+In this work, we rely on more recent developments, including the CLT for Lazy OQWs presented in [10], as well as the general CLT for homogeneous OQWs (HOQWs) on lattices with finite internal degrees of freedom given in [6].
 This work is structured as follows.
 In Section 2 we define the open quantum walks induced by a coin and their transition probabilities.
-Section 3 is devoted to presenting a recurrence criterion for a class of Lazy OQWs, and one restricted to Lazy OQWs of dimension 2.
+Section 3 is devoted to presenting a recurrence criterion for a class of Lazy HOQWs, and one restricted to Lazy HOQWs of dimension 2.
 In Section 4 we focus on the case where the coin is non-lazy and we present a general recurrence criterion for finite-dimensional coins.
@@
 2 General Settings
-Let us consider a set of vertices $V$, which might be finite or countable infinite.
-Let $\mathcal{K}=\mathbb{C}^{V}$ be the state space of a quantum system with as many degrees of freedom as the number of vertices.
-In the case where $V$ is infinite countable, we assume $\mathcal{K}$ to be any separable Hilbert space with an orthonormal basis indexed by $V$, denote by $\left\{\lvert i\rangle\right\}_{i\in V}$.
+Let us consider the set of vertices $\mathbb{Z},$ and $\mathcal{K}=\mathbb{C}^{\mathbb{Z}}$ be the state space of a quantum system with as many degrees of freedom as the number of vertices.
 To emphasize the position of the particle of our quantum walk, we consider a quantum system described by the separable complex Hilbert space
-$\mathcal{H}=\bigoplus_{i\in V}\mathfrak{h}_{i},$
+$\mathcal{H}=\bigoplus_{i\in\mathbb{Z}}\mathfrak{h}_{i},$
 where the $\mathfrak{h}_{i}$ are separable Hilbert spaces.
-The label $i\in V$ represents the position of the particle and, when the particle is located at the vertex $i\in V$ , its internal state is encoded in the space $\mathfrak{h}_{i},$ as it will be detailed below.
-We consider the set of density operators over $\mathcal{H},$
-$\mathcal{D}(\mathcal{H})=\{\rho\in\mathcal{H}:\rho\geq 0,\mathrm{Tr}(\rho)=1\},$
+The label $i\in\mathbb{Z}$ represents the position of the particle and, when the particle is located at the vertex $i\in\mathbb{Z}$, its internal state is encoded in the space $\mathfrak{h}_{i},$ as it will be detailed below.
+We consider the set of density operators over a finite Hilbert space $\mathcal{W},$
+$\mathcal{D}(\mathcal{W})=\{\rho:\mathcal{W}\rightarrow\mathcal{W}:\;\rho\geq 0,\;\mathrm{Tr}(\rho)=1\},$
 and introduce
-$\mathcal{S}(\mathcal{H}\otimes\mathcal{K})=\left\{\tau\in\mathcal{I}(\mathcal{H}\otimes\mathcal{K}):\textstyle\tau=\displaystyle\sum_{j\in\mathbb{Z}}\rho(j)\otimes\lvert j\rangle\langle j\rvert\right\},$
-the set of density operators acting on $\mathcal{H}\otimes\mathcal{K}.$ Fixed a set of vertices $V,$ an open quantum random walk (OQW) on $V$ is a completely positive map
-$\Phi:\mathcal{H}\otimes\mathcal{K}\mapsto\mathcal{H}\otimes\mathcal{K},$
-given by
-$\Phi(\tau)=\sum_{i\in V}\Bigg{(}\sum_{j\in V}B_{i}^{j}\rho_{j}B_{i}^{j^{*}}\Bigg{)}\otimes|i\rangle\langle i|,\;\;\;\tau\in\mathcal{S}(\mathcal{H}\otimes\mathcal{K}),$
-where $B_{i}^{j}$ are bounded operators on $\mathcal{H}$ corresponding to the transition effect from site $j$ to $i$, and
-$\sum_{i\in V}B_{i}^{j*}B_{i}^{j}=I,\;\;\;j\in V,$
-which guarantees that $\mathrm{Tr}(\Phi(\tau))=\mathrm{Tr}(\tau)$, for every $\tau\in\mathcal{D}(\mathcal{H})$, meaning that $\Phi$ is trace-preserving, thus $\Phi$ preserves the set $\mathcal{S}(\mathcal{H}\otimes\mathcal{K}),$ i.e., $\Phi(\tau)\in\mathcal{S}(\mathcal{H}\otimes\mathcal{K})\;\forall\tau\in\mathcal{S}(\mathcal{H}\otimes\mathcal{K}).$
-In this work, we focus on the OQW model presented in [2], where the set of vertices is $V=\mathbb{Z}.$ For simplicity, we will say integer line, or simply the line, to refer to the set $\mathbb{Z}$ of integer numbers.
-An OQW on $\mathbb{Z}$ concerns the evolution of density operators of the form
-$\tau=\sum_{j\in\mathbb{Z}}\rho(j)\otimes\lvert j\rangle\langle j\rvert\;\in\;\mathcal{H}\otimes\mathbb{C}^{\mathbb{Z}},$
-where the $V=\mathbb{Z}$ represents the position of the walker and $\mathfrak{h}_{i}$ is the Hilbert space describing the internal degrees of freedom for site $\lvert i\rangle$, which we call the coin space and will be denoted by $\mathfrak{h}_{i}=\mathbb{C}^{k}$.
-Finally, we consider homogeneous lazy open quantum walks: at each step either the particle on site $\lvert j\rangle$ moves exactly one position to some of its two nearest-neighbours, in the left ($\lvert j-1\rangle$) or in the right ($\lvert j+1\rangle$), or it remains at same site ($\lvert j\rangle$), and the coin is constant, i.e., the same transition rule is applied at each site.
-Therefore, the linear operator $\Phi$ can be written by
-$\Phi(\tau)=\sum_{i\in\mathbb{Z}}\sum_{j=i-1}^{i+1}M_{i}^{j}\tau(M_{i}^{j})^{*},$
-where
-$M_{i}^{i-1}=L\otimes\lvert i-1\rangle\langle i\rvert,\quad M_{i}^{i}=B\otimes\lvert i\rangle\langle i\rvert,\quad M_{i}^{i+1}=R\otimes\lvert i+1\rangle\langle i\rvert.$
-The linear operators $L,B,R\in\mathbb{M}_{k}({\mathbb{C}})$ are matrices such that $L^{*}L+B^{*}B+R^{*}R=I_{k},$ and simplified by
-$\Phi(\tau)=\sum_{i\in\mathbb{Z}}\Big{(}L\rho(i+1)L^{*}+B\rho(i)B^{*}+R\rho(i-1)R^{*}\Big{)}\otimes|i\rangle\langle i|,\;\;\;\tau\in\mathcal{S}(\mathcal{H}\otimes\mathcal{K}).$ (1)
+$\mathcal{S}(\mathcal{H}\otimes\mathcal{K})=\left\{\tau\in\mathcal{I}(\mathcal{H}\otimes\mathcal{K}):\textstyle\tau=\displaystyle\sum_{j\in\mathbb{Z}}\rho(j)\otimes\lvert j\rangle\langle j\rvert,\quad\rho(j)\geq 0,\quad\sum_{j\in\mathbb{Z}}\mathrm{Tr}\left(\rho(j)\right)=1\right\},$
+the set of density operators acting on $\mathcal{H}\otimes\mathcal{K}.$ We are interested in HOQWs of the form
+$\Phi(\tau)=\sum_{i\in\mathbb{Z}}\Big{(}L\rho(i+1)L^{*}+B\rho(i)B^{*}+R\rho(i-1)R^{*}\Big{)}\otimes|i\rangle\langle i|,\;\;\;\tau\in\mathcal{S}(\mathcal{H}\otimes\mathcal{K}),$ (1)
+where $L,B,R\in\mathbb{M}_{d}({\mathbb{C}})$ satisfy $L^{*}L+B^{*}B+R^{*}R=I_{d}.$
 We say that the triple of matrices $(L,B,R)$ is a coin if eq.
@@
 If $B\neq 0,$ the quantum walk exhibits a lazy behavior, meaning that in the context of quantum walks there is a non-zero probability for the walker to stay in the same vertex after a step.
-We refer to the OQW above as an OQW induced by the coin $(L,B,R)$ in which the walk is inherently lazy.
+We refer to the HOQW above as a HOQW induced by the coin $(L,B,R)$ in which the walk is inherently lazy.
+Therefore, this coin represents a homogeneous, nearest-neighbor OQW on the integer line with the left and right transitions given by $L$ and $R$, respectively, and $B$ represents the rate of no jump at some instant.
 This lazy behavior introduces distinct properties compared to its non-lazy counterpart.
 Notably, the presence of the matrix $B$ influences the walk’s propensity to remain in its current vertex, thereby affecting the overall dynamics.
-If $B=0,$ then $\Phi$ will be called an OQW induced by a coin $(L,R)$ or an non-lazy OQW induced by a coin $(L,R).$
-The dynamic described above is represented by the graph in Figure 2.
+If $B=0,$ then $\Phi$ will be called a HOQW induced by a coin $(L,R)$ or a non-lazy HOQW induced by a coin $(L,R).$ The dynamic described above is represented by the graph in Figure 2.
 […]
 Figure 2: Lazy Quantum Walk on $\mathbb{Z}$.
-The OQW induced by a coin $(L,B,R)$ is the homogeneous, nearest-neighbor OQW on the integer line with the left and right transitions given by $L$ and $R$, respectively, and $B$ represents the rate of no jump at some instant.
-In this sense, the quantum trajectory of this kind of OQW on the line, induced by a coin $(L,B,R)$ and starting from a state $\tau$ of the form $\tau=\sum_{i\in\mathbb{Z}}\rho(i)\otimes|i\rangle\langle i|,$ is any path generated by the Markov chain $(X_{n},\tau_{n})_{n\geq 0}$, where $X_{n}$ denotes the position of the particle at time $n$ and $\tau_{n}$ its internal degree.
+The quantum trajectory of this kind of OQW on the line, induced by a coin $(L,B,R)$ and starting from a state $\tau$ of the form $\tau=\sum_{i\in\mathbb{Z}}\tau_{i}\otimes|i\rangle\langle i|,$ is any path generated by the Markov chain $(X_{n},\tau_{n})_{n\geq 0}$, where $X_{n}$ denotes the position of the particle at time $n$ and $\tau_{n}$ its internal degree.
 The transition probabilities are given by
-$\begin{split}\mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left(i-1,\frac{L\sigma L^{*}}{\mathrm{Tr}(L\sigma L^{*})}\right)\;\Bigg{|}\;(x_{n},\rho_{n})=(i,\sigma)\,\right)&=\mathrm{Tr}(L\sigma L^{*}),\\ \mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left(i,\frac{B\sigma B^{*}}{\mathrm{Tr}(B\sigma B^{*})}\right)\;\Bigg{|}\;(x_{n},\rho_{n})=(i,\sigma)\,\right)&=\mathrm{Tr}(B\sigma B^{*}),\\ \mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left(i+1,\frac{R\sigma R^{*}}{\mathrm{Tr}(R\sigma R^{*})}\right)\;\Bigg{|}\;(x_{n},\rho_{n})=(i,\sigma)\,\right)&=\mathrm{Tr}(R\sigma R^{*}).\end{split}$
-for every $i\in\mathbb{Z}$, $\sigma\in\mathcal{S}(\mathfrak{h}_{i})$, and initial law
-$\mathbb{P}\left(\,(X_{0},\rho_{0})=\left(i,\frac{\rho(i)}{\mathrm{Tr}\,\rho(i)}\right)\;\right)=\mathrm{Tr}(\rho(i)),$
+$\begin{split}\mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left(i-1,\frac{L\sigma L^{*}}{\mathrm{Tr}(L\sigma L^{*})}\right)\;\Bigg{|}\;(x_{n},\tau_{n})=(i,\sigma)\,\right)&=\mathrm{Tr}(L\sigma L^{*}),\\ \mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left(i,\frac{B\sigma B^{*}}{\mathrm{Tr}(B\sigma B^{*})}\right)\;\Bigg{|}\;(x_{n},\tau_{n})=(i,\sigma)\,\right)&=\mathrm{Tr}(B\sigma B^{*}),\\ \mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left(i+1,\frac{R\sigma R^{*}}{\mathrm{Tr}(R\sigma R^{*})}\right)\;\Bigg{|}\;(x_{n},\tau_{n})=(i,\sigma)\,\right)&=\mathrm{Tr}(R\sigma R^{*}),\end{split}$
+for every $i\in\mathbb{Z}$, $\sigma\in\mathcal{D}(\mathfrak{h})$, and initial law
+$\mathbb{P}\left(\,(X_{0},\tau_{0})=\left(i,\frac{\tau_{i}}{\mathrm{Tr}\,\tau_{i}}\right)\;\right)=\mathrm{Tr}(\tau_{i}),$
 and all other transition probabilities are null.
+Therefore, at time $n$, if the system is located at site $\lvert i\rangle$ with an initial density $\rho\otimes\lvert i\rangle\langle i\rvert$, then at the next time step $n+1$, one of the following occurs:
+  • The particle moves to site $\lvert i-1\rangle$ (left) with probability $\operatorname{Tr}(L_{i}\rho L_{i}^{*})$, and the density evolves to
+    $\frac{L_{i}\rho L_{i}^{*}}{\operatorname{Tr}(L_{i}\rho L_{i}^{*})}\otimes\lvert i-1\rangle\langle i-1\rvert;$
+  • The particle remains at site $\lvert i\rangle$ with probability $\operatorname{Tr}(B_{i}\rho B_{i}^{*})$, and the new density matrix is
+    $\frac{B_{i}\rho B_{i}^{*}}{\operatorname{Tr}(B_{i}\rho B_{i}^{*})}\otimes\lvert i\rangle\langle i\rvert;$
+  • The particle moves to site $\lvert i+1\rangle$ (right) with probability $\operatorname{Tr}(R_{i}\rho R_{i}^{*})$, leading to the new density matrix
+    $\frac{R_{i}\rho R_{i}^{*}}{\operatorname{Tr}(R_{i}\rho R_{i}^{*})}\otimes\lvert i+1\rangle\langle i+1\rvert.$
 3 Recurrence criteria for Lazy Open Quantum Walks
 3.1 Faithfulness of Density Operators
-Density operators play a fundamental role in open quantum random walks, providing a statistical description of this type of quantum system.
-Faithful and non-faithful density operators are two categories that arise in the study of quantum states, being characterized by their positive definiteness and positive-semidefiniteness, respectively.
+Density operators play a fundamental role in open quantum systems.
+In this work, they provide statistical descriptions of HOQWs.
+Faithful and non-faithful density operators are characterized by their positive definiteness and positive semi-definiteness, respectively, when they evolve in finite-dimensional systems.
 Faithful density operators are defined as being positive definite.
-This means that the eigenvalues of a faithful density operator are strictly positive and it is invertible.
-In essence, faithfulness ensures a faithful representation of the quantum system, capturing the full richness of its physical properties.
-On the other hand, non-faithful density operators are positive-semidefinite, thus they have $\lambda=0$ as one of their eigenvalues.
-Thereby, that there may be ambiguity in the quantum state description, as multiple states can correspond to the same non-faithful density operator.
-The positivity-semidefiniteness criterion relaxes the strictness of positive definiteness, permitting a broader range of quantum state representations.
+This means that the eigenvalues of a faithful density operator are strictly positive, and the operator is invertible.
+Essentially, faithfulness ensures a complete representation of the quantum system, capturing the full richness of its physical properties.
+On the other hand, non-faithful density operators are positive semi-definite, thus they have $\lambda=0$ as one of their eigenvalues.
+The presence of zero eigenvalues implies that the non-faithful density operator is not invertible and may represent specific mixed states or situations where information loss occurs.
+The positive semi-definiteness criterion relaxes the strictness of positive definiteness, allowing for a broader range of quantum state representations, while still maintaining the one-to-one correspondence between quantum states and density operators.
 An important subset of the set of non-faithful density operators is composed of the pure density operators.
-A pure density is a density $\rho$ that is a projection of rank 1, i.e., there exists some vector $\lvert u\rangle\in\mathbb{C}^{k}$ such that $\rho=\lvert u\rangle\langle u\rvert$.
+A pure density is a density $\rho$ that is a projection of rank 1, i.e., there exists some vector $\lvert u\rangle\in\mathbb{C}^{d}$ such that $\rho=\lvert u\rangle\langle u\rvert$.
 Let us consider an OQW starting the walk on vertex $\lvert i\rangle$ with initial density operator $\rho.$ For evolution on time $n=0,1,2,\ldots,$ we define
@@
 where
-$\Phi^{n}(\rho\otimes\lvert i\rangle\langle i\rvert)=\sum_{s\in V}\rho_{n}(s)\otimes\lvert s\rangle\langle s\rvert.$
+$\Phi^{n}(\rho\otimes\lvert i\rangle\langle i\rvert)=\sum_{s\in\mathbb{Z}}\rho_{n}(s)\otimes\lvert s\rangle\langle s\rvert.$
 With this notation,
 $p_{ji;\rho}(n)=p_{n}(\rho\otimes|i\rangle\to|j\rangle)=\mathrm{\mathrm{Tr}}(\rho_{n}(j)\otimes\lvert j\rangle\langle j\rvert)=\mathrm{\mathrm{Tr}}\left(\Phi^{n}(\rho\otimes\lvert i\rangle\langle i\rvert)(I\otimes\lvert j\rangle\langle j\rvert)\right)$
-denote the probability of being at site $j$ at step $n,$ given that we started at site $i,$ with initial density $\rho$ concentrated at $i.$ Therefore, the dynamics starts with a density operator $\rho$ concentrated at some vertex $\lvert i\rangle,$ takes the evolution up to time ($n$) through the linear operator $\Phi^{n},$ producing a new density operator
+denote the probability of being at site $\lvert j\rangle$ at step $n,$ given that we started at site $\lvert i\rangle,$ with initial density $\rho$ concentrated at $\lvert i\rangle.$ Therefore, the dynamics starts with a density operator $\rho$ concentrated at some vertex $\lvert i\rangle,$ takes the evolution up to time ($n$) through the linear operator $\Phi^{n},$ producing a new density operator
 $\rho_{n}=\sum_{s}\rho_{n}(s)\otimes\lvert s\rangle\langle s\rvert=\Phi^{n}(\rho\otimes\lvert i\rangle\langle i\rvert),\quad\mathrm{Tr}\left(\sum_{s}\rho_{n}(s)\right)=1,$
@@
 which represents the data concentrated at vertex $\lvert j\rangle$ at time $n.$
-Fixed some $i\in V,$ we denote the set of density operators acting on $\mathfrak{h}_{i}$ by
-$\mathcal{S}(\mathfrak{h}_{i})=\left\{\rho:\mathfrak{h}_{i}\rightarrow\mathfrak{h}_{i}:\;\mathrm{Tr}(\rho)=1,\;\rho=\rho^{*};\;\rho\geq 0\right\}.$
-Consider an OQW on some set of vertices $V$, let $\lvert i\rangle\in V$ and $\rho\in\mathcal{S}(\mathfrak{h}_{i})$. We say that a vertex $\lvert i\rangle$ is
+Consider an OQW on some set of vertices $\mathbb{Z}$, let $i\in\mathbb{Z}$ and $\rho\in\mathcal{D}(\mathfrak{h})$. We say that a vertex $\lvert i\rangle$ is
   1. 1.
     $\rho$-recurrent if
-    $\sum_{n=0}^{\infty}p_{ii;\rho}(n)=\infty.$
-    Otherwise, $i$ is said to be $\rho$-transient;
+    $\sum_{n=0}^{\infty}p_{ii;\rho}(n)=\infty.$ (2)
+    Otherwise, $\lvert i\rangle$ is said to be $\rho$-transient;
   2. 2.
-    recurrent, if $i$ is $\rho-$recurrent for all $\rho\in\mathcal{S}(\mathfrak{h}_{i});$
+    recurrent, if $\lvert i\rangle$ is $\rho-$recurrent for all $\rho\in\mathcal{D}(\mathfrak{h});$
   3. 3.
-    transient, if $i$ is $\rho-$transient for all $\rho\in\rho\in\mathcal{S}(\mathfrak{h}_{i}).$
+    transient, if $\lvert i\rangle$ is $\rho-$transient for all $\rho\in\rho\in\mathcal{D}(\mathfrak{h}).$
 Concerning all vertices, an OQW is called: recurrent if every vertex is recurrent; transient if every vertex is transient.
-The following Proposition describes $\rho$-recurrence in a concentrated vertex $\lvert i\rangle\in V$ when assumed that $\lvert i\rangle$ is recurrent for some density and $\mathfrak{h}_{i}$ has a finite internal degree of freedom.
+Note that equation (2) computes the mean number of returns to site $\lvert i\rangle.$
+The following Proposition describes $\rho$-recurrence in a concentrated vertex $i\in\mathbb{Z}$ when assumed that $\lvert i\rangle$ is recurrent for some density and $\mathfrak{h}$ has a finite internal degree of freedom.
 An immediate consequence will also be given for the case where $\lvert i\rangle$ is $\sigma$-transient for some non-faithful density.
 Theorem 3.1.
-Consider an OQW on some set of vertices $V,$ let $i\in V$ and assume that $\mathfrak{h}_{i}$ is finite-dimensional.
-If $\lvert i\rangle$ is $\tau$-recurrent for some $\tau\in\mathcal{S}(\mathfrak{h}_{i}),$ then
+Consider an OQW on some set of vertices $V,$ let $i\in V$ and assume that $\mathfrak{h}$ is finite-dimensional.
+If $\lvert i\rangle$ is $\tau$-recurrent for some $\tau\in\mathcal{D}(\mathfrak{h}),$ then
   1. 1.
-    vertex $\lvert i\rangle$ is $\rho$-recurrent for all faithful $\rho\in\mathcal{S}(\mathfrak{h}_{i})$;
+    vertex $\lvert i\rangle$ is $\rho$-recurrent for all faithful $\rho\in\mathcal{D}(\mathfrak{h})$;
   2. 2.
-    there exists a non-faithful density operator $\tilde{\rho}$ such that $\lvert i\rangle$ is $\tilde{\rho}$-recurrent;
-  3. 3.
-    there exists a pure density operator $\lvert w\rangle\langle w\rvert$ such that the OQW is $\lvert w\rangle\langle w\rvert$-recurrent, where $\lvert w\rangle$ is an unitary vector of $\mathfrak{h}_{i}.$
+    there exists a pure density operator $\lvert w\rangle\langle w\rvert$ such that the OQW is $\lvert w\rangle\langle w\rvert$-recurrent, where $\lvert w\rangle$ is an unit vector of $\mathfrak{h}.$
 Proof.
-1. Let $\rho\in\mathcal{S}(\mathfrak{h}_{i})$ be faithful. By the compactness of $\mathcal{S}(\mathfrak{h}_{i}),$ there exists $c>0$ such that $\rho>c\tau,$ thus
-$\begin{split}\sum_{n=0}^{\infty}p_{ii;\rho}(n)=&\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\rho\otimes\lvert i\rangle\langle i\rvert\right)\right(I\otimes\lvert i\rangle\langle i\rvert)]\\ >&c\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\tau\otimes\lvert i\rangle\langle i\rvert\right)(I\otimes\lvert i\rangle\langle i\rvert)\right]\\ =&c\sum_{n=0}^{\infty}p_{ii;\tau}(n)\\ =&\infty.\end{split}$
-2. Let $\rho\in\mathcal{S}(\mathfrak{h}_{i})$ be non-faithful. By the Spectral Theorem, $\rho$ can be written as
-$\rho=\sum_{x=1}^{n}\lambda_{x}\lvert x\rangle\langle x\rvert,$ (2)
-where the vectors $\lvert x\rangle$ are the eigenvectors of $\rho$ with corresponding eigenvalues $\lambda_{x}$.
-Since $\rho$ is non-faithful, there is at least one null eigenvalue and the remainder of eigenvalues are non-negative summing $1$ Thus, (2) can be rewritten as
-$\rho=\sum_{x\in S}\lambda_{x}\lvert x\rangle\langle x\rvert,\quad S\varsubsetneq\{1,\ldots,n\}.$
-Take a sequence of positive numbers $(\alpha_{r})_{r\in R},$ where $R:=\{1,\ldots,n\}/S\neq\emptyset,$ whose sum is $1$
-Defining
-$\rho_{X}=\sum_{x\in S}\frac{\lambda_{x}}{2}\lvert x\rangle\langle x\rvert+\sum_{x\in R}\frac{\alpha_{x}}{2}\lvert x\rangle\langle x\rvert=\sum_{x=1}^{n}\frac{\tilde{\alpha}_{x}}{2}\lvert x\rangle\langle x\rvert,\quad\tilde{\alpha}_{x}=\begin{cases}\lambda_{x},&\mbox{if }x\in S\\ \alpha_{x},&\mbox{if }x\in R\end{cases},$
-we get by the first item that $\lvert i\rangle$ is $\rho_{X}$-recurrent, since $\rho_{X}$ is faithful.
-Now, define
-$\rho_{Y}=\sum_{x\in R}\alpha_{x}\lvert x\rangle\langle x\rvert,$
-which is a non-faithful density operator. Since $2\rho_{X}=\rho+\rho_{Y},$ we obtain by linearity the equality
-$\sum_{n=0}^{\infty}p_{ii;\rho}(n)+\sum_{n=0}^{\infty}p_{ii;\rho_{Y}}(n)=\sum_{n=0}^{\infty}\left(p_{ii;\rho}(n)+p_{ii;\rho_{Y}}(n)\right)=2\sum_{n=0}^{\infty}p_{ii;\rho_{X}}(n).$
-The sum on the right hand diverges once $\lvert i\rangle$ is $\rho_{X}$-recurrent, implying that at least one of the sums on the left-hand side diverges.
-Therefore, $\lvert i\rangle$ is $\rho$-recurrent or $\rho_{Y}$-recurrent, both being non-faithful.
-3. Let $\tilde{\rho}$ be one non-faithful density given by item (2), then we have
-$\tilde{\rho}=\sum_{j}\lambda_{j}\lvert e_{j}\rangle\langle e_{j}\rvert,$
-where each $\lambda_{j}\geq 0$ is an eigenvalue of $\tilde{\rho}$ with corresponding eigenvector $\lvert e_{j}\rangle.$ If $\tilde{\rho}$ has only one eigenvector, then we are done.
-If $\tilde{\rho}$ has at least two eigenvectors, then
-$\infty=\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\rho\otimes\lvert i\rangle\langle i\rvert\right)\right(I\otimes\lvert i\rangle\langle i\rvert)]=\sum_{j,\lambda_{j}\neq 0}\left(\lambda_{j}\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\lvert e_{j}\rangle\langle e_{j}\rvert\otimes\lvert i\rangle\langle i\rvert\right)\right(I\otimes\lvert i\rangle\langle i\rvert)]\right),$
-where at least one of the members of the sum must be infinity, and some $\lambda_{j}$ is positive. This completes the proof.
+Let $\rho\in\mathcal{D}(\mathfrak{h})$ be faithful. Since the state space $\mathcal{D}(\mathfrak{h})$ is compact in the trace-norm topology, the function
+$f(\tau)=\inf\{c>0\mid\rho>c\tau\}$
+is continuous and thus attains its minimum and maximum. In particular, there exists $c>0$ such that
+$\rho>c\tau,\quad\forall\rho\in\mathcal{D}(\mathfrak{h}).$
+thus
+$\begin{split}\sum_{n=0}^{\infty}p_{ii;\rho}(n)=&\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\rho\otimes\lvert i\rangle\langle i\rvert\right)\right(I\otimes\lvert i\rangle\langle i\rvert)]\\ >&c\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\tau\otimes\lvert i\rangle\langle i\rvert\right)(I\otimes\lvert i\rangle\langle i\rvert)\right]\\ =&c\sum_{n=0}^{\infty}p_{ii;\tau}(n)\\ =&\infty,\end{split}$
+giving the first item.
+For the second one, let
+$\tau=\sum_{j}\lambda_{j}\lvert e_{j}\rangle\langle e_{j}\rvert,$
+where each $\lambda_{j}>0$ is an eigenvalue of $\tau$ with corresponding eigenvector $\lvert e_{j}\rangle.$ Then
+$\infty=\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\tau\otimes\lvert i\rangle\langle i\rvert\right)\right(I\otimes\lvert i\rangle\langle i\rvert)]=\sum_{j}\left(\lambda_{j}\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\lvert e_{j}\rangle\langle e_{j}\rvert\otimes\lvert i\rangle\langle i\rvert\right)\right(I\otimes\lvert i\rangle\langle i\rvert)]\right),$
+where at least one of the members of the sum must be infinity. This completes the proof.
 $\square$
@@
 Corollary 3.2.
-Let $\Phi$ be an OQW on some set of vertices $V$ and $i\in V,$ and $\dim(\mathfrak{h}_{i})<\infty.$ If $\lvert i\rangle$ is $\tau$-transient for some faithful $\tau\in\mathcal{S}(\mathfrak{h}_{i}),$ then $\lvert i\rangle$ is transient.
+Let $\Phi$ be an OQW on some set of vertices $V,$ $i\in V,$ and $\dim(\mathfrak{h})<\infty.$ If $\lvert i\rangle$ is $\tau$-transient for some faithful $\tau\in\mathcal{D}(\mathfrak{h}),$ then $\lvert i\rangle$ is transient.
 Another direct consequence of Theorem 3.1 is related to the homogeneous case.
 Corollary 3.3.
-Let $\Phi$ be a homogeneous OQW on $\mathbb{Z}$ with $\dim(\mathfrak{h})=d<\infty.$ If $\Phi$ is $\tau$-recurrent for some $\tau\in\mathcal{S}(\mathfrak{h}),$ then
+Let $\Phi$ be a HOQW on $\mathbb{Z}$ with $\dim(\mathfrak{h})=d<\infty.$ If $\Phi$ is $\tau$-recurrent for some $\tau\in\mathcal{D}(\mathfrak{h}),$ then
   1. 1.
-    $\Phi$ is $\rho$-recurrent for all faithful $\rho\in\mathcal{S}(\mathfrak{h})$;
+    $\Phi$ is $\rho$-recurrent for all faithful $\rho\in\mathcal{D}(\mathfrak{h})$;
   2. 2.
-    there exists a non-faithful density operator $\tilde{\rho}$ such that $\Phi$ is $\tilde{\rho}$-recurrent;
-  3. 3.
-    there exists a pure density operator $\lvert w\rangle\langle w\rvert$ such that $\Phi$ is $\lvert w\rangle\langle w\rvert$-recurrent, where $\lvert w\rangle$ is an unitary vector of $\mathbb{C}^{d}.$
+    there exists a pure density operator $\lvert w\rangle\langle w\rvert$ such that $\Phi$ is $\lvert w\rangle\langle w\rvert$-recurrent, where $\lvert w\rangle$ is an unit vector of $\mathbb{C}^{d}.$
 3.2 Enclosures and Invariant States
-Firstly, a subspace $\mathcal{Y}$ of $\mathfrak{h}$ is called an enclosure for a quantum channel $\mathcal{L}$ if for every positive $\sigma\in L^{1}(\mathfrak{h}),$ we have
+Firstly, a subspace $\mathcal{Y}$ of $\mathfrak{h}$ is called an enclosure for a quantum channel $\mathcal{L}$ if for every $\sigma\in\mathcal{D}\mathfrak{h}),$ we have
 $\mathrm{supp}(\sigma)\subset\mathcal{Y}\Rightarrow\mathrm{supp}\left(\mathcal{L}(\sigma)\right)\subset\mathcal{Y}.$
@@
 For more on enclosures of OQWs, we refer the reader to [6, 7, 3].
-Following this property and using the homogeneity of OQWs induced by a coin, we split $\mathcal{S}(\mathfrak{h})$ into two subsets $\mathcal{S}(\mathcal{Y})$ and $\mathcal{S}(\mathcal{X})$ in such a way that the walk is transient for $\mathcal{S}(\mathcal{Y})$ and recurrent for $\mathcal{S}(\mathcal{X}),$ where $\mathcal{Y}$ is a enclosure of $\mathcal{L}:$
+Following this property and using the homogeneity of HOQWs induced by a coin, we split $\mathfrak{h}$ into two subsets $\mathcal{Y}$ and $\mathcal{X}$ in such a way that the walk is transient for $\mathcal{D}(\mathcal{Y})$ and recurrent everywhere else, with $\mathcal{Y}$ being an enclosure of the quantum channel with Kraus operators $L,B,R.$
 Proposition 3.4.
-Consider an OQW induced by a coin $(L,B,R)$ of dimension $d<\infty$. We are in one (and only one) of the following situations:
+Consider a HOQW induced by a coin $(L,B,R)$ of dimension $d$. We are in one (and only one) of the following situations:
   • $(L,B,R)$ is recurrent;
   • $(L,B,R)$ is transient;
-  • $(L,B,R)$ is recurrent with respect to all densities except for a convex set $\mathcal{S}(\mathcal{Y}).$ In this case, $\mathcal{Y}$ is an enclosure of $\mathcal{L},$
-    $\mathfrak{h}=\mathcal{Y}\oplus\mathcal{X}$
+  • $(L,B,R)$ is recurrent with respect to all densities except for a convex set $\mathcal{D}(\mathcal{Y}).$ In this case, $\mathcal{Y}$ is an enclosure of $\mathcal{L},$
+    $\mathfrak{h}=\mathcal{Y}\oplus\mathcal{X}$ (3)
     for some subspace $\mathcal{X}$ of $\mathfrak{h},$ with $\mathcal{Y},\mathcal{X}\neq\{0\}.$
 Proof.
+We claim $(L,B,R)$ is $\rho$-transient if and only if it is $\rho_{n}$-transient for every $n,$ where $\rho_{n}$ is the density at the $n$-th step of the walk.
+Indeed, let us suppose the walk starts at vertex $\lvert 0\rangle$ with initial density $\rho.$ For instance, if the first jump is to $\lvert-1\rangle,$ the density, after this first jump, is
+$\rho_{1}=\frac{L\rho L^{*}}{\mathrm{Tr}(L\rho L^{*})}.$
+Since the walk is homogeneous, it is $\rho$-transient if and only if it is $\rho_{1}$-transient. After $n$ jumps, the density is
+$\rho_{n}=\frac{J_{\pi}\rho J_{\pi}^{*}}{\mathrm{Tr}(J_{\pi}\rho J_{\pi}^{*})},$ (4)
+where $J_{\pi}$ is a concatenation of $n$ operators, each one belonging to $\{L,B,R\},$ proving the claim, once this is valid for any $n=1,2,3,\ldots$
+Moreover, by equation (4) we have that $\cup_{n}\;\text{supp}(\rho_{n})$ is a subspace of $\mathfrak{h}.$ Joining all the possible initial states in which $(L,B,R)$ is transient, we obtain a decomposition (3), where $\mathcal{Y}$ or $\mathcal{X}$ may be null, and by definition $\mathcal{Y}$ is an enclosure.
 Let us suppose that $(L,B,R)$ is neither recurrent nor transient.
-By Proposition 3.1, there exist unitary vectors $\lvert u\rangle,\lvert v\rangle\in\mathfrak{h}$ such that the walk is $\lvert u\rangle\langle u\rvert$-recurrent and $\lvert v\rangle\langle v\rvert$-transient.
-Let $\mathcal{Y}$ be the subset of $\mathfrak{h}$ such that $\mathcal{S}(\mathcal{Y})$ is the set of densities of $\mathcal{B}(\mathfrak{h})$ in which $(L,B,R)$ is transient, thus we pick $\rho\in\mathcal{S}(\mathcal{Y})$ and we can affirm that $\rho_{n}\in\mathcal{S}(\mathcal{Y}),$ where $(X_{n},\rho_{n})_{n}$ are the quantum trajectories of the OQW.
-Indeed, if we had $\rho_{n}\in\mathcal{S}(\mathfrak{h})\setminus\mathcal{S}(\mathcal{Y})$ for some $n,$ then we could refresh the walk and the vertex $\lvert X_{n}\rangle$ would be $\rho_{n}$-recurrent, contradicting the hypothesis of homogeneity, thus we must have that $\mathcal{Y}$ is an enclosure of $\mathcal{L}.$ Therefore, take $\mathcal{X}=\mathcal{Y}^{\perp}$ and the trichotomy holds.
+By Theorem 3.1, there exist unit vectors $\lvert u\rangle,\lvert v\rangle\in\mathfrak{h}$ such that the walk is $\lvert u\rangle\langle u\rvert$-recurrent and $\lvert v\rangle\langle v\rvert$-transient.
+Let $\mathcal{Y}$ be the subspace of $\mathfrak{h}$ such that $\mathcal{D}(\mathcal{Y})$ is the set of densities of $\mathcal{B}(\mathfrak{h})$ in which $(L,B,R)$ is transient, thus we pick $\rho\in\mathcal{D}(\mathcal{Y})$ and we can affirm that $(L,B,R)$ is $\rho_{n}$-transient for every $n.$ Finally, take $\mathcal{X}=\mathcal{Y}^{\perp}$ and the trichotomy holds.
 $\square$
@@
 Theorem 3.5.
-Consider a Lazy OQW induced by a coin $(L,B,R)$ of dimension $d<\infty.$ If the auxiliary map $\mathcal{L}$ admits only one invariant state $\rho_{\infty},$ then
-$\frac{X_{n}}{n}\stackrel{{\scriptstyle n\rightarrow\infty}}{{\longrightarrow}}m,\qquad\text{where }\qquad m=\mathrm{Tr}\left(L\rho_{\infty}L^{*}\right)-\mathrm{Tr}\left(R\rho_{\infty}R^{*}\right),$
+Consider a Lazy HOQW induced by a coin $(L,B,R)$ of dimension $d.$ If the auxiliary map $\mathcal{L}$ admits only one invariant state $\rho_{\infty},$ then
+$\frac{X_{n}}{n}\stackrel{{\scriptstyle n\rightarrow\infty}}{{\longrightarrow}}m,\qquad\text{where }\qquad m=\mathrm{Tr}\left(R\rho_{\infty}R^{*}\right)-\mathrm{Tr}\left(L\rho_{\infty}L^{*}\right),$
 in probability.
 Proof.
-This result can be obtained by using a Doob decomposition as described in [8, Equation 27] and [1, Theorem 5.2] for a Martingale $M_{n}$ satisfying
+This result can be obtained by using a Doob decomposition as described in [10, Equation 27] and [1, Theorem 5.2] for a Martingale $M_{n}$ satisfying
 $X_{n}-n.m=X_{0}+M_{n}+\beta_{n},$
@@
 $\square$
-Theorem 3.6 (Chung-Fuchs Theorem for OQWs with ergodic auxiliary map).
-Consider an OQW induced by a coin $(L,B,R)$ such that $\mathcal{Y}$ is an enclosure Suppose that the initial density operator is $\rho\in\mathcal{S}(\mathcal{Y})$ with quantum trajectories $(X_{n},\rho_{n})_{n\geq 0}$ and $X_{n}/n\rightarrow 0$ in probability, then $(L,B,R)$ is $\sigma$-recurrent for some $\sigma\in\mathcal{S}(\mathcal{Y}).$
+Theorem 3.6 (Chung-Fuchs Theorem for HOQWs with ergodic auxiliary map).
+Consider a HOQW induced by a coin $(L,B,R),\;\mathcal{L}$ its auxiliary map and $\mathcal{Y}$ an enclosure.
+Suppose that the initial density operator is $\rho\in\mathcal{D}(\mathcal{Y})$ with quantum trajectories $(X_{n},\rho_{n})_{n\geq 0}$ and $X_{n}/n\rightarrow 0$ in probability, then $(L,B,R)$ is $\sigma$-recurrent for some $\sigma\in\mathcal{D}(\mathcal{Y}).$
 Proof.
-The proof is analogous to [9, Theorem 7] and we will proceed by remarking on how it can be generalized.
-Firstly, the maximum appearing in [9, Lemma 6] can be taken in $\mathcal{D}=\mathcal{S}(\mathcal{Y}).$ Indeed, if the initial state is $\rho\in\mathcal{S}(\mathcal{Y}),$ then $\rho_{n}\in\mathcal{S}(\mathcal{Y})$ for every $n,$ thereby the maximum taken in the proof of [9, Lemma 6] can be reduced to $\mathcal{S}(\mathcal{Y}).$ Therefore, for the same reason, the density operator considered in [9, Theorem 7] belongs to $\mathcal{S}(\mathcal{Y}).$
+The proof is analogous to [11, Theorem 7] and we will proceed by remarking on how it can be generalized.
+Firstly, the maximum appearing in [11, Lemma 6] can be taken in $\mathcal{D}=\mathcal{D}(\mathcal{Y}).$ Indeed, if the initial state is $\rho\in\mathcal{D}(\mathcal{Y}),$ then $\rho_{n}\in\mathcal{D}(\mathcal{Y})$ for every $n,$ thereby the maximum taken in the proof of [11, Lemma 6] can be reduced to $\mathcal{D}(\mathcal{Y}).$ Therefore, by the same reason, the density operator considered in [11, Theorem 7] belongs to $\mathcal{D}(\mathcal{Y}),$ proving the Theorem for the non-lazy case.
+Moreover, the lazy case holds as an extension of this proof, since all the results work for homogeneous OQWs on $\mathbb{Z}$.
 $\square$
+The following theorem presents the first recurrence criterion for non-irreducible HOQWs of arbitrary finite dimension.
+This result extends [11, Corollary 8 and Proposition 8] for reducible CTOQWs and can also be applied to Lazy HOQWs.
 Theorem 3.7 (Recurrence Criteria).
-Let us consider an OQW on $\mathbb{Z}$ induced by a coin $(L,B,R)$ of dimension $d$ such that the auxiliary map $\mathcal{L}$ is ergodic.
+Let us consider a HOQW on $\mathbb{Z}$ induced by a coin $(L,B,R)$ of dimension $d$ such that the auxiliary map $\mathcal{L}$ is ergodic.
 Then $\mathcal{L}$ has a unique invariant state $\rho_{\infty}$ and
@@
 We suppose that the walk is neither recurrent nor transient.
-In this case, the walk is recurrent for some density and transient for all densities in some $\mathcal{S}(\mathcal{Y}),$ where $\mathcal{Y}$ is an enclosure (Proposition 3.4).
-Moreover, the unique invariant state $\rho_{\infty}$ of $\mathcal{L}$ must belong to $\mathcal{S}(\mathcal{Y}).$
-If $\;\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)=\textmd{Tr}\left(R\rho_{\infty}R^{*}\right),$ then $X_{n}/n\rightarrow 0$ by Theorem 3.5, thus $(L,B,R)$ is $\rho$-recurrent for some $\rho\in\mathcal{S}(\mathcal{Y})$ by Theorem 3.6, which is a contradiction.
+In this case, the walk is recurrent for some density and transient for all densities in some $\mathcal{D}(\mathcal{Y}),$ where $\mathcal{Y}$ is an enclosure (Proposition 3.4).
+Moreover, the unique invariant state $\rho_{\infty}$ of $\mathcal{L}$ must belong to $\mathcal{D}(\mathcal{Y}).$
+If $\;\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)=\textmd{Tr}\left(R\rho_{\infty}R^{*}\right),$ then $X_{n}/n\rightarrow 0$ by Theorem 3.5, thus $(L,B,R)$ is $\rho$-recurrent for some $\rho\in\mathcal{D}(\mathcal{Y})$ by Theorem 3.6, which is a contradiction.
 […]
@@
 See example 5.3 below.
-3.3 General Recurrence Criteria for OQWs induced by a coin $(L,B,R)$ of dimension 2
-The following Proposition is a generalisation of [9, Proposition 12, item (3)].
-We can prove it in an analogous demonstration given by [9], however, we would like to do it in terms of the non-ergodicity of the auxiliary map induced by this coin.
-Proposition 3.8.
-Consider an OQW induced by a coin $(L,B,R)$ of dimension 2 such that $L,B,R$ share two unitary linearly independent eigenvectors.
-Then for any $\lvert u\rangle$ orthogonal to one of these two eigenvectors, we have that $\lvert u\rangle$ is also a common eigenvector of $L,B,R.$
-Proof.
-Suppose $L,B,R$ share two unitary and linearly independent unitary eigenvectors $\lvert v\rangle,\lvert w\rangle,$ then $\tau_{1}=\lvert v\rangle\langle v\rvert$ and $\tau_{2}=\lvert w\rangle\langle w\rvert$ are distinct invariant states of the auxiliary map $\mathcal{L},$ thus $\mathcal{L}$ is not ergodic.
-This means that there are two minimal enclosures $\mathcal{Y}_{1},\mathcal{Y}_{2},$ and thus $\mathcal{Y}_{1}\perp\mathcal{Y}_{2}.$ The dimension of the coin gives that $\mathcal{Y}_{1}$ and $\mathcal{Y}_{2}$ are one-dimensional and therefore there exist $\lvert u_{1}\rangle,\lvert u_{2}\rangle\in\mathbb{C}^{2}$ such that $\mathcal{Y}_{1}=\mbox{span}(\lvert u_{1}\rangle)$ and $\mathcal{Y}_{2}=\mbox{span}(\lvert u_{2}\rangle).$ Thereby $\lvert u_{1}\rangle$ and $\lvert u_{2}\rangle$ are common orthonormal eigenvectors of $L,B,R.$
-$\square$
+3.3 General Recurrence Criteria for HOQWs induced by a coin $(L,B,R)$ of dimension 2
+The following theorem is due to [4].
 Theorem 3.8.
-Consider an OQW on $\mathbb{Z}$ induced by a coin $(L,B,R)$ of dimension 2.
+For a finite quantum channel $\mathcal{L}:\mathcal{B}(\mathfrak{h})\rightarrow\mathcal{B}(\mathfrak{h})$, the following are equivalent:
+i.
+$\mathcal{L}$ is ergodic;
+ii.
+$\mathcal{L}$ does not have two invariant subspaces $S_{1}\neq\{0\}$ and $S_{2}\neq\{0\}$ such that $S_{1}\cap S_{2}=\{0\};$
+iii.
+$\mathcal{L}$ has a minimal invariant subspace, that is, a subspace $S\neq\{0\}$ such that $S\subset S^{\prime}$ for every invariant subspace $S^{\prime}\neq\{0\}.$
+Building upon the established preceding proposition, we will advance to the following theorem, delineating a comprehensive criterion for 2-dimensional coins, extending [11, Theorem 17] to encompass the Lazy case.
+This criterion will be structured into two distinct parts: the first adhering to the logic of the previous theorem, and the second articulated in terms of the eigenvalues of the operators of the coin, independent of the invariant densities of the auxiliary map.
+Theorem 3.9.
+Consider a HOQW on $\mathbb{Z}$ induced by a coin $(L,B,R)$ of dimension 2.
 (1) If $L,B,R$ have at most one common eigenvector, let $\rho_{\infty}$ be the unique invariant density of the auxiliary map. Then, we have
@@
 $(1.2)\quad\textmd{Tr}(L\rho_{\infty}L^{*})=\textmd{Tr}(R\rho_{\infty}R^{*})\Rightarrow(L,B,R)\mbox{ is recurrent}.$
-(2) If $L,B,R$ have two linearly independent eigenvectors in common, $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ unitaries, then they satisfy $\lvert u_{1}\rangle\bot\lvert u_{2}\rangle,$ we can put
+(2) If $L,B,R$ have two linearly independent eigenvectors in common, $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ unit, then they satisfy $\lvert u_{1}\rangle\bot\lvert u_{2}\rangle,$ we can put
 $L\lvert u_{j}\rangle=l_{j}\lvert u_{j}\rangle,\qquad B\lvert u_{j}\rangle=b_{j}\lvert u_{j}\rangle,\qquad R\lvert u_{j}\rangle=r_{j}\lvert u_{j}\rangle,\qquad j=1,2,$
@@
 $(2.2)\quad|l_{1}|\neq|r_{1}|\mbox{ and }|l_{2}|\neq|r_{2}|$$\Rightarrow$ $(L,B,R)\mbox{ is transient,}$,
-$(2.3)\quad|l_{i}|=|r_{i}|\mbox{ and }|l_{j}|\neq|r_{j}|$$\Rightarrow$ $(L,B,R)\mbox{ is transient with respect to }\sigma_{j}=\lvert u_{j}\rangle\langle u_{j}\rvert\mbox{ and it is}$
-recurrent with respect to all densities but $\mbox{recurrent with respect to all densities but }\sigma_{j}.$
+$(2.3)\quad|l_{i}|=|r_{i}|\mbox{ and }|l_{j}|\neq|r_{j}|$$\Rightarrow$ $(L,B,R)\mbox{ is transient with respect to }\sigma_{j}=\lvert u_{j}\rangle\langle u_{j}\rvert\mbox{ and it is recurrent}$
+[…]
 Proof.
 If the elements of the coin do not share at least two orthogonal eigenvectors, then the auxiliary map is ergodic, thus item (1) is a particular case of Theorem 3.7.
-Now we suppose that $L,B,R$ share two linearly independent eigenvectors, then we can apply Proposition 3.8 to obtain $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ common orthonormal eigenvectors of $L,B,R.$ We assume that $|l_{1}|=|r_{1}|$ and $|l_{2}|=|r_{2}|,$ then for $\rho=\lvert u\rangle\langle u\rvert,$ where $\lvert u\rangle$ is some common eigenvector of $L,B,R,$ we have $p_{j,j+1;\rho}(n)=p_{j,j-1;\rho}(n)=w$ for all $j\in\mathbb{Z}$ and all $n\geq 0$ for some constant $w$ and therefore, in this case, we have a classical homogeneous walk which is recurrent.
-Since it is valid for any common eigenvector, the walk is recurrent by Proposition 3.1.
-Analogously, if $|l_{1}|\neq|r_{1}|$ and $|l_{2}|\neq|r_{2}|,$ then for $\rho_{1}=\lvert u_{1}\rangle\langle u_{1}\rvert,\;\rho_{2}=\lvert u_{2}\rangle\langle u_{2}\rvert$ where $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ are orthonormal common eigenvectors of $L,B,R,$ we have $p_{j,j+1;\rho}(n)<p_{j,j-1;\rho}(n)$ (or $>$) for all $j\in\mathbb{Z}$ and all $n\geq 0$ and therefore in this case we have a classical homogeneous walk which is transient for both $\rho_{1},\rho_{2}$.
-Since the walk is transient for two distinct eigenvectors, the walk is transient by Proposition 3.1.
+Now we suppose that $L,B,R$ share two linearly independent eigenvectors, then we can apply Theorem 3.8 to obtain $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ common orthonormal eigenvectors of $L,B,R.$ We assume that $|l_{1}|=|r_{1}|$ and $|l_{2}|=|r_{2}|,$ then for $\rho=\lvert u\rangle\langle u\rvert,$ where $\lvert u\rangle$ is some common eigenvector of $L,B,R,$ we have $p_{j,j+1;\rho}(n)=p_{j,j-1;\rho}(n)=w$ for all $j\in\mathbb{Z}$ and all $n\geq 0$ for some constant $w$ and therefore, in this case, we have a classical homogeneous walk which is recurrent.
+If $(L,B,R)$ was not transient, then we would be in third situation of proposition 3.4, with
+$\mathfrak{h}=\mathcal{Y}\oplus\mathcal{X},\quad\text{where}\quad\text{dim}(\mathcal{Y})\geq 1,\;\;\text{ and }\;\;\text{dim}(\mathcal{X})=2,$
+a contradiction. Therefore, we are in the first situation of proposition 3.4 and the walk is recurrent.
+Analogously, if $|l_{1}|\neq|r_{1}|$ and $|l_{2}|\neq|r_{2}|,$ then for $\rho_{1}=\lvert u_{1}\rangle\langle u_{1}\rvert,$ and $\rho_{2}=\lvert u_{2}\rangle\langle u_{2}\rvert,$ where $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ are orthonormal common eigenvectors of $L,B,R,$ we have $p_{j,j+1;\rho}(n)<p_{j,j-1;\rho}(n)$ (or $>$) for all $j\in\mathbb{Z}$ and all $n\geq 0$ and therefore in this case we have a classical homogeneous walk which is transient for both $\rho_{1},\rho_{2}$.
+Since the walk is transient for two distinct eigenvectors, the walk is transient by corollary 3.2.
 For the last case, $|l_{i}|=|r_{i}|$ and $|l_{j}|\neq|r_{j}|,$ $(L,B,R)$ is $\rho_{i}$-recurrent and $\rho_{j}$-transient, where $\rho_{i}=\lvert u_{i}\rangle\langle u_{i}\rvert$ and $\rho_{j}=\lvert u_{j}\rangle\langle u_{j}\rvert.$ This situation is a particular case of the last item in Proposition 3.4.
@@
 4 A General Recurrence Criteria for Open Quantum Walks
-Consider a coin $(L,R)$ of dimension $n<\infty.$ By [5], we can obtain a decomposition of $\mathfrak{h}$ in the following way:
-$\mathfrak{h}=\mathfrak{R}\oplus\mathfrak{T},$
-where
-$\mathfrak{R}=\sup\{\mbox{supp}(\omega)\;|\;\omega\mbox{ is an invariant state for }\mathcal{L}\}.$
+Following the approach described in [6, Section 4], as well as the references therein, we can obtain for a coin $(L,R)$ a decomposition
+$\mathfrak{h}=\mathfrak{R}\oplus\mathfrak{T},\qquad\mathfrak{R}=\sup\{\mbox{supp}(\omega)\;|\;\omega\mbox{ is an invariant state for }\mathcal{L}\}.$
 Since $\mathfrak{h}$ is finite-dimensional, any minimal enclosure is included in $\mathfrak{R},$ and is the support of a unique extremal invariant state.
@@
 […]
-Further, we can write
-$\mathfrak{R}=\bigoplus_{\alpha\in A}\mathfrak{X}_{\alpha}=\bigoplus_{\alpha\in A}\bigoplus_{\beta\in I_{\alpha}}\mathfrak{V}_{\alpha,\beta},$
-where every $\mathfrak{X}_{\alpha}$ either is a minimal enclosure or assumes a decomposition (not unique) as the sum of mutually orthogonal isomorphic minimal enclosures $\{\mathfrak{V}_{\alpha,\beta},\;\beta\in I_{\alpha}\}$.
-Given a minimal enclosure $\mathfrak{V}_{\alpha},$ let $\tau_{\alpha}$ be the unique invariant state with support on $\mathfrak{V}_{\alpha}.$ Introducing the parameter
-$m_{\alpha}=\mathrm{Tr}\left(L\tau_{\alpha}L^{*}\right)-\mathrm{Tr}\left(R\tau_{\alpha}R^{*}\right),$
-we have by [5, Lema 4.2] that $m_{\alpha}$ is independent of the particular minimal enclosure $\mathfrak{V}_{\alpha}$ in $\mathfrak{X}_{\alpha}.$ Therefore, we can combine Proposition 3.4, Theorem 3.7 and the discussion above to state the property
+Further, every $\mathfrak{X}_{\alpha}$ either is a minimal enclosure or assumes a decomposition (not unique) as the sum of mutually orthogonal isomorphic minimal enclosures $\{\mathfrak{V}_{\alpha,\beta},\;\beta\in I_{\alpha}\}:$
+$\mathfrak{X}_{\alpha}=\bigoplus_{\beta\in I_{\alpha}}\mathfrak{V}_{\alpha,\beta}.$
+Given a minimal enclosure $\mathfrak{V}_{\beta},$ let $\tau_{\beta}$ be the unique invariant state with support on $\mathfrak{V}_{\beta}.$ Introducing the parameter
+$m_{\beta}=\mathrm{Tr}\left(L\tau_{\beta}L^{*}\right)-\mathrm{Tr}\left(R\tau_{\beta}R^{*}\right),$
+we have by [6, Lemma 4.2] that $m_{\beta}$ is independent of the particular minimal enclosure $\mathfrak{V}_{\beta}$ in $\mathfrak{X}_{\alpha}.$ Therefore, we can combine Proposition 3.4, Theorem 3.7 and the discussion above to state the property
 Proposition 4.1.
@@
 where $A$ is a finite set, $\mathcal{Y}_{\alpha}$ is a subspace of $\mathfrak{h}$ being the support of exactly one invariant state $\tau_{\alpha}$ for $\mathcal{L},$ $\mathcal{X}$ contains no enclosure, and
-  • $\textmd{Tr}\left(L\tau_{\alpha}L^{*}\right)=\dfrac{1}{2}\Rightarrow$ the walk is $\tau_{\alpha}$-recurrent for all $\tau\in\mathcal{S}(\mathcal{Y}_{\alpha})$;
-  • $\textmd{Tr}\left(L\tau_{\alpha}L^{*}\right)\neq\dfrac{1}{2}\Rightarrow$ the walk is $\tau_{\alpha}$-transient for all $\tau\in\mathcal{S}(\mathcal{Y}_{\alpha}).$
+  • $\textmd{Tr}\left(L\tau_{\alpha}L^{*}\right)=\dfrac{1}{2}\Rightarrow$ the walk is $\tau_{\alpha}$-recurrent for all $\tau\in\mathcal{D}(\mathcal{Y}_{\alpha})$;
+  • $\textmd{Tr}\left(L\tau_{\alpha}L^{*}\right)\neq\dfrac{1}{2}\Rightarrow$ the walk is $\tau_{\alpha}$-transient for all $\tau\in\mathcal{D}(\mathcal{Y}_{\alpha}).$
 Once the decomposition on Equation (5) is obtained, we can say that $\mathcal{Y}_{\alpha}$ is either recurrent or transient.
-The following result is due to [5] and will be crucial to our criteria.
+The following result is due to [6] and will be crucial to our last criterion.
 To state it, consider a coin $(L,R),$ and let $\mathcal{Y}$ be an enclosure for $\mathcal{L}.$ Denoting $p_{\mathcal{Y}}$ the orthogonal projection onto $\mathcal{Y},$ we can define the associated absorption operator (see [5]):
@@
 Corollary 4.2.
-Consider a coin $(L,R)$ with the decomposition of $\mathfrak{h}$ given on Equation (3), and let $a_{\alpha}(\rho)=\mathbb{E}_{\rho}\left[\mathrm{Tr}\left(A(\mathcal{Y}_{\alpha})\rho_{0}\right)\right].$ Then
+Consider a coin $(L,R)$ with the decomposition of $\mathfrak{h}$ given on Equation (5), and let $a_{\alpha}(\rho)=\mathbb{E}_{\rho}\left[\mathrm{Tr}\left(A(\mathcal{Y}_{\alpha})\rho_{0}\right)\right].$ Then we have the convergence in law
 $\frac{X_{n}-X_{0}}{n}\stackrel{{\scriptstyle n\rightarrow\infty}}{{\longrightarrow}}\sum_{\alpha\in A}a_{\alpha}(\rho)\delta_{m_{\alpha}},$ (6)
 where $\delta_{m_{\alpha}}$ is the Dirac measure concentrated in $m_{\alpha}.$
-Let us start an OQW induced by a finite-dimensional coin $(L,R)$ and we assume that, in the $n$-th step, the density operator is $\rho_{n}.$ Note that each $\mathcal{Y}_{\alpha}$ on the Proposition is an enclosure, and any $\rho_{n}\in\mathcal{S}(\mathfrak{h})$ can be written in the form
-$\rho_{n}=\sum_{\alpha\in A}a_{\alpha}^{(n)}\tau_{\alpha}^{(n)}+\tau^{(n)},$
-where $\tau_{\alpha}^{(n)}$ is a positive operator with support on $\mathcal{Y}_{\alpha}$ and $\tau^{(n)}$ is either null or a positive operator with support on $\mathcal{X}$ as in Equation (3).
-There are three possibilities concerning recurrence when $\tau^{(0)}$ is non-null and its support is not disjoint from $\mathcal{X}.$
+We remark that the convergence in law brought by corollary 4.2 can be upgraded to a convergence a.s., as demonstrated in [9, Theorem 3.5.2].
+Let us start a HOQW induced by a coin $(L,R).$ There are two possibilities concerning recurrence when $\text{supp}(\rho_{0})\not\perp\mathcal{X}:$
 1.
-$(L,R)$ is recurrent for $\mathcal{Y}_{\alpha}$ for some $\alpha$ and supp$\left(\mathcal{L}(\rho)\right)\cap\mathcal{Y}_{\alpha}\neq\emptyset$: in this case, the walk is $\rho$-recurrent.
-Indeed, concerning the quantum trajectories $(\rho_{n})_{n}$, we could refresh the walk on some instant $n$ at some vertex $\lvert X_{n}\rangle,$ which is $\rho_{n}$-recurrent, since supp$\left(\rho_{n}\right)\cap\mathcal{Y}_{\alpha}\neq\emptyset$ for a sufficient large $n$.
-If there was some density $\sigma$ with supp$(\sigma)\cap\mathcal{X}$ and $\sigma$-transient, there would be some enclosure contained on $\mathcal{X},$ which is a contradiction by Proposition 3.4.
-The conclusion is held by the homogeneity of the coin.
+$(L,R)$ is recurrent for $\mathcal{Y}_{\alpha}$ for some $\alpha$ and supp$\left(\mathcal{L}^{n}(\rho_{0})\right)\not\perp\mathcal{Y}_{\alpha}$ for some $n\in\mathbb{N}$: in this case, the walk is $\rho_{0}$-recurrent.
+Indeed, concerning the quantum trajectories $(\rho_{n})_{n}$, we could refresh the walk on some instant $n$ at some vertex $\lvert X_{n}\rangle,$ which is $\rho_{n}$-recurrent, since supp$\left(\mathcal{L}^{n}(\rho_{0})\right)\not\perp\mathcal{Y}_{\alpha}$ for a sufficiently large $n$.
+If there was some density $\sigma$ with supp$(\sigma)\not\perp\mathcal{X}$ and $\sigma$-transient, there would be some enclosure not orthogonal to $\mathcal{X},$ which is a contradiction by Proposition 3.4.
+The conclusion is held by the homogeneity of the coin, since, in this case, the walk either is $\rho_{s}$-recurrent for every $s=0,1,2,\ldots$ or it is $\rho_{s}$-transient for every $s=0,1,2,\ldots$, analogously to the proof of Proposition 3.4.
 2.
-$(L,R)$ is transient for all $\mathcal{Y}_{\alpha}$ satisfying supp$\left(\mathcal{L}(\rho)\right)\cap\mathcal{Y}_{\alpha}\neq\emptyset$: the walk is $\sigma$-transient for all $\sigma\in\mathcal{S}(\mathcal{X})$.
+$(L,R)$ is transient for all $\mathcal{Y}_{\alpha}$ satisfying supp$\left(\mathcal{L}^{n}(\rho_{0})\right)\not\perp\mathcal{Y}_{\alpha}$ for every $n\in\mathbb{N}$: the walk is $\sigma$-transient for all $\sigma\in\mathcal{D}(\mathcal{X})$.
 Indeed, in this case, starting the walk on vertex $X_{0}=0,$ $X_{n}/n$ converges (when $n\rightarrow\infty$) to
@@
 as given by Equation (6), where the second equality follows from $m_{\alpha}\neq 0,$ since $(L,R)$ is $\mathcal{Y}_{\alpha}$ transient.
-Thus, $X_{n}$ converges either to $\infty$- ∞ or $\infty$,∞ , in probability, imposing that the walk is $\rho$-transient and thus $(L,R)$ is transient on $\mathcal{S}(\mathcal{X}).$
+Thus, $X_{n}$ converges either to $\infty$- ∞ or $\infty$,∞ , in probability, imposing that the walk is $\rho$-transient and thus $(L,R)$ is transient on $\mathcal{D}(\mathcal{X}).$
 Therefore, we have shown the following recurrence criterion:
 Theorem 4.3 (A Generalized Criteria).
-Consider a coin $(L,R)$ of dimension $d<\infty.$ Consider the decomposition of $\mathfrak{h}$ as given in Equation (3), and $\tau_{1},\ldots,\tau_{m}$ the respectively invariant states of $\mathcal{L}$ in $\mathcal{Y}_{\alpha},\alpha=1,\ldots,m.$
+Consider a coin $(L,R)$ of dimension $d.$ Consider the decomposition of $\mathfrak{h}$ as given in Equation (5), and $\tau_{1},\ldots,\tau_{m}$ the respectively invariant states of $\mathcal{L}$ in $\mathcal{Y}_{\alpha},\alpha=1,\ldots,m.$
 $(1)\quad\mathrm{Tr}(L\tau_{j}L^{*})\neq\dfrac{1}{2}\;\forall j\in\{1,\ldots m\}\Rightarrow(L,R)\mbox{ is transient;}$;
-$(2)\quad\mathrm{Tr}(L\tau_{j_{k}}L^{*})=\dfrac{1}{2}\;$ for a set of indices $j_{k}\in J\subset\{1,\ldots,m\}$ and $\mathrm{Tr}(L\tau_{j_{l}}L^{*})\neq\dfrac{1}{2}\;$ for the remaining indices $j_{l}\in J^{C},$ and $\sigma$ is a density such that supp$(\sigma)\cap\mathcal{X}\neq\emptyset.$
-(2.1) supp$\left(\mathcal{L}(\sigma)\right)\cap\mathcal{Y}_{\alpha}\neq\emptyset$ for some $\alpha\in J,$ then (L,R) is
-$\begin{cases}\rho\text{-transient for }\rho\in\mathcal{D}_{1}:=\mathcal{S}(\oplus_{k\in J^{C}}\mathcal{Y}_{k})\\ \rho\text{-recurrent for }\rho\in\mathcal{D}_{1}^{C}.\end{cases}$
-(2.2) supp$\left(\mathcal{L}(\rho)\right)\cap\mathcal{Y}_{\alpha}=\emptyset$ for all $\alpha\in J,$ then (L,R) is
-$\begin{cases}\rho\text{-transient for }\rho\in\mathcal{D}_{2}:=\mathcal{S}(\oplus_{k\in J^{C}}\mathcal{Y}_{k}\oplus\mathcal{X})\\ \rho\text{-recurrent for }\mathcal{D}_{2}^{C}\end{cases}$
+$(2)\quad\mathrm{Tr}(L\tau_{j_{k}}L^{*})=\dfrac{1}{2}\;$ for a set of indices $j_{k}\in J\subset\{1,\ldots,m\}$ and $\mathrm{Tr}(L\tau_{j_{l}}L^{*})\neq\dfrac{1}{2}\;$ for the remaining indices $j_{l}\in J^{C},$ and $\sigma$ is a density such that supp$(\sigma)\not\perp\mathcal{X}.$
+(2.1) supp$\left(\mathcal{L}^{n}(\sigma)\right)\not\perp\mathcal{Y}_{\alpha}$ for some $\alpha\in J$ and some $n\in\mathbb{N},$ then $(L,R)$ is
+$\begin{cases}\rho\text{-transient for }\rho\in\mathcal{D}_{1}:=\mathcal{D}(\oplus_{k\in J^{C}}\mathcal{Y}_{k})\\ \rho\text{-recurrent for }\rho\in\mathcal{D}_{1}^{C}.\end{cases}$
+(2.2) supp$\left(\mathcal{L}^{n}(\rho)\right)\perp\mathcal{Y}_{\alpha}$ for all $\alpha\in J$ and all $n\in\mathbb{N},$ then $(L,R)$ is
+$\begin{cases}\rho\text{-transient for }\rho\in\mathcal{D}_{2}:=\mathcal{D}(\oplus_{k\in J^{C}}\mathcal{Y}_{k}\oplus\mathcal{X})\\ \rho\text{-recurrent for }\mathcal{D}_{2}^{C}\end{cases}$
 $(3)\quad\mathrm{Tr}(L\tau_{\alpha}L^{*})=\dfrac{1}{2}\;$ for every $\alpha\Rightarrow$ the walk is recurrent.
+Remark 4.4.
+It is worth observing that $\mathcal{D}_{2}$ in Theorem 4.3 actually arises from an enclosure due to the orthogonality between supp$\left(\mathcal{L}^{n}(\rho)\right)$ and $\mathcal{Y}_{\alpha}$; otherwise, it would contradict Proposition 3.4.
 5 Examples
 Example 5.1.
-Our first examples will consider lazy OQWs of dimension 2.
+Our first examples will consider lazy HOQWs of dimension 2.
   1. 1.
@@
 Example 5.2.
-Consider the Lazy OQW induced by the following coin of dimension 3: $(L,B,R),$ where
+Consider the Lazy HOQW induced by the following coin of dimension 3: $(L,B,R),$ where
 $L=\frac{1}{8}\begin{bmatrix}2(1+\sqrt{2})&0&2(1-\sqrt{2})\\ 0&\sqrt{31}&0\\ 2(1-\sqrt{2})&0&2(1+\sqrt{2})\end{bmatrix},\quad B=\frac{1}{16}\begin{bmatrix}\sqrt{30}&2i&\sqrt{30}\\ 2&0&2\\ \sqrt{30}&-2i&\sqrt{30}\end{bmatrix},\quad R=\frac{1}{8}\begin{bmatrix}2(1-\sqrt{2})&0&2(1+\sqrt{2})\\ 0&\sqrt{31}&0\\ 2(1+\sqrt{2})&0&2(1-\sqrt{2})\end{bmatrix}.$
@@
 Example 5.4.
-Let us consider the coin given in [5, Section 6.2]:
+Let us consider the coin given in [6, Section 6.2]:
 $R=\begin{bmatrix}\sqrt{\frac{3}{8}}&0&0&0\\ -\sqrt{\frac{p_{1}}{2}}&\frac{1}{\sqrt{2}}&0&0\\ -\sqrt{\frac{p_{2}}{2}}&0&\frac{1}{\sqrt{2}}&0\\ \sqrt{\frac{2p_{3}}{3}}&0&0&\frac{1}{\sqrt{3}}\end{bmatrix},\qquad L=\begin{bmatrix}\frac{1}{2\sqrt{2}}&0&0&0\\ \sqrt{\frac{p_{1}}{2}}&\frac{1}{\sqrt{2}}&0&0\\ \sqrt{\frac{p_{2}}{2}}&0&\frac{1}{\sqrt{2}}&0\\ -\sqrt{\frac{p_{3}}{3}}&0&0&\frac{\sqrt{2}}{\sqrt{3}}\end{bmatrix},$
@@
 Example 5.5.
-Consider an OQW induced by the coin $(L,R),$ where
+Consider a HOQW induced by the coin $(L,R),$ where
 $L=\begin{bmatrix}\frac{\sqrt{5}}{5}&0&-\frac{\sqrt{5}}{5}\\ 0&\frac{2\sqrt{5}}{5}&\frac{\sqrt{5}}{10}\\ 0&0&\frac{1}{2}\end{bmatrix},\qquad R=\begin{bmatrix}\frac{2\sqrt{5}}{5}&0&\frac{\sqrt{5}}{10}\\ 0&\frac{\sqrt{5}}{5}&-\frac{\sqrt{5}}{5}\\ 0&0&\frac{1}{2}\end{bmatrix}.$
@@
 6 Conclusion
-We present new general recurrence criteria for classes of homogeneous OQWs, whose dynamics are governed by Kraus operators that are induced by a finite-dimensional coin.
-In the context of these OQWs, we presented three distinct criteria for analyzing recurrence.
+We present new general recurrence criteria for classes of HOQWs, whose dynamics are governed by Kraus operators that are induced by a finite-dimensional coin.
+In the context of these HOQWs, we presented three distinct criteria for analyzing recurrence.
 Firstly, Theorem 3.7 introduces a criterion for ergodic Lazy coins, where the recurrence is determined by the operators $L$ and $R$ and by the only invariant state $\rho_{\infty}$ associated with the corresponding auxiliary map.
-Next, Theorem 3.9 focuses on Lazy OQWs of dimension 2, offering a complete and easy-to-apply criterion.
+Next, Theorem 3.9 focuses on Lazy HOQWs of dimension 2, offering a complete and easy-to-apply criterion.
 The simplicity of this criterion comes from the fact that it depends on the knowledge of the eigenvectors of the operators $L$, $B$, and $R$, allowing a practical analysis of the recurrence of these processes that are not necessarily ergodic.
 Finally, Theorem 4.3 extends the analysis to non-lazy coins, using a decomposition of the vector space where the operators $L$, $B$, and $R$ act.
-By checking recurrence in vector subspaces that contain unique invariant states, this general criterion is extended to any initial density, significantly expanding the scope of recurrence analysis for OQWs in more complex configurations, since it is valid for any finite dimensional coin.
+By checking recurrence in vector subspaces that contain unique invariant states, this general criterion is extended to any initial density, significantly expanding the scope of recurrence analysis for HOQWs in more complex configurations, since it is valid for any finite-dimensional coin.
 […]

### v3

@@
 Abstract
-In this paper, we study the recurrence of homogeneous Open Quantum Walks (HOQWs) induced by finite-dimensional coins $(L,B,R)$.
-We present three distinct recurrence criteria, each adapted to different types of coins.
-The first criterion applies to coins whose auxiliary map has a unique invariant state, resulting in the first recurrence criterion for Lazy HOQWs.
-The second one applies to Lazy HOQWs of dimension 2, where we provide a complete characterization of the recurrence for this low-dimensional case.
-Finally, we present a general criterion for finite-dimensional coins in the non-lazy case $(B=0)$, which generalizes many of the previously known results.
-This new criterion holds for irreducible and reducible HOQWs.
+In this paper, we study the recurrence of Open Quantum Walks induced by finite-dimensional coins on the line ($\mathbb{Z}$) and on the grid ($\mathbb{Z}^{2}$).
+Two versions are considered: discrete-time open quantum walks (OQW) and continuous-time open quantum walks (CTOQW).
+We present three distinct recurrence criteria for OQWs on $\mathbb{Z}$, each adapted to different types of coins.
+The first criterion applies to coins whose auxiliary map has a unique invariant state, resulting in the first recurrence criterion for Lazy OQWs.
+The second one applies to Lazy OQWs of dimension 2, where we provide a complete characterization of the recurrence for this low-dimensional case.
+The third one presents a general criterion for finite-dimensional coins in the non-lazy case, which generalizes many of the previously known results for OQWs on $\mathbb{Z}$.
+Also, we present a general recurrence criterion for OQWs on $\mathbb{Z}^{2}$ via the open quantum jump chain, obtained from a recurrence criterion for CTOQWs on $\mathbb{Z}^{2}$.
 keywords:
@@
 Open Quantum Walks (OQWs) naturally extend classical Markov chains within the framework of open quantum systems.
-Originally proposed by [2], OQWs describe open quantum systems in which the repeated measurement of the particle’s position, together with the space of particle states, gives rise to a Markov chain.
-In this context, we consider homogeneous OQWs (HOQWs) defined on the set of integers, that is, the graph of this type of random walk is the set of vertices $V=\mathbb{Z}$, where 3 linear operators $L,B,R$ play the roles of transitions between the vertices of the quantum walk, which are constant and each jump only occurs from one vertex to its nearest neighbors, or to itself.
-HOQWs on more general lattices are explored and detailed in [7].
-Recurrence is an important probabilistic property of any random walk, often characterized by the expected number of returns to the initial vertex, emerging as a central question of interest in the quantum context, and can be directly connected with the classical recurrence criteria for random walks in $\mathbb{Z}$.
-For example, consider the classical homogeneous random walk on $\mathbb{Z}$ with graph given by Figure 1.
-It is well known that this lazy random walk with $r,l\geq 0,\;b>0,\;l+b+r=1,$ is recurrent if and only if the transition parameters $r$ (rate of jumping to right) and $l$ (rate of jumping to left) are equal.
+Originally proposed by [2], OQWs describe open quantum systems in which repeated measurements of the particle’s position, together with the particle’s state space, give rise to a Markov chain.
+In this context, we first consider homogeneous Open Quantum Walks (HOQWs) defined on the set of integers, that is, the graph of this type of walk has the set of sites $V=\mathbb{Z}$, where three linear operators $L,B,R$ correspond to transitions between sites in the quantum walk, remaining constant, with each step moving only to the nearest neighbors $(R,L)$ or staying at the same site $(B)$.
+Recurrence is an important probabilistic property of any walk, often characterized by the expected number of returns to the initial site, bringing a central concern in the quantum context, and can be directly related to classical recurrence criteria for walks in $\mathbb{Z}$.
+For example, consider the classical homogeneous walk on $\mathbb{Z}$ with graph as shown in Figure 1.
+It is well known that this lazy random walk with $r,l\geq 0,\;b>0,\;l+b+r=1,$ is recurrent if and only if the transition parameters $r$ (rate of jumping to the right) and $l$ (rate of jumping to the left) are equal.
 Also, if $b=0,$ then the walk is non-lazy and it is recurrent if and only if $r=1/2=l$ [21].
@@
 This is equivalent to saying that the average number of returns to that state is infinite, indicating that the chain will revisit the state indefinitely [21].
-With this description, it is natural to seek criteria to determine under what conditions a HOQW is recurrent given $L,B,R.$ This problem has been addressed in the literature with the development of specific criteria, which are essential to understanding the dynamics and long-term behavior of these systems.
-See for instance [8, 13, 12].
+With this description, it is natural to seek criteria to determine under what conditions an HOQW is recurrent, given $L,B,R.$ This problem has been addressed in the literature with the development of specific criteria, which are essential to understanding the dynamics and long-term behavior of these systems.
+See, for instance, [12, 17, 16].
 These works define recurrence in terms of probability of return, where a state is said to be recurrent if the probability of the walk returning to it is equal to 1.
-However, it has been shown that for OQWs, the two definitions of recurrence based on returning with probability 1 is not equivalent to the one based on the average number of returns being infinite [3].
+However, it has been shown that for OQWs, the two definitions of recurrence based on returning with probability 1 are not equivalent to the one based on the average number of returns being infinite [4].
 This property highlights a distinct characteristic of OQWs compared to classical Markov chains.
-Once this difference is understood, we discuss the recurrence criteria for OQWs, adopting the definition proposed in [3].
-Specifically, we analyze the recurrence of OQWs based on the average number of returns.
+Once this difference is understood, we discuss the recurrence criteria for HOQWs, adopting the definition proposed in [4].
+Specifically, we analyze the recurrence of HOQWs based on the average number of returns.
 Furthermore, recurrence criteria concerning the mean number of returns for higher-order HOQWs have been presented in [15].
@@
 A first Central Limit Theorem (CLT) for Open Quantum Walks (OQWs) was established in [1].
-In this work, we rely on more recent developments, including the CLT for Lazy OQWs presented in [10], as well as the general CLT for homogeneous OQWs (HOQWs) on lattices with finite internal degrees of freedom given in [6].
+In this work, we rely on more recent developments, including the CLT for Lazy HOQWs presented in [14], as well as the general CLT for HOQWs on lattices with finite internal degrees of freedom given in [10].
+By making use of the ideas above and generalizing results from the classical version to recurrence, we also obtain a recurrence criterion for the HOQW on the grid, which is the case where the set of vertices is $\mathbb{Z}^{2}$; however, this is achieved through continuous-time OQWs (CTOQWs).
+For this purpose, we use recent results by [20], where the introduction of the open quantum jump chain allowed for the association between recurrence in discrete-time and continuous-time versions of OQWs.
 This work is structured as follows.
-In Section 2 we define the open quantum walks induced by a coin and their transition probabilities.
+In Section 2, we define the open quantum walks induced by a coin and their transition probabilities.
 Section 3 is devoted to presenting a recurrence criterion for a class of Lazy HOQWs, and one restricted to Lazy HOQWs of dimension 2.
-In Section 4 we focus on the case where the coin is non-lazy and we present a general recurrence criterion for finite-dimensional coins.
-Section 5 illustrates the results with comprehensive examples.
+In Section 4, we focus on the case where the coin is non-lazy, and we present a general recurrence criterion for finite-dimensional coins.
+Section 5 illustrates a first pack of comprehensive examples.
+Section 6 discusses OQWs on the grid, and Section 7 brings examples of HOQWs on those walks.
 2 General Settings
@@
 In this article, we consider only the case $d<\infty.$
-If $B\neq 0,$ the quantum walk exhibits a lazy behavior, meaning that in the context of quantum walks there is a non-zero probability for the walker to stay in the same vertex after a step.
+If $B\neq 0,$ the quantum walk exhibits a lazy behavior, meaning that in the context of quantum walks, there is a non-zero probability for the walker to stay in the same vertex after a step.
 We refer to the HOQW above as a HOQW induced by the coin $(L,B,R)$ in which the walk is inherently lazy.
-Therefore, this coin represents a homogeneous, nearest-neighbor OQW on the integer line with the left and right transitions given by $L$ and $R$, respectively, and $B$ represents the rate of no jump at some instant.
+Therefore, this coin represents a homogeneous, nearest-neighbor OQW on the integer line with the left and right transitions given by $L$ and $R$, respectively, while $B$ describes the rate of no jump at some instant.
 This lazy behavior introduces distinct properties compared to its non-lazy counterpart.
 Notably, the presence of the matrix $B$ influences the walk’s propensity to remain in its current vertex, thereby affecting the overall dynamics.
-If $B=0,$ then $\Phi$ will be called a HOQW induced by a coin $(L,R)$ or a non-lazy HOQW induced by a coin $(L,R).$ The dynamic described above is represented by the graph in Figure 2.
+If $B=0,$ then $\Phi$ will be called a HOQW induced by a coin $(L,R),$ or a non-lazy HOQW induced by a coin $(L,R).$ The dynamic described above is represented by the graph in Figure 2.
 […]
 Figure 2: Lazy Quantum Walk on $\mathbb{Z}$.
-The quantum trajectory of this kind of OQW on the line, induced by a coin $(L,B,R)$ and starting from a state $\tau$ of the form $\tau=\sum_{i\in\mathbb{Z}}\tau_{i}\otimes|i\rangle\langle i|,$ is any path generated by the Markov chain $(X_{n},\tau_{n})_{n\geq 0}$, where $X_{n}$ denotes the position of the particle at time $n$ and $\tau_{n}$ its internal degree.
+The quantum trajectory of this kind of HOQW on the line, induced by a coin $(L,B,R)$ and starting from a state $\tau$ of the form $\tau=\sum_{i\in\mathbb{Z}}\tau_{i}\otimes|i\rangle\langle i|,$ is any path generated by the Markov chain $(X_{n},\tau_{n})_{n\geq 0}$, where $X_{n}$ denotes the position of the particle at time $n$ and $\tau_{n}$ its internal degree.
 The transition probabilities are given by
@@
 A pure density is a density $\rho$ that is a projection of rank 1, i.e., there exists some vector $\lvert u\rangle\in\mathbb{C}^{d}$ such that $\rho=\lvert u\rangle\langle u\rvert$.
-Let us consider an OQW starting the walk on vertex $\lvert i\rangle$ with initial density operator $\rho.$ For evolution on time $n=0,1,2,\ldots,$ we define
-$p_{n}(j)=\mathbb{P}(\mbox{the quantum particle, at time }n,\mbox{ is in site}\;\lvert j\rangle)=\mathrm{Tr}(\rho_{n}(j)),$
+Let us consider an HOQW starting the walk on vertex $\lvert i\rangle$ with initial density operator $\rho.$ For evolution on time $n=0,1,2,\ldots,$ we define
+[…]
 where
@@
 With this notation,
-$p_{ji;\rho}(n)=p_{n}(\rho\otimes|i\rangle\to|j\rangle)=\mathrm{\mathrm{Tr}}(\rho_{n}(j)\otimes\lvert j\rangle\langle j\rvert)=\mathrm{\mathrm{Tr}}\left(\Phi^{n}(\rho\otimes\lvert i\rangle\langle i\rvert)(I\otimes\lvert j\rangle\langle j\rvert)\right)$
+[…]
 denote the probability of being at site $\lvert j\rangle$ at step $n,$ given that we started at site $\lvert i\rangle,$ with initial density $\rho$ concentrated at $\lvert i\rangle.$ Therefore, the dynamics starts with a density operator $\rho$ concentrated at some vertex $\lvert i\rangle,$ takes the evolution up to time ($n$) through the linear operator $\Phi^{n},$ producing a new density operator
@@
 which represents the data concentrated at vertex $\lvert j\rangle$ at time $n.$
-Consider an OQW on some set of vertices $\mathbb{Z}$, let $i\in\mathbb{Z}$ and $\rho\in\mathcal{D}(\mathfrak{h})$. We say that a vertex $\lvert i\rangle$ is
+Consider an HOQW on some set of vertices $\mathbb{Z}$, let $i\in\mathbb{Z}$ and $\rho\in\mathcal{D}(\mathfrak{h})$. We say that a vertex $\lvert i\rangle$ is
   1. 1.
@@
   3. 3.
-    transient, if $\lvert i\rangle$ is $\rho-$transient for all $\rho\in\rho\in\mathcal{D}(\mathfrak{h}).$
+    transient, if $\lvert i\rangle$ is $\rho$-transient for all $\rho\in\mathcal{D}(\mathfrak{h}).$
 Concerning all vertices, an OQW is called: recurrent if every vertex is recurrent; transient if every vertex is transient.
@@
 where each $\lambda_{j}>0$ is an eigenvalue of $\tau$ with corresponding eigenvector $\lvert e_{j}\rangle.$ Then
-$\infty=\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\tau\otimes\lvert i\rangle\langle i\rvert\right)\right(I\otimes\lvert i\rangle\langle i\rvert)]=\sum_{j}\left(\lambda_{j}\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\lvert e_{j}\rangle\langle e_{j}\rvert\otimes\lvert i\rangle\langle i\rvert\right)\right(I\otimes\lvert i\rangle\langle i\rvert)]\right),$
+[…]
 where at least one of the members of the sum must be infinity. This completes the proof.
@@
 Corollary 3.3.
-Let $\Phi$ be a HOQW on $\mathbb{Z}$ with $\dim(\mathfrak{h})=d<\infty.$ If $\Phi$ is $\tau$-recurrent for some $\tau\in\mathcal{D}(\mathfrak{h}),$ then
+Let $\Phi$ be an HOQW on $\mathbb{Z}$ with $\dim(\mathfrak{h})=d<\infty.$ If $\Phi$ is $\tau$-recurrent for some $\tau\in\mathcal{D}(\mathfrak{h}),$ then
   1. 1.
@@
 A minimal enclosure with finite dimension is the support of a unique invariant state of the channel, that is, a density operator $\sigma$ such that $\mathcal{L}(\sigma)=\sigma$.
-For more on enclosures of OQWs, we refer the reader to [6, 7, 3].
+For more on enclosures of HOQWs, we refer the reader to [10, 11, 4].
 Following this property and using the homogeneity of HOQWs induced by a coin, we split $\mathfrak{h}$ into two subsets $\mathcal{Y}$ and $\mathcal{X}$ in such a way that the walk is transient for $\mathcal{D}(\mathcal{Y})$ and recurrent everywhere else, with $\mathcal{Y}$ being an enclosure of the quantum channel with Kraus operators $L,B,R.$
+Remark 3.4.
+From now on we disregard the trivial cases where at least one of the operators $L,B,R$ has some eigenvalue $\lambda$ of modulus 1.
+To clarify, let $T\lvert v\rangle=\lambda\lvert v\rangle,$ where $\lvert v\rangle$ is an unitary vector and $T\in\{L,B,R\}.$ In this case, starting the walk with density operator $\rho=\lvert v\rangle\langle v\rvert$, the walker is either absorbed by a vertex $(T=B)$ or the particle always jumps to the same side $(T=L$ or $T=R).$
 Proposition 3.5.
-Consider a HOQW induced by a coin $(L,B,R)$ of dimension $d$. We are in one (and only one) of the following situations:
+Consider an HOQW induced by a coin $(L,B,R)$ of dimension $d$. We are in one (and only one) of the following situations:
   • $(L,B,R)$ is recurrent;
@@
 where $J_{\pi}$ is a concatenation of $n$ operators, each one belonging to $\{L,B,R\},$ proving the claim, once this is valid for any $n=1,2,3,\ldots$
-Moreover, by equation (4) we have that $\cup_{n}\;\text{supp}(\rho_{n})$ is a subspace of $\mathfrak{h}.$ Joining all the possible initial states in which $(L,B,R)$ is transient, we obtain a decomposition (3), where $\mathcal{Y}$ or $\mathcal{X}$ may be null, and by definition $\mathcal{Y}$ is an enclosure.
+Moreover, by equation (4) we have that $\cup_{n}\;\text{supp}(\rho_{n})$ is a subspace of $\mathfrak{h}.$ Joining all the possible initial states in which $(L,B,R)$ is transient, we obtain a decomposition (3), where $\mathcal{Y}$ or $\mathcal{X}$ may be $\{0\}$, and by definition $\mathcal{Y}$ is an enclosure.
 Let us suppose that $(L,B,R)$ is neither recurrent nor transient.
@@
 Proof.
-This result can be obtained by using a Doob decomposition as described in [10, Equation 27] and [1, Theorem 5.2] for a Martingale $M_{n}$ satisfying
+This result can be obtained by using a Doob decomposition as described in [14, Equation 27] and [1, Theorem 5.2] for a Martingale $M_{n}$ satisfying
 $X_{n}-n.m=X_{0}+M_{n}+\beta_{n},$
@@
 Theorem 3.7 (Chung-Fuchs Theorem for HOQWs with ergodic auxiliary map).
-Consider a HOQW induced by a coin $(L,B,R),\;\mathcal{L}$ its auxiliary map and $\mathcal{Y}$ an enclosure.
+Consider an HOQW induced by a coin $(L,B,R),\;\mathcal{L}$ its auxiliary map and $\mathcal{Y}$ an enclosure.
 Suppose that the initial density operator is $\rho\in\mathcal{D}(\mathcal{Y})$ with quantum trajectories $(X_{n},\rho_{n})_{n\geq 0}$ and $X_{n}/n\rightarrow 0$ in probability, then $(L,B,R)$ is $\sigma$-recurrent for some $\sigma\in\mathcal{D}(\mathcal{Y}).$
 Proof.
-The proof is analogous to [11, Theorem 7] and we will proceed by remarking on how it can be generalized.
-Firstly, the maximum appearing in [11, Lemma 6] can be taken in $\mathcal{D}=\mathcal{D}(\mathcal{Y}).$ Indeed, if the initial state is $\rho\in\mathcal{D}(\mathcal{Y}),$ then $\rho_{n}\in\mathcal{D}(\mathcal{Y})$ for every $n,$ thereby the maximum taken in the proof of [11, Lemma 6] can be reduced to $\mathcal{D}(\mathcal{Y}).$ Therefore, by the same reason, the density operator considered in [11, Theorem 7] belongs to $\mathcal{D}(\mathcal{Y}),$ proving the Theorem for the non-lazy case.
-Moreover, the lazy case holds as an extension of this proof, since all the results work for homogeneous OQWs on $\mathbb{Z}$.
+The proof is analogous to [15, Theorem 7] and we will proceed by remarking on how it can be generalized.
+Firstly, the maximum appearing in [15, Lemma 6] can be taken in $\mathcal{D}=\mathcal{D}(\mathcal{Y}).$ Indeed, if the initial state is $\rho\in\mathcal{D}(\mathcal{Y}),$ then $\rho_{n}\in\mathcal{D}(\mathcal{Y})$ for every $n,$ thereby the maximum taken in the proof of [15, Lemma 6] can be reduced to $\mathcal{D}(\mathcal{Y}).$ Therefore, by the same reason, the density operator considered in [15, Theorem 7] belongs to $\mathcal{D}(\mathcal{Y}),$ proving the Theorem for the non-lazy case.
+Moreover, the lazy case holds as an extension of this proof, since all the results work for HOQWs on $\mathbb{Z}$.
 $\square$
 The following theorem presents the first recurrence criterion for non-irreducible HOQWs of arbitrary finite dimension.
-This result extends [11, Corollary 8 and Proposition 8] for reducible CTOQWs and can also be applied to Lazy HOQWs.
+This result extends [15, Corollary 8 and Proposition 8] for reducible CTOQWs and can also be applied to Lazy HOQWs.
 Theorem 3.8 (Recurrence Criteria).
-Let us consider a HOQW on $\mathbb{Z}$ induced by a coin $(L,B,R)$ of dimension $d$ such that the auxiliary map $\mathcal{L}$ is ergodic.
+Let us consider an HOQW on $\mathbb{Z}$ induced by a coin $(L,B,R)$ of dimension $d$ such that the auxiliary map $\mathcal{L}$ is ergodic.
 Then $\mathcal{L}$ has a unique invariant state $\rho_{\infty}$ and
@@
 Moreover, the unique invariant state $\rho_{\infty}$ of $\mathcal{L}$ must belong to $\mathcal{D}(\mathcal{Y}).$
-If $\;\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)=\textmd{Tr}\left(R\rho_{\infty}R^{*}\right),$ then $X_{n}/n\rightarrow 0$ by Theorem 3.5, thus $(L,B,R)$ is $\rho$-recurrent for some $\rho\in\mathcal{D}(\mathcal{Y})$ by Theorem 3.6, which is a contradiction.
-[…]
+If $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)=\textmd{Tr}\left(R\rho_{\infty}R^{*}\right)$ then $X_{n}/n\rightarrow 0$ by Theorem 3.6, thus $(L,B,R)$ is $\rho$-recurrent for some $\rho\in\mathcal{D}(\mathcal{Y})$ by Theorem 3.7, which is a contradiction.
+On the other hand, if $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)\neq\textmd{Tr}\left(R\rho_{\infty}R^{*}\right),$ then $X_{n}/n\rightarrow m\neq 0.$ Therefore $\mathbb{P}_{0,\rho}(X_{n}=0\mbox{ i.o.})=0,$ showing that the mean number of returns to the initial site is finite for any $\rho$, thus the walk is transient, which is also a contradiction.
+We have shown that either the walk is recurrent or transient.
 Therefore, if $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)=\textmd{Tr}\left(R\rho_{\infty}R^{*}\right),$ then the walk is $\rho$-recurrent for some density and thus it is recurrent.
@@
 $\mathcal{L}$ has a minimal invariant subspace, that is, a subspace $S\neq\{0\}$ such that $S\subset S^{\prime}$ for every invariant subspace $S^{\prime}\neq\{0\}.$
-Building upon the established preceding proposition, we will advance to the following theorem, delineating a comprehensive criterion for 2-dimensional coins, extending [11, Theorem 17] to encompass the Lazy case.
+Building upon the established preceding proposition, we will advance to the following theorem, delineating a comprehensive criterion for 2-dimensional coins, extending [15, Theorem 17] to encompass the Lazy case.
 This criterion will be structured into two distinct parts: the first adhering to the logic of the previous theorem, and the second articulated in terms of the eigenvalues of the operators of the coin, independent of the invariant densities of the auxiliary map.
 Theorem 3.10.
-Consider a HOQW on $\mathbb{Z}$ induced by a coin $(L,B,R)$ of dimension 2.
+Consider an HOQW on $\mathbb{Z}$ induced by a coin $(L,B,R)$ of dimension 2.
 (1) If $L,B,R$ have at most one common eigenvector, let $\rho_{\infty}$ be the unique invariant density of the auxiliary map. Then, we have
@@
 and
-$(2.1)\quad|l_{1}|=|r_{1}|\mbox{ and }|l_{2}|=|r_{2}|$$\Rightarrow$ $(L,B,R)\mbox{ is recurrent,}$,
-$(2.2)\quad|l_{1}|\neq|r_{1}|\mbox{ and }|l_{2}|\neq|r_{2}|$$\Rightarrow$ $(L,B,R)\mbox{ is transient,}$,
-$(2.3)\quad|l_{i}|=|r_{i}|\mbox{ and }|l_{j}|\neq|r_{j}|$$\Rightarrow$ $(L,B,R)\mbox{ is transient with respect to }\sigma_{j}=\lvert u_{j}\rangle\langle u_{j}\rvert\mbox{ and it is recurrent}$
+$(2.1)\quad|l_{1}|=|r_{1}|\mbox{ and }|l_{2}|=|r_{2}|$ $\Rightarrow$ $(L,B,R)\mbox{ is recurrent,}$
+$(2.2)\quad|l_{1}|\neq|r_{1}|\mbox{ and }|l_{2}|\neq|r_{2}|$ $\Rightarrow$ $(L,B,R)\mbox{ is transient,}$
+[…]
 […]
@@
 a contradiction. Therefore, we are in the first situation of proposition 3.5 and the walk is recurrent.
-Analogously, if $|l_{1}|\neq|r_{1}|$ and $|l_{2}|\neq|r_{2}|,$ then for $\rho_{1}=\lvert u_{1}\rangle\langle u_{1}\rvert,$ and $\rho_{2}=\lvert u_{2}\rangle\langle u_{2}\rvert,$ where $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ are orthonormal common eigenvectors of $L,B,R,$ we have $p_{j,j+1;\rho}(n)<p_{j,j-1;\rho}(n)$ (or $>$) for all $j\in\mathbb{Z}$ and all $n\geq 0$ and therefore in this case we have a classical homogeneous walk which is transient for both $\rho_{1},\rho_{2}$.
+Analogously, if $|l_{1}|\neq|r_{1}|$ and $|l_{2}|\neq|r_{2}|,$ then for $\rho_{1}=\lvert u_{1}\rangle\langle u_{1}\rvert$, and $\rho_{2}=\lvert u_{2}\rangle\langle u_{2}\rvert$, where $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ are orthonormal common eigenvectors of $L,B,R,$ we have $p_{j,j+1;\rho}(n)<p_{j,j-1;\rho}(n)$ (or >>) for all $j\in\mathbb{Z}$ and all $n\geq 0$ and therefore in this case we have a classical homogeneous walk which is transient for both $\rho_{1},\rho_{2}$.
 Since the walk is transient for two distinct eigenvectors, the walk is transient by corollary 3.2.
-For the last case, $|l_{i}|=|r_{i}|$ and $|l_{j}|\neq|r_{j}|,$ $(L,B,R)$ is $\rho_{i}$-recurrent and $\rho_{j}$-transient, where $\rho_{i}=\lvert u_{i}\rangle\langle u_{i}\rvert$ and $\rho_{j}=\lvert u_{j}\rangle\langle u_{j}\rvert.$ This situation is a particular case of the last item in Proposition 3.4.
+For the last case, $|l_{i}|=|r_{i}|$ and $|l_{j}|\neq|r_{j}|,$ $(L,B,R)$ is $\rho_{i}$-recurrent and $\rho_{j}$-transient, where $\rho_{i}=\lvert u_{i}\rangle\langle u_{i}\rvert$ and $\rho_{j}=\lvert u_{j}\rangle\langle u_{j}\rvert$.
+This situation is a particular case of the last item in Proposition 3.5.
 $\square$
 4 A General Recurrence Criteria for Open Quantum Walks
-Following the approach described in [6, Section 4], as well as the references therein, we can obtain for a coin $(L,R)$ a decomposition
-$\mathfrak{h}=\mathfrak{R}\oplus\mathfrak{T},\qquad\mathfrak{R}=\sup\{\mbox{supp}(\omega)\;|\;\omega\mbox{ is an invariant state for }\mathcal{L}\}.$
+Following the approach described in [10, Section 4], as well as the references therein, we can obtain for a coin $(L,R)$ a decomposition
+[…]
 Since $\mathfrak{h}$ is finite-dimensional, any minimal enclosure is included in $\mathfrak{R},$ and is the support of a unique extremal invariant state.
@@
 […]
-Further, every $\mathfrak{X}_{\alpha}$ either is a minimal enclosure or assumes a decomposition (not unique) as the sum of mutually orthogonal isomorphic minimal enclosures $\{\mathfrak{V}_{\alpha,\beta},\;\beta\in I_{\alpha}\}:$
-$\mathfrak{X}_{\alpha}=\bigoplus_{\beta\in I_{\alpha}}\mathfrak{V}_{\alpha,\beta}.$
-Given a minimal enclosure $\mathfrak{V}_{\beta},$ let $\tau_{\beta}$ be the unique invariant state with support on $\mathfrak{V}_{\beta}.$ Introducing the parameter
+Further, every $\mathfrak{X}_{\alpha}$ either is a minimal enclosure or assumes a decomposition (not unique) as the sum of mutually orthogonal isomorphic minimal enclosures $\{\mathcal{Y}_{\alpha,\beta},\;\beta\in I_{\alpha}\}:$
+$\mathfrak{X}_{\alpha}=\bigoplus_{\beta\in I_{\alpha}}\mathcal{Y}_{\alpha,\beta}.$
+Given a minimal enclosure $\mathcal{Y}_{\beta},$ let $\tau_{\beta}$ be the unique invariant state with support on $\mathcal{Y}_{\beta}.$ Introducing the parameter
 $m_{\beta}=\mathrm{Tr}\left(L\tau_{\beta}L^{*}\right)-\mathrm{Tr}\left(R\tau_{\beta}R^{*}\right),$
-we have by [6, Lemma 4.2] that $m_{\beta}$ is independent of the particular minimal enclosure $\mathfrak{V}_{\beta}$ in $\mathfrak{X}_{\alpha}.$ Therefore, we can combine Proposition 3.4, Theorem 3.7 and the discussion above to state the property
+we have by [10, Lemma 4.2] that $m_{\beta}$ is independent of the particular minimal enclosure $\mathcal{Y}_{\beta}$ in $\mathfrak{X}_{\alpha}.$ Therefore, we can combine Proposition 3.5, Theorem 3.8 and the discussion above to state the property
 Proposition 4.1.
@@
 where $\delta_{m_{\alpha}}$ is the Dirac measure concentrated in $m_{\alpha}.$
-We remark that the convergence in law brought by corollary 4.2 can be upgraded to a convergence a.s., as demonstrated in [9, Theorem 3.5.2].
-Let us start a HOQW induced by a coin $(L,R).$ There are two possibilities concerning recurrence when $\text{supp}(\rho_{0})\not\perp\mathcal{X}:$
-1.
-$(L,R)$ is recurrent for $\mathcal{Y}_{\alpha}$ for some $\alpha$ and supp$\left(\mathcal{L}^{n}(\rho_{0})\right)\not\perp\mathcal{Y}_{\alpha}$ for some $n\in\mathbb{N}$: in this case, the walk is $\rho_{0}$-recurrent.
-Indeed, concerning the quantum trajectories $(\rho_{n})_{n}$, we could refresh the walk on some instant $n$ at some vertex $\lvert X_{n}\rangle,$ which is $\rho_{n}$-recurrent, since supp$\left(\mathcal{L}^{n}(\rho_{0})\right)\not\perp\mathcal{Y}_{\alpha}$ for a sufficiently large $n$.
-If there was some density $\sigma$ with supp$(\sigma)\not\perp\mathcal{X}$ and $\sigma$-transient, there would be some enclosure not orthogonal to $\mathcal{X},$ which is a contradiction by Proposition 3.4.
-The conclusion is held by the homogeneity of the coin, since, in this case, the walk either is $\rho_{s}$-recurrent for every $s=0,1,2,\ldots$ or it is $\rho_{s}$-transient for every $s=0,1,2,\ldots$, analogously to the proof of Proposition 3.4.
-2.
-$(L,R)$ is transient for all $\mathcal{Y}_{\alpha}$ satisfying supp$\left(\mathcal{L}^{n}(\rho_{0})\right)\not\perp\mathcal{Y}_{\alpha}$ for every $n\in\mathbb{N}$: the walk is $\sigma$-transient for all $\sigma\in\mathcal{D}(\mathcal{X})$.
-Indeed, in this case, starting the walk on vertex $X_{0}=0,$ $X_{n}/n$ converges (when $n\rightarrow\infty$) to
+We remark that the convergence in law brought by corollary 4.2 can be upgraded to a convergence a.s., as demonstrated in [13, Theorem 3.5.2].
+Let us start an HOQW induced by a coin $(L,R).$ There are two possibilities concerning recurrence when $\text{supp}(\rho_{0})\not\perp\mathcal{X}:$
+[…]
+[…]
 $m=\sum_{\alpha\in A}a_{\alpha}(\rho)\delta_{m_{\alpha}}=\sum_{\alpha\in A}a_{\alpha}(\rho),$
 as given by Equation (6), where the second equality follows from $m_{\alpha}\neq 0,$ since $(L,R)$ is $\mathcal{Y}_{\alpha}$ transient.
-Thus, $X_{n}$ converges either to $\infty$- ∞ or $\infty$,∞ , in probability, imposing that the walk is $\rho$-transient and thus $(L,R)$ is transient on $\mathcal{D}(\mathcal{X}).$
+Thus, $X_{n}$ converges either to -$\infty$ or $\infty,$ in probability, imposing that the walk is $\rho$-transient and thus $(L,R)$ is transient on $\mathcal{D}(\mathcal{X}).$
 Therefore, we have shown the following recurrence criterion:
@@
 Consider a coin $(L,R)$ of dimension $d.$ Consider the decomposition of $\mathfrak{h}$ as given in Equation (5), and $\tau_{1},\ldots,\tau_{m}$ the respectively invariant states of $\mathcal{L}$ in $\mathcal{Y}_{\alpha},\alpha=1,\ldots,m.$
-$(1)\quad\mathrm{Tr}(L\tau_{j}L^{*})\neq\dfrac{1}{2}\;\forall j\in\{1,\ldots m\}\Rightarrow(L,R)\mbox{ is transient;}$;
-$(2)\quad\mathrm{Tr}(L\tau_{j_{k}}L^{*})=\dfrac{1}{2}\;$ for a set of indices $j_{k}\in J\subset\{1,\ldots,m\}$ and $\mathrm{Tr}(L\tau_{j_{l}}L^{*})\neq\dfrac{1}{2}\;$ for the remaining indices $j_{l}\in J^{C},$ and $\sigma$ is a density such that supp$(\sigma)\not\perp\mathcal{X}.$
-(2.1) supp$\left(\mathcal{L}^{n}(\sigma)\right)\not\perp\mathcal{Y}_{\alpha}$ for some $\alpha\in J$ and some $n\in\mathbb{N},$ then $(L,R)$ is
+$(1)\quad\mathrm{Tr}(L\tau_{j}L^{*})\neq\dfrac{1}{2}\;\forall j\in\{1,\ldots m\}\Rightarrow(L,R)\mbox{ is transient;}$
+[…]
+[…]
 $\begin{cases}\rho\text{-transient for }\rho\in\mathcal{D}_{1}:=\mathcal{D}(\oplus_{k\in J^{C}}\mathcal{Y}_{k})\\ \rho\text{-recurrent for }\rho\in\mathcal{D}_{1}^{C}.\end{cases}$
-(2.2) supp$\left(\mathcal{L}^{n}(\rho)\right)\perp\mathcal{Y}_{\alpha}$ for all $\alpha\in J$ and all $n\in\mathbb{N},$ then $(L,R)$ is
+[…]
 $\begin{cases}\rho\text{-transient for }\rho\in\mathcal{D}_{2}:=\mathcal{D}(\oplus_{k\in J^{C}}\mathcal{Y}_{k}\oplus\mathcal{X})\\ \rho\text{-recurrent for }\mathcal{D}_{2}^{C}\end{cases}$
 $(3)\quad\mathrm{Tr}(L\tau_{\alpha}L^{*})=\dfrac{1}{2}\;$ for every $\alpha\Rightarrow$ the walk is recurrent.
-Remark 4.4.
-It is worth observing that $\mathcal{D}_{2}$ in Theorem 4.3 actually arises from an enclosure due to the orthogonality between supp$\left(\mathcal{L}^{n}(\rho)\right)$ and $\mathcal{Y}_{\alpha}$; otherwise, it would contradict Proposition 3.4.
-5 Examples
+Remark $4.$
+[…]
+5 Examples - 1D
 Example 5.1.
@@
     The only invariant state for the auxiliary map is $\rho_{\infty}=I/2.$ Therefore,
-    $\mathrm{Tr}(L\rho_{\infty}L^{*})=\frac{n^{2}}{2}=\mathrm{Tr}(R\rho_{\infty}R^{*})\Rightarrow\;(L,B,R)\mbox{ is recurrent,}$,
+    $\mathrm{Tr}(L\rho_{\infty}L^{*})=\frac{n^{2}}{2}=\mathrm{Tr}(R\rho_{\infty}R^{*})\Rightarrow\;(L,B,R)\mbox{ is recurrent,}$
     by Item (1) of Theorem 3.10.
@@
     We are in Item (2) of Theorem 3.10, thus we conclude that
-    $(1)\quad|x_{1}|=\frac{1}{\sqrt{3}}\mbox{ and }|x_{2}|=\frac{1}{2}$$\Rightarrow$ $(L,B,R)\mbox{ is recurrent,}$,
-    $(2)\quad|x_{1}|\neq\frac{1}{\sqrt{3}}\mbox{ and }|x_{2}|\neq\frac{1}{2}$$\Rightarrow$ $(L,B,R)\mbox{ is transient,}$,
-    $(3)\quad|x_{1}|=\frac{1}{\sqrt{3}}\mbox{ and }|x_{2}|\neq\frac{1}{2}$$\Rightarrow$ $(L,B,R)\mbox{ is transient with respect to }\sigma=\begin{bmatrix}0&0\\ 0&1\end{bmatrix}\mbox{ and it is}$
-    recurrent with respect to all densities but $\mbox{recurrent with respect to all densities but }\sigma,$
-    $(4)\quad|x_{1}|\neq\frac{1}{\sqrt{3}}\mbox{ and }|x_{2}|=\frac{1}{2}$$\Rightarrow$ $(L,B,R)\mbox{ is transient with respect to }\sigma=\begin{bmatrix}1&0\\ 0&0\end{bmatrix}\mbox{ and it is}$
-    recurrent with respect to all densities but $\mbox{recurrent with respect to all densities but }\sigma.$
+    $(1)\quad|x_{1}|=\frac{1}{\sqrt{3}}\mbox{ and }|x_{2}|=\frac{1}{2}$ $\Rightarrow$ $(L,B,R)\mbox{ is recurrent,}$
+    $(2)\quad|x_{1}|\neq\frac{1}{\sqrt{3}}\mbox{ and }|x_{2}|\neq\frac{1}{2}$ $\Rightarrow$ $(L,B,R)\mbox{ is transient,}$
+[…]
+    $\mbox{recurrent with respect to all densities but }\sigma,$
+[…]
+    $\mbox{recurrent with respect to all densities but }\sigma.$
 $\Diamond$
@@
 $\Diamond$
-We remark that the choices of the coins above were taken in such a way that $L,B^{\prime}$ and $R$ had a common invariant eigenvector, namely $[1,0]^{T},$ which was not an eigenvector of $B.$ This is due to the hard job of, given non-trivial $R$ and $L,$ to find $B^{\prime}$ in such a way that $(L,B^{\prime},R)$ is recurrent.
+We remark that the choices of the coins above were taken in such a way that $L,B^{\prime}$ and $R$ had a common invariant eigenvector, namely $1,0]^{T},$ which was not an eigenvector of $B.$ This is due to the hard job of, given non-trivial $R$ and $L,$ to find $B^{\prime}$ in such a way that $(L,B^{\prime},R)$ is recurrent.
 Attempt were conduct using mathematic software with many $L$’s and $R$’s having simple entries, but even though the entries of $B$’s that made $(L,B,R)$ to be transient were very long to write in this paper.
 Example 5.4.
-Let us consider the coin given in [6, Section 6.2]:
+Let us consider the coin given in [10, Section 6.2]:
 $R=\begin{bmatrix}\sqrt{\frac{3}{8}}&0&0&0\\ -\sqrt{\frac{p_{1}}{2}}&\frac{1}{\sqrt{2}}&0&0\\ -\sqrt{\frac{p_{2}}{2}}&0&\frac{1}{\sqrt{2}}&0\\ \sqrt{\frac{2p_{3}}{3}}&0&0&\frac{1}{\sqrt{3}}\end{bmatrix},\qquad L=\begin{bmatrix}\frac{1}{2\sqrt{2}}&0&0&0\\ \sqrt{\frac{p_{1}}{2}}&\frac{1}{\sqrt{2}}&0&0\\ \sqrt{\frac{p_{2}}{2}}&0&\frac{1}{\sqrt{2}}&0\\ -\sqrt{\frac{p_{3}}{3}}&0&0&\frac{\sqrt{2}}{\sqrt{3}}\end{bmatrix}$,
@@
 Example 5.5.
-Consider a HOQW induced by the coin $(L,R),$ where
+Consider an HOQW induced by the coin $(L,R),$ where
 $L=\begin{bmatrix}\frac{\sqrt{5}}{5}&0&-\frac{\sqrt{5}}{5}\\ 0&\frac{2\sqrt{5}}{5}&\frac{\sqrt{5}}{10}\\ 0&0&\frac{1}{2}\end{bmatrix},\qquad R=\begin{bmatrix}\frac{2\sqrt{5}}{5}&0&\frac{\sqrt{5}}{10}\\ 0&\frac{\sqrt{5}}{5}&-\frac{\sqrt{5}}{5}\\ 0&0&\frac{1}{2}\end{bmatrix}$.
@@
 By item (3) of Theorem 4.3, the walk is recurrent.
-6 Conclusion
-We present new general recurrence criteria for classes of HOQWs, whose dynamics are governed by Kraus operators that are induced by a finite-dimensional coin.
-In the context of these HOQWs, we presented three distinct criteria for analyzing recurrence.
-Firstly, Theorem 3.7 introduces a criterion for ergodic Lazy coins, where the recurrence is determined by the operators $L$ and $R$ and by the only invariant state $\rho_{\infty}$ associated with the corresponding auxiliary map.
-Next, Theorem 3.9 focuses on Lazy HOQWs of dimension 2, offering a complete and easy-to-apply criterion.
-The simplicity of this criterion comes from the fact that it depends on the knowledge of the eigenvectors of the operators $L$, $B$, and $R$, allowing a practical analysis of the recurrence of these processes that are not necessarily ergodic.
-Finally, Theorem 4.3 extends the analysis to non-lazy coins, using a decomposition of the vector space where the operators $L$, $B$, and $R$ act.
-By checking recurrence in vector subspaces that contain unique invariant states, this general criterion is extended to any initial density, significantly expanding the scope of recurrence analysis for HOQWs in more complex configurations, since it is valid for any finite-dimensional coin.
+6 2D Open Quantum Walks
+We start this section with the continuous-time OQW.
+6.1 Continuous-Time Open Quantum Walks
+In order to define the CTOQWs, we recall that an operator semigroup $\mathcal{T}$ on a Hilbert space $\mathcal{K}$ is a family of bounded linear operators $(T_{t})$ acting on $\mathcal{K}$, $t\geq 0,$ such that
+$T_{t}T_{s}=T_{t+s},\;\forall s,t\in\mathbb{R}^{+}\;\mbox{ and }\;T_{0}=I_{\mathcal{K}}.$
+If $t\mapsto T_{t}$ is continuous for the operator norm of $\mathcal{K}$, then $\mathcal{T}$ is said to be uniformly continuous, which is equivalent to the following assertion [7]: There exists a bounded linear operator $L$ on $\mathcal{K}$ such that $T_{t}=e^{tL},\;\forall t\geq 0.$ In this case,
+$L=\lim_{t\rightarrow\infty}\frac{1}{t}(T_{t}-I_{\mathcal{K}})$
+and the operator $L$ is called the generator of $\mathcal{T}.$
 […]
-Acknowledgements This work was supported by Universidade Federal do Pampa (UNIPAMPA).
+Definition 6.1.
+Let $V$ be a finite or countable infinite set and $\mathcal{H}$ be a Hilbert space of the form (2).
+A Continuous-time Open Quantum Walk in $V$ is an uniformly continuous QMS on $\mathcal{I}_{1}(\mathcal{H})$ with Lindblad operator of the form
+$\begin{split}\mathcal{\mathcal{L}}:\mathcal{I}_{1}(\mathcal{H})&\rightarrow\mathcal{I}_{1}(\mathcal{H})\\ \rho&\mapsto-i[\textbf{H},\rho]+\sum_{i,j\in V}\left(S_{i}^{j}\rho S_{i}^{j^{*}}-\frac{1}{2}\{S_{i}^{j*}S_{i}^{j},\rho\}\right),\end{split}$ (7)
+where $A,B]\equiv AB-BA$ and $\{A,B\}\equiv AB+BA.$
+To rearrange Equation 7 We can write $S_{i}^{j}=R_{i}^{j}\otimes\lvert j\rangle\langle i\rvert$ for bounded operators $R_{i}^{j}\in\mathcal{B}(\mathfrak{h}_{i},\mathfrak{h}_{j}),$ the bounded operators H and $S_{i}^{j}$ on $\mathcal{H}$ are of the form $\textbf{H}=\sum_{i\in V}H_{i}\otimes\lvert i\rangle\langle i\rvert$, $H_{i}$ is self-adjoint on $\mathfrak{h}_{i},$ and $\sum_{i,j\in V}S_{i}^{j*}S_{i}^{j}$ converges in the strong sense.
+Letting $\rho=\sum_{i\in V}\rho(i)\otimes\lvert i\rangle\langle i\rvert\in\mathcal{D},$ denoting
+$e^{t\mathcal{L}}(\rho)=\mathcal{T}_{t}(\rho)=\sum_{i\in V}\rho_{t}(i)\otimes\lvert i\rangle\langle i\rvert,\forall t\geq 0,$
+we have
+$\frac{d}{dt}\rho_{t}(i)=-i[H_{i},\rho_{t}(i)]+\sum_{j\in V}\left(R_{j}^{i}\rho_{t}(j)R_{j}^{i^{*}}-\frac{1}{2}\{R_{i}^{j*}R_{i}^{j},\rho_{t}(i)\}\right).$
+Equation (7) can be written as given in [3, Equation 18.7]:
+$\mathcal{L}(\rho)=\sum_{i\in V}\left(G_{i}\rho(i)+\rho(i)G_{i}^{*}+\sum_{j\in V}R_{j}^{i}\rho(j)R_{j}^{i*}\right)\otimes\lvert i\rangle\langle i\rvert,\qquad G_{i}=-iH_{i}-\frac{1}{2}\sum_{j\in V}R_{i}^{j*}R_{i}^{j}.$ (8)
+In this paper we deal with the quantum trajectories of CTOQWs, which are more complex than the trajectories od discrete-time HOQWs.
+Thus, we recall that for a givenfinite or countably infinite set of vertices $V,$ a CTOQW is a stochastic process evolving on a Hilbert space of the form (2), the label $i\in V$ represents the position of the walker and when the walker is located at $i\in V,$ its internal state is encoded in $\mathfrak{h}_{i},$ that is, $\mathfrak{h}_{i}$ describes the internal degrees of freedom of the walker when it is at site $i\in V.$
+Given $i\in V,$ we denote the set of density operators concentrated at vertex $i$ by
+$\mathcal{D}_{i}=\{\rho\in\mathcal{I}(\mathfrak{h}_{i}),\;\rho\geq 0,\;\mathrm{Tr}(\rho)=1\}.$
+Thus, for $\rho\in\mathcal{D}_{i},$ we have $\rho\otimes\lvert i\rangle\langle i\rvert\in\mathcal{D}.$
+Starting the walk on site $\lvert i\rangle$ with initial density operator $\rho\in\mathcal{D}_{i},$ the quantum measurement of the “position” gives rise to a probability distribution $p_{0}$ on $V,$ such that
+$p_{0}(j)=\mathbb{P}(\mbox{the quantum particle is in site}\;\lvert j\rangle)=\mathrm{Tr}(\rho(j)),$
+and for evolution on time $t\geq 0,$
+[…]
+where
+$e^{t\mathcal{L}}(\rho)=\sum_{k\in V}\rho_{t}(i)\otimes\lvert k\rangle\langle k\rvert$.
+6.2 Quantum Trajectories
+Let $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$ be a probability space where independent Poisson point processes $N^{ij},i,j\in V,\;i\neq j$ ($N^{ii}=0$ by convention) on $\mathbb{R}^{2}$ are defined.
+The jump from site $i$ to site $j$ on the graph $V$ will be governed by these Poisson point processes [3]:
+Definition 6.2.
+Consider a CTOQW with generator of the form (8) and an initial density operator $\mu=\sum_{i\in V}\rho(i)\otimes\lvert i\rangle\langle i\rvert\in\mathcal{D}.$ The quantum trajectory describing the indirect measurement of the position of the CTOQW is the Markov chain represented by the density operators $(\mu_{t})_{t\geq 0}$ such that $\mu_{0}=\rho_{0}\otimes\lvert X_{0}\rangle\langle X_{0}\rvert$, where $X_{0}$ and $\rho_{0}$ are random variables with distribution
+$\mathbb{P}\left((X_{0},\rho)=\left(i,\frac{\rho(i)}{\mathrm{Tr}(\rho(i))}\right)\right)=\mathrm{Tr}\left(\rho(i)\right)\mbox{ for all }i\in V,$
+and $\mu_{t}=:\rho_{t}\otimes\lvert X_{t}\rangle\langle X_{t}\rvert$ satisfies the stochastic differential equation
+$\begin{split}\mu_{t}=&\mu_{0}+\int_{0}^{t}M(\mu_{s^{-}})ds\\ +&\sum_{ij}\int_{0}^{t}\int_{\mathbb{R}}\left(\frac{S_{i}^{j}\mu_{s^{-}}S_{i}^{j*}}{\mathrm{Tr}(SS_{i}^{j*}\mu_{s^{-}}S_{i}^{j*})}-\mu_{s^{-}}\right)1_{0<y<\mathrm{Tr}(S_{i}^{j}\mu_{s^{-}}S_{i}^{j*})}N^{ij}(dy,ds)\end{split}$ (10)
+for all $t\geq 0,$ where
+$M(u)=\mathcal{L}(u)-\sum_{ij}\left(S_{i}^{j}\mu S_{i}^{j*}-\mu\mathrm{Tr}(S_{i}^{j}\mu S_{i}^{j*})\right).$
+Hence, for a fixed $\mu=\sum_{i}\rho(i)\otimes\lvert i\rangle\langle i\rvert\in\mathcal{D},$
+$M(\mu)=\sum_{i}\left(G_{i}\rho(i)+\rho(i)G_{i}^{*}-\rho(i)\mathrm{Tr}\left(G_{i}\rho(i)+\rho(i)G_{i}^{*}\right)\right)\otimes\lvert i\rangle\langle i\rvert$.
+For simplicity, suppose $X_{0}=i_{0}$ for some $i_{0}\in V$ and $\rho_{0}\in V(\mathfrak{h}_{i_{0}}).$ For all $t\geq 0,$ consider the solution of Equation (10)
+$\eta_{t}=\rho_{0}+\int_{0}^{t}\left(G_{i_{0}}\eta_{s}+\eta_{s}G_{i_{0}}^{*}-\eta_{s}\mathrm{Tr}\left(G_{i_{0}}\eta_{s}+\eta_{s}G_{i_{0}}^{*}\right)\right)ds,$
+which is a density operator on $\mathfrak{h}_{i_{0}}.$ For $j\neq i_{0},$ define
+$T_{1}^{j}=\mathrm{inf}\{t\geq 0;N^{i_{0},j}\left({u,y|0\leq u\leq t,0\leq y\leq\mathrm{Tr}(R_{i_{0}}^{j}\eta_{u}R_{i_{0}}^{j*})}\right)\geq 1\}.$
+Since the random variables $T_{1}^{j}$ are mutually independent and nonatomic, we can define $T_{1}=\inf_{j\neq i_{0}}\{T_{1}^{j}\}$ once there exists a unique $j\in V$ such that $T_{1}^{j}=T_{1}.$ The random variable $T_{1}$ is said to be the first jump time of the CTOQW conditional on $X_{0}=i_{0}.$
+The first jump time to site $\lvert j\rangle$ is then denoted by $T_{1}^{j}$ and has distribution
+$\mathbb{P}(T_{1}^{j}>\varepsilon)=e^{-\int_{0}^{\varepsilon}\mathrm{Tr}(R_{i_{0}}^{j}\eta_{u}R_{i_{0}}^{j*})du},$
+thus
+$\mathbb{P}(T_{1}\leq\varepsilon)\leq\varepsilon\sum_{j\neq i_{0}}\|R_{i_{0}}^{j*}R_{i_{0}}^{j}\|.$
+The strongly convergence of $\sum_{ij}S_{i}^{j*}S_{i}^{j}$ implies that $\mathbb{P}(T_{1}>0)=1.$ Thereby, on $0,T_{1}],$ we can define the solution $(X_{t},\rho_{t})_{t\geq 0}$ as
+$(X_{t},\rho_{t})$ $=$ $(i_{0},\eta_{t})\mbox{ for }t\in[0,T_{1})\mbox{ and }$
+$(X_{T_{1}},\rho_{T_{1}})$ $=$ $\left(j,\frac{R_{i}^{j}\eta_{T_{1}-}R_{i}^{j*}}{\mathrm{Tr}(R_{i}^{j}\eta_{T_{1}-}R_{i}^{j*})}\right)\;\mbox{ if }T_{1}=T_{1}^{j}.$
+Now we solve
+$\eta_{t}=\rho_{T_{1}}+\int_{0}^{t}\left(G_{j}\eta_{s}+\eta_{s}G_{j}^{*}-\eta_{s}\mathrm{Tr}\left(G_{j}\eta_{s}+\eta_{s}G_{j}^{*}\right)\right)ds$
+and then obtain the second jump time $T_{2}.$ So on we obtain an increasing sequence of jumps $(T_{n})_{n}$ with $\lim_{n\rightarrow\infty}T_{n}=\infty$ almost surely (see section 18.2.3 of [3] for more details).
+This means that the walk does not explode, thus the walker has a finite number of jumps in a finite interval.
+For details concerning explosions of classical Markov chains, see [21, Section 2.2].
+6.3 Recurrence of CTOQWs
+Movin forward, we deal with the following class of 2-dimensional CTOQWs, that is, we will consider the nearest-neighbor quantum random walk on $V=\mathbb{Z}^{2}.$
+Definition 6.3.
+Consider a CTOQW on $V=\mathbb{Z}^{2}$. If there exists $A_{1},A_{2},A_{3},A_{4},H\in\mathcal{B}(\mathbb{C}^{n})$ such that
+$H_{k}=H\;\forall k\in\mathbb{Z}^{2},\quad A_{1}=R_{(i,j)}^{(i+1,j)},\;A_{2}=R_{(i,j)}^{(i,j+1)},\;A_{3}=R_{(i,j)}^{(i-1,j+1)},\;A_{4}=R_{(i,j)}^{(i,j-1)},$
+and $R_{x}^{y}=0_{n}$ for the remaining $x,y\in\mathbb{Z}^{2},$ then the CTOQW is said to be induced by a coin $(A,H)$ of dimension $n.$
+For this kind of CTOQW we can let $\mathfrak{h}_{i}=\mathfrak{h}$ for all $i\in\mathbb{Z}^{2}.$ If $\dim(\mathfrak{h})=n,$ then $A_{1},A_{2},A_{3},A_{4}$ and $H$ can be represented by square matrices of order $n.$
+The auxiliary map of a coin $(A,H)$ is defined as
+$\mathbb{L}(\rho)=G\rho+\rho G^{*}+\sum_{j=1}^{4}A_{j}\rho A_{j}^{*},\qquad G=-iH-\frac{1}{2}\sum_{k=1}^{4}A_{k}^{*}A_{k}.$
+The map $\mathbb{L}$ has at least one stationary state $\rho_{\infty},$ that is, $\mathbb{L}(\rho_{\infty})=0$ [5]. When the stationary state is unique, we define
+$m=\begin{bmatrix}m_{1}\\ m_{2}\end{bmatrix},\quad m_{1}=\mathrm{Tr}\left(A_{1}\rho_{\infty}A_{1}^{*}\right)-\mathrm{Tr}\left(A_{3}\rho_{\infty}A_{3}^{*}\right),\quad m_{2}=\mathrm{Tr}\left(A_{2}\rho_{\infty}A_{2}^{*}\right)-\mathrm{Tr}\left(A_{4}\rho_{\infty}A_{4}^{*}\right).$
+To demonstrate the recurrence criteria of this work for CTOQWs, we recall the following statistical results for those random walks.
+Theorem 6.4 ([6]).
+Consider a coin $(A,H)$, where its auxiliary map has a unique invariant state $\rho_{\infty}.$ Then
+$\frac{X_{t}-tm}{\sqrt{t}}\stackrel{{\scriptstyle t\rightarrow\infty}}{{\longrightarrow}}N(0,\sigma^{2})$ (11)
+for some $\sigma>0.$
+Proposition 6.5 (Law of Large Numbers for CTOQWs ).
+[19, Theorem 15] If the auxiliary map of a coin $(D)$ has a unique stationary state, then
+$\frac{X_{t}}{t}\rightarrow m\quad\mbox{a.s}.$
+Let us recall the classical definition of recurrence of CTOQWs.
+For this, $p_{ji;\rho}(t)$ will denote the probability of being at site $j$ at time $t,$ given that we started at site $i,$ with initial density $\rho$ concentrated at $i$:
+[…]
+Therefore, the walk starts with a density operator $\rho$ concentrated at some vertex $\lvert i\rangle,$ takes the evolution up to time $t>0$ through the quantum Markov semigroup generated by $\mathcal{L},$ producing a new density operator
+$\rho_{t}=\sum_{k}\rho_{t}(k)\otimes\lvert k\rangle\langle k\rvert=e^{t\mathcal{L}}(\rho\otimes\lvert i\rangle\langle i\rvert),\quad\mathrm{Tr}\left(\sum_{k}\rho_{t}(k)\right)=1.$
+Now we project $\rho_{t}$ onto the subspace generated by vertex $\lvert j\rangle$ giving
+$e^{t\mathcal{L}}(\rho\otimes\lvert i\rangle\langle i\rvert)(I\otimes\lvert j\rangle\langle j\rvert)=\rho_{t}(j)\otimes\lvert j\rangle\langle j\rvert$,
+which represents the data concentrated at vertex $\lvert j\rangle$ at time $t.$
+Let $i\in V,\rho\in\mathcal{D}_{i}.$ We say that vertex $i$ is
+  • $\rho$-recurrent if
+    $\int_{0}^{\infty}p_{ii;\rho}(t)dt=\infty.$
+    Otherwise, $i$ is said to be $\rho$-transient
+  • recurrent, if $i$ is $\rho$-recurrent for all $\rho\in\mathcal{D}_{i};$
+  • transient, if $i$ is $\rho$-transient for all $\rho\in\mathcal{D}_{i}.$
+By definition, $i$ is $\rho$-recurrent when the mean number of returns to $i$ is infinite.
+Definition 6.6.
+A CTOQW is said to be:
+  • recurrent if every vertex is recurrent;
+  • transient if every vertex is transient.
+Proposition 6.7.
+A CTOQW induced by a coin is $\rho$-recurrent if and only if it is $\rho_{T_{n}}$-recurrent for any possible jump time $T_{n}.$
+Proof.
+Note that if we start the walk with $(X_{0},\rho_{0})=(i,\rho),$ then at some time $t>0,$ we have $(X_{t},\rho_{t})=(j,\rho_{T}).$ If we refresh the walk at this instant, then the walk is still recurrent with respect to the density $\rho_{t},$ by the homogeneity in space.
+The other hand follows the same idea, assuming by contraposition that the walk is $\rho$-transient.
+$\square$
+Proposition 6.8.
+If $(A,H)$ is $\tau$-recurrent for some density $\tau\in\mathcal{D}$ and $\mathfrak{h}$ has at most one non-trivial invariant subspace for $\Phi:\mathfrak{h}\rightarrow\mathfrak{h}$ given by
+$\Phi(.)=\sum_{k}A_{k}(.)A_{k}^{*},$
+then $(A,H)$ is recurrent.
+Proof.
+Let us start supposing that $(A,H)$ is recurrent with respect to some density operator $\tau$ with support on $\mathcal{V}.$ Let $\rho\in\mathcal{D}$ be another density operator.
+Any $\rho_{T_{n}}$ is of the form
+$\rho_{T_{n}}=\frac{A_{k}\eta A_{k}^{*}}{\mathrm{Tr}\left(A_{k}\eta A_{k}^{*}\right)}\text{ for some }k\in\{1,2,3,4\},$ (12)
+[…]
+$\sigma:=\rho_{T_{n}}-\frac{\tau}{a}\geq 0.$
+Therefore,
+$\frac{1}{\mathrm{Tr}(\sigma)}\int_{0}^{\infty}p_{ii;\rho_{T_{n}}}(t)\;\text{d}t=\int_{0}^{\infty}p_{ii;\frac{\sigma}{\mathrm{Tr}(\sigma)}}(t)\;\text{d}t+\frac{1}{a\mathrm{Tr}(\sigma)}\int_{0}^{\infty}p_{ii;\tau}(t)\;\text{d}t=\infty.$
+By Proposition 6.7, $(A,H)$ is $\rho$-recurrent.
+$\square$
+6.4 Recurrence Criterion
+This section is devoted to exhibit a recurrence criterion for finite dimensional coins $(A,H),$ where the auxiliary map has at most one common invariant subspace.
+The result uses a Central Limit Theorem for CTOQWs and the following Lemma.
+Lemma 6.9.
+Let us consider a homogeneous CTOQW on $\mathbb{Z}^{d}.$ Let $\varepsilon>0$ and $m\geq 1$ be an integer. Then, for every $\rho\in\mathcal{D}$,
+$\int_{0}^{\infty}\mathbb{P}_{0,\rho}(|X_{t}|<m\varepsilon)dt\leq(2m)^{d}\cdot\max_{\sigma\in\mathcal{D}}\int_{0}^{\infty}\mathbb{P}_{0,\sigma}(|X_{t}|<\varepsilon)dt.$
+Proof. Firstly, note that
+$\int_{0}^{\infty}\mathbb{P}_{0,\rho}\left(|X_{t}|<m.\varepsilon\right)dt\leq\int_{0}^{\infty}\sum_{k}\mathbb{P}_{0,\rho}\left(X_{t}\in k.\varepsilon+[0,\varepsilon)^{d}\right)dt,$ (13)
+where the sum over $k$ is on the interval $\{-m,-m+1,\ldots,m-2,m-1\}^{d}$.
+Let
+$T_{k}=\inf\left\{z\geq 0:X_{z}\in k.\varepsilon+\left[0,\varepsilon\right)^{d}\right\}\;\text{ and }\;\zeta_{\rho}^{k,\varepsilon}=\int_{0}^{\infty}\mathbb{P}_{0,\rho}\left(X_{t}\in k.\varepsilon+[0,\varepsilon)^{d}\right)dt.$
+. By the Law of Total Probability, and Fubini’s Theorem, we can rewrite $\zeta_{\rho}^{k,\varepsilon}$ as
+$\begin{split}\zeta_{\rho}^{k,\varepsilon}=&\int_{0}^{\infty}\int_{0}^{t}\mathbb{P}_{0,\rho}\left(X_{t}\in k.\varepsilon+[0,\varepsilon)^{d},\;T_{k}=y\right)dy\;dt\\ =&\int_{0}^{\infty}\int_{0}^{\infty}\mathbb{P}_{0,\rho}(|X_{t}-X_{y}|<\varepsilon,T_{k}=y)dy\;dt\\ =&\int_{0}^{\infty}\int_{y}^{\infty}\mathbb{P}_{0,\rho}(|X_{t}-X_{y}|<\varepsilon,T_{k}=y)dt\;dz\\ =&\int_{0}^{\infty}\mathbb{P}_{0,\rho}(T_{k}=z)\int_{y}^{\infty}\mathbb{P}_{0,\rho}(|X_{t}-X_{y}|<\varepsilon)dt\;dy,\end{split}$ (14)
+[…]
+Denote $\mathbb{P}_{0,\rho}^{k,y}=\mathbb{P}_{0,\rho}\left(T_{k}=y\right).$ An application of Fubini’s Theorem gives
+$\begin{split}\zeta_{\varepsilon,k}^{\rho}=&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}\int_{y}^{\infty}\mathbb{P}_{0,\rho}\left(|X_{t}-X_{y}|<\varepsilon\right)dtdy\\ =&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}\int_{y}^{\infty}\sum_{\sigma\in\mathcal{D},j\in\mathbb{Z}}\mathbb{P}_{0,\rho}\left(|X_{t}-X_{y}|<\varepsilon,(X_{y},\rho_{y})=(j,\sigma)\right)dtdy\\ =&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}\sum_{\sigma\in\mathcal{D},j\in\mathbb{Z}}\int_{y}^{\infty}\mathbb{P}_{0,\rho}\left(|X_{t}-X_{y}|<\varepsilon|(X_{y},\rho_{y})=(j,\sigma)\right)\mathbb{P}_{0,\rho}\left((X_{y},\rho_{y})=(j,\sigma)\right)dtdy\\ \zeta_{\varepsilon,k}^{\rho}=&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}\sum_{\sigma\in\mathcal{D},j\in\mathbb{Z}}\int_{y}^{\infty}\mathbb{P}_{j,\sigma}\left(|X_{t-y}-X_{0}|<\varepsilon\right)dt\mathbb{P}_{0,\rho}\left((X_{y},\rho_{y})=(j,\sigma)\right)dy\\ =&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}\sum_{\sigma\in\mathcal{D},j\in\mathbb{Z}}\int_{0}^{\infty}\mathbb{P}_{j,\sigma}\left(|X_{s}-j|<\varepsilon\right)ds\mathbb{P}_{0,\rho}\left((X_{y},\rho_{y})=(j,\sigma)\right)dy\\ =&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}\sum_{\sigma\in\mathcal{D}}\int_{0}^{\infty}\mathbb{P}_{0,\sigma}\left(|X_{s}|<\varepsilon\right)ds\sum_{j\in\mathbb{Z}}\mathbb{P}_{0,\rho}\left((X_{y},\rho_{y})=(j,\sigma)\right)dy\\ \leq&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}\max_{\tau\in\mathcal{D}}\int_{0}^{\infty}\mathbb{P}_{0,\tau}\left(|X_{s}|<\varepsilon\right)ds\sum_{\sigma\in\mathcal{D},j\in\mathbb{Z}}\mathbb{P}_{0,\rho}\left((X_{y},\rho_{y})=(j,\sigma)\right)dy\\ =&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}dy\max_{\tau\in\mathcal{D}}\int_{0}^{\infty}\mathbb{P}_{0,\tau}\left(|X_{s}|<\varepsilon\right)ds\\ =&\max_{\tau\in\mathcal{D}}\int_{0}^{\infty}\mathbb{P}_{0,\tau}\left(|X_{s}|<\varepsilon\right)ds.\end{split}$
+Therefore, Equation (13) and the inequality above give
+$\int_{0}^{\infty}\mathbb{P}_{0,\rho}(|X_{t}|<m\varepsilon)dt\leq\sum_{k}\zeta_{\rho}^{k,\varepsilon}\leq\sum_{k}\max_{\tau\in\mathcal{D}}\int_{0}^{\infty}\mathbb{P}_{0,\tau}(|X_{s}|<\varepsilon)ds=(2m)^{d}\cdot\max_{\tau\in\mathcal{D}}\int_{0}^{\infty}\mathbb{P}_{0,\tau}(|X_{s}|<\varepsilon)ds.$
+Theorem 6.10.
+If $X_{t}$ denotes a CTOQW on $\mathbb{Z}^{2}$ and $X_{t}/t^{1/2}\Rightarrow$ a Gaussian measure, then the CTOQW is recurrent for some density.
+Proof.
+Let us denote $u(t,m)=\mathbb{P}(|X_{t}|<m).$ By Lemma 6.9,
+$\max_{\tau\in D}\int_{0}^{\infty}\mathbb{P}(|X_{t}|<1)dt=\max_{\tau\in D}\int_{0}^{\infty}u(t,1)dt\geq(4m^{2})^{-1}\int_{0}^{\infty}u(t,m)dt.$
+Assume that $m/\sqrt{t}\rightarrow c,$ then
+$u(t,m)\rightarrow\int_{[-c,c]^{2}}f(x)dx,$ (15)
+where $f(x)$ represents the density of the limiting normal distribution. We denote the right-hand side of Equation (15) by $g(c)$ and let $t=\theta m^{2}$ to obtain
+$u(\theta m^{2},m)\rightarrow g(\theta^{-1/2}).$
+Now we use the chenge of variables
+$m^{-2}\int_{0}^{\infty}u(t,m)dt=\int_{0}^{\infty}u(\theta m^{2},m)d\theta,$
+let $m\rightarrow\infty,$ and use Fatou’s lemma to get
+$\liminf_{m\rightarrow\infty}(4m^{2})^{-1}\int_{0}^{\infty}u(t,m)dt\geq 4^{-1}\int_{0}^{\infty}g(\theta^{-1/2})d\theta.$ (16)
+Since the normal density is positive and continuous at $0,$
+$g(c)=\int_{[-c,c]^{2}}n(x)dx\sim n(0)(2c)^{2}$
+as $c\rightarrow 0.$ So $g(\theta^{-1/2})\sim 4n(0)/\theta$ as $\theta\rightarrow\infty,$ the integral in Equation (16) diverges.
+As a result, $\max_{\tau\in D}\int_{0}^{\infty}\mathbb{P}(|X_{t}|<1)dt=\infty,$ leading us to conclude that the walk is $\tau$-recurrent.
+$\square$
+Note that the maximum in Lemma 6.9 can be assumed to be taken over the jump times $(T_{n})_{n}.$ Therefore, by Proposition 6.7 and Theorem 6.10 we obtain the following result.
+Theorem 6.11.
+If $X_{t}$ denotes a CTOQW induced by a coin $(A,H),$ there is at most one minimal enclosure for $\Phi$ and $X_{t}/t^{1/2}\Rightarrow$ a Gaussian measure, then the CTOQW is recurrent.
+We already have a sufficient condition to recurrence given by the Theorem 6.11. For transience, we also have a sufficient condition:
+Theorem 6.12.
+Let us consider a CTOQW induced by a coin $(A,H),$ having at most one minimal enclosure, and let $\rho_{\infty}$ be the unique stationary state of the auxiliary map.
+If $m\neq 0,$ then the CTOQW is transient.
+Proof.
+Let us start the CTOQW at some vertex $\lvert i\rangle\in\mathbb{Z}^{d}$ with initial density operator $\rho.$ By the Law of Large Numbers, Proposition 6.5,
+$\frac{X_{t}}{t}\rightarrow m\neq 0^{d}.$
+This means that at least one coordinate of $X_{t}$ is unbounded, from where we conclude that
+$\mathbb{P}\left(X_{t}=X_{0}\;\text{i.o.}\right)=0.$
+Hence, the CTOQW is transient.
+$\square$
+Finally, we can present a recurrence criteria for CTOQWs on $\mathbb{Z}^{2}$ induced by a coin $(A,H).$
+Theorem 6.13 (Recurrence criteria for CTOQWs on $\mathbb{Z}^{2}$).
+Let us consider a CTOQW induced by a coin $(A,H),$ having at most one minimal enclosure, and let $\rho_{\infty}$ be the unique stationary state of the auxiliary map. Then
+  • $m=0\;\Rightarrow\;$ the CTOQW is recurrent;
+  • $m\neq 0\;\Rightarrow\;$ the CTOQW is transient.
+Proof.
+For $m\neq 0,$ the result is given in Theorem 6.12. It remains to proof that $m=0$ implies that the CTOQW is recurrent. Firstly, if $m=0,$ then by Theorem 6.4
+$\frac{X_{t}}{\sqrt{t}}=\frac{X_{t}-t.m}{\sqrt{t}}\Rightarrow\text{ a non-degenerate normal distribution,}$
+thus we use Theorem 6.10 to obtain a density operator $\tau$ such that the CTOQW is $\tau$-recurrent. Therefore, $(A,H)$ is recurrent by Theorem 6.11.
+$\square$
+6.5 Recurrence criteria for HOQWs
+6.5.1 Criterion for auxiliary map having at most one minimal enclosure
+To present a complete criterion for the recurrence of finite-dimensional coins $(A)$, we will start this section with an easily verifiable criterion based on the quantum jump chains associated with CTOQWs, as presented in [20].
+The complete recurrence criterion will be developed from a decomposition of the Hilbert space $\mathfrak{h}$ into subspaces that support invariant states.
+The central idea is identifying and characterizing these subspaces, thus we will be able to determine the recurrence of quantum walks by analyzing the evolution within each invariant component.
+Furthermore, these invariant states will play an indispensable role in formulating the criterion.
+Let $\Pi$ denote a discrete HOQW given by
+$\Pi(\rho)=\sum_{i,j\in V}M_{j}^{i}\rho(j)M_{j}^{i*}\otimes\lvert i\rangle\langle i\rvert$.
+This HOQW is the open quantum jump chain of the CTOQW with Lindblad generator
+$\mathbb{L}=\Psi-I,\;\;\Psi(\rho)=\sum_{i,j\in V}B_{j}^{i}\rho(j)B_{j}^{i*},\;B_{j}^{i}=M_{j}^{i}\otimes\lvert i\rangle\langle j\rvert$, (17)
+where $I$ is the identity operator acting on $\mathcal{H}$ [20].
+Theorem 6.14.
+[20, Theorem 5.6] A vertex $i\in V$ is $\rho$-recurrent for the HOQW $\Pi$ if and only if it is $\rho$-recurrent for the CTOQW generated by $\mathbb{L}$ given by Equation (17).
+We consider an HOQW with a set of vertices $V=\mathbb{Z}^{2}.$ Analogously to the continuous-time version, if there exists $D_{1},D_{2},D_{3},D_{4}\in\mathcal{B}(\mathbb{C}^{n})$ such that
+$D_{1}=M_{(i,j)}^{(i+1,j)},\;D_{2}=M_{(i,j)}^{(i,j+1)},\;D_{3}=M_{(i,j)}^{(i-1,j+1)},\;D_{4}=M_{(i,j)}^{(i,j-1)},$
+and $M_{x}^{y}=0_{n}$ for the remaining $x,y\in\mathbb{Z}^{2},$ then the HOQW $\Pi$ is said to be induced by a coin $(D)$ of dimension $n.$
+The HOQW induced by a coin $(D)$ is the homogeneous, nearest-neighbor HOQW on the grid, with the right, up, left and down transitions given by $D_{1},D_{2},D_{3},D_{4},$ respectively.
+Therefore, the quantum trajectory of those walks, starting from a state $\tau$ of the form $\tau=\sum_{i\in\mathbb{Z}}\rho(i)\otimes|i\rangle\langle i|,$ is any path generated by the Markov chain $(X_{n},\tau_{n})_{n\geq 0}$, where $X_{n}$ denotes the position of the particle at time $n$ and $\tau_{n}$ its internal degree.
+The transition probabilities are given by
+$\begin{split}\mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left((i+1,j),\frac{D_{1}\sigma D_{1}^{*}}{\mathrm{Tr}(D_{1}\sigma D_{1}^{*})}\right)\;\Bigg|\;(X_{n},\rho_{n})=((i,j),\sigma)\,\right)&=\mathrm{Tr}(D_{1}\sigma D_{1}^{*}),\\ \mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left((i,j+1),\frac{D_{2}\sigma D_{2}^{*}}{\mathrm{Tr}(D_{2}\sigma D_{2}^{*})}\right)\;\Bigg|\;(X_{n},\rho_{n})=((i,j),\sigma)\,\right)&=\mathrm{Tr}(D_{2}\sigma D_{2}^{*}),\\ \mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left((i-1,j),\frac{D_{3}\sigma D_{3}^{*}}{\mathrm{Tr}(D_{3}\sigma D_{3}^{*})}\right)\;\Bigg|\;(X_{n},\rho_{n})=((i,j),\sigma)\,\right)&=\mathrm{Tr}(D_{3}\sigma D_{3}^{*}),\\ \mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left((i,j-1),\frac{D_{4}\sigma D_{4}^{*}}{\mathrm{Tr}(D_{4}\sigma D_{4}^{*})}\right)\;\Bigg|\;(X_{n},\rho_{n})=((i,j),\sigma)\,\right)&=\mathrm{Tr}(D_{4}\sigma D_{4}^{*}),\end{split}$
+for every $(i,j)\in\mathbb{Z}^{2}$, $\sigma\in\mathcal{D}$, and initial law
+$\mathbb{P}\left(\,(X_{0},\rho_{0})=\left((i,j),\frac{\rho(i,j)}{\mathrm{Tr}\,\rho(i,j)}\right)\;\right)=\mathrm{Tr}(\rho(i,j)),$
+and all other transition probabilities are null. The graph of this kind of HOQW is represented in Figure 3.
+$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}\vdots\vdots\vdots\vdots\ldots\vdots\vdots\vdots\vdots$
+Figure 3: Homogeneous Open Quantum walk on $\mathbb{Z}^{2}$.
+Now we can give a recurrence criteria for HOQWs on $\mathbb{Z}^{2}.$ To this, note that if $\rho$ is an invariant state for the auxiliary map $\Phi$ of $(D),$ defined as
+$\Phi(.)=\sum_{j=1}^{4}D_{j}(.)D_{j}^{*},$
+then it is a stationary state for its open quantum jump chain $\mathbb{L}.$ Consequently, if $\Phi$ has at most one minimal enclosure, then $\Phi$ has a unique invariant state $\rho_{\infty}$ and we can let
+$m=\begin{bmatrix}m_{1}\\ m_{2}\end{bmatrix},\quad m_{1}=\mathrm{Tr}\left(D_{1}\rho_{\infty}D_{1}^{*}\right)-\mathrm{Tr}\left(D_{3}\rho_{\infty}D_{3}^{*}\right),\quad m_{2}=\mathrm{Tr}\left(D_{2}\rho_{\infty}D_{2}^{*}\right)-\mathrm{Tr}\left(D_{4}\rho_{\infty}D_{4}^{*}\right)$
+to obtain, as a consequence of Theorems 6.13 and 6.14, the following criterion.
+Theorem 6.15 (Recurrence criterion for HOQWs on $\mathbb{Z}^{2}$).
+Let us consider an HOQW induced by a coin $(D),$ where the map
+$\Phi(.)=\sum_{j=1}^{4}D_{j}(.)D_{j}^{*}$
+has at most one minimal enclosure, and let $\rho_{\infty}$ be the unique invariant state of $\Phi.$ Then
+  • $m=0\;\Rightarrow\;$ the HOQW is recurrent;
+  • $m\neq 0\;\Rightarrow\;$ the HOQW is transient.
+6.5.2 A generalized recurrence criteria for HOQWs on $\mathbb{Z}^{2}$
+We have already presented several recurrence criteria for discrete-time and continuous-time Open Quantum Walks on the set of vertices $\mathbb{Z}^{2}$, constrained by a minimal enclosure.
+These criteria are interesting tools for understanding the long-term behavior of open quantum walks.
+Building on these consequences, we aim to derive a stronger recurrence criterion for discrete-time HOQWs, allowing for a more comprehensive analysis of their statistics.
+At this point we will allow the auxiliary map of a coin $(D)$ to have more than one invariant state, thus we consider the decomposition of $\mathfrak{h}$ given on Section 4.
+Therefore, given a minimal enclosure $\mathcal{Y}_{\alpha},$ let $\tau_{\alpha}$ be the unique invariant state with support on $\mathcal{Y}_{\alpha}.$ Introducing the parameter
+$\tilde{m}_{\alpha}=\begin{bmatrix}m_{1,\alpha}\\ m_{2,\alpha}\end{bmatrix},\quad m_{1,\alpha}=\mathrm{Tr}\left(D_{1}\tau_{\alpha}D_{1}^{*}\right)-\mathrm{Tr}\left(D_{3}\tau_{\alpha}D_{3}^{*}\right),\quad m_{2,\alpha}=\mathrm{Tr}\left(D_{2}\tau_{\alpha}D_{2}^{*}\right)-\mathrm{Tr}\left(D_{4}\tau_{\alpha}D_{4}^{*}\right),$
+we have by [10, Lema 4.2] that $\tilde{m}_{\alpha}$ is independent of the particular minimal enclosure $\mathcal{Y}_{\alpha}$ in $\mathfrak{X}_{\alpha}.$
+Therefore, we can combine Theorem 6.15 and results concerning CTOQWs above to state the next property, which is a continuous-time version on the grid of Proposition 4.1.
+Proposition 6.16.
+Consider a coin $(D).$ We have a decomposition
+$\mathfrak{h}=\left(\bigoplus_{\alpha\in A}\mathcal{Y}_{\alpha}\right)\oplus\mathcal{X},$ (18)
+where $A$ is a finite set, $\mathcal{Y}_{\alpha}$ is a subspace of $\mathfrak{h}$ being the support of exactly one invariant state $\tau_{\alpha}$ for $\Phi,$ $\mathcal{X}$ contains no enclosure, and
+  • $\tilde{m}_{\alpha}=0\Rightarrow$ the walk is $\tau_{\alpha}$-recurrent for all $\tau\in\mathcal{S}(\mathcal{Y}_{\alpha})$;
+  • $\tilde{m}_{\alpha}\neq 0\Rightarrow$ the walk is $\tau_{\alpha}$-transient for all $\tau\in\mathcal{S}(\mathcal{Y}_{\alpha}).$
+Proof.
+Fixed a subspace $\mathcal{Y}_{\alpha}$ which is the support of exactly one invariant state $\tau_{\alpha}\in\mathcal{D},$ we assume that $\tilde{m}_{\alpha}=0.$ Then $(D)$ is $\rho$-recurrent for every $\rho$ by Theorem 6.11.
+If $\tilde{m}_{\alpha}\neq 0,$ then $(D)$ is transient by Theorem 6.12.
+$\square$
+With the decomposition on Equation (18) established, we can say that $\mathcal{Y}_{\alpha}$ is either recurrent or transient.
+It remains to understand what happens when a quantum state has support outside of $\oplus_{\alpha}\mathcal{Y}_{\alpha}.$
+Analogously to the one dimensional case, our recurrence criterion is:
+Theorem 6.17 (A Generalized Criterion for coins $(D)$).
+Consider a coin $(D)$ of dimension $d<\infty.$ Consider the decomposition of $\mathfrak{h}$ as given in Equation (18), and $\tau_{1},\ldots,\tau_{w}$ the respectively invariant states of $\Phi$ in $\mathcal{Y}_{\alpha},\;\alpha=1,\ldots,w.$
+$(1)\quad\tilde{m}_{\alpha}\neq 0\;\forall j\in\{1,\ldots w\}\Rightarrow(D)\mbox{ is transient;}$
+[…]
+[…]
+$\begin{cases}\rho\text{-transient for }\rho\in\mathcal{D}_{1}:=\mathcal{D}(\oplus_{k\in J^{C}}\mathcal{Y}_{k})\\ \rho\text{-recurrent for }\rho\in\mathcal{D}_{1}^{C}.\end{cases}$
+[…]
+$\begin{cases}\rho\text{-transient for }\rho\in\mathcal{D}_{2}:=\mathcal{D}(\oplus_{k\in J^{C}}\mathcal{Y}_{k}\oplus\mathcal{X})\\ \rho\text{-recurrent for }\mathcal{D}_{2}^{C}\end{cases}$
+$(3)\quad\tilde{m}_{\alpha}=0\;$ for every $\alpha\Rightarrow$ the walk is recurrent.
+7 Examples
+Example 7.1.
+We start with a CTOQW induced by a coin $(A,H),$
+$A_{1}=\begin{bmatrix}3&-1\\ 0&0\end{bmatrix},\quad A_{2}=\begin{bmatrix}1&-2\\ 2i&0\end{bmatrix},\quad A_{3}=\begin{bmatrix}1&1\\ -2&2\end{bmatrix},\quad A_{4}=\begin{bmatrix}-2i&i\\ 0&2\end{bmatrix},\quad H=\begin{bmatrix}-1&h\\ \overline{h}&2\end{bmatrix},\;h\in\mathbb{C}.$
+In this case,
+$m=\frac{2\;\text{Re}(h)-10\;\text{Im}(h)+19}{\alpha}\;\begin{bmatrix}4\\ 1\end{bmatrix}$,
+where $\alpha$ depends on $h,$ but it is never null. Therefore, by Theorem 6.13, $(A,H)$ is recurrent if and only if
+$\text{Re}(h)=5\;\text{Im}(h)-\frac{19}{2}.$
+$\Diamond$
+Example 7.2.
+Consider an HOQW induced by a coin $(D),$ where
+$D_{1}=\frac{1}{2\sqrt{30}}\begin{bmatrix}2&0&-2\\ 0&4&1\\ 0&0&\sqrt{5}\end{bmatrix},\;D_{2}=\frac{1}{2\sqrt{6}}\begin{bmatrix}2&0&-2\\ 0&4&1\\ 0&0&\sqrt{5}\end{bmatrix},\;D_{3}=\frac{1}{2\sqrt{30}}\begin{bmatrix}4&0&1\\ 0&2&-2\\ 0&0&\sqrt{5}\end{bmatrix},\;D_{4}=\frac{1}{2\sqrt{6}}\begin{bmatrix}4&0&1\\ 0&2&-2\\ 0&0&\sqrt{5}\end{bmatrix}$.
+The invariant states of the auxiliary map are of the form $\rho_{\infty}=\omega_{1}\sigma_{1}+\omega_{2}\sigma_{2},$ where
+$\sigma_{1}=\begin{bmatrix}1&0&0\\ 0&0&0\\ 0&0&0\end{bmatrix},\quad\sigma_{2}=\begin{bmatrix}0&0&0\\ 0&1&0\\ 0&0&0\end{bmatrix},\quad\omega_{i},\omega_{j}\geq 0,\;\;\omega_{1}+\omega_{2}=1.$
+We have
+$\tilde{m}_{1}=\frac{1}{10}\begin{bmatrix}-1\\ -5\end{bmatrix},\qquad\tilde{m}_{2}=\frac{1}{10}\begin{bmatrix}1\\ 5\end{bmatrix}$.
+By item (1) of Theorem 6.17, the walk is transient.
+Example 7.3.
+Consider the coins $(A,H_{1})$ and $(A,H_{2}),$ where $A_{k}=D_{k}$ for $k=1,2,3,4$ as in Example 7.2, and
+$H_{1}=\begin{bmatrix}0&0&0\\ 0&0&0\\ 0&0&0\end{bmatrix},\qquad H_{2}=\begin{bmatrix}0&1&0\\ 1&0&0\\ 0&0&0\end{bmatrix}$.
+Its straightforward that $(A,H_{1})$ is transient, since we cain join the results of Example 7.2 and Theorem 6.14. However, it is very interesting to see that $(A,H_{2})$ is recurrent. Indeed,
+$\rho_{\infty}=\frac{1}{2}\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&0\end{bmatrix}$
+is the only stationary state and $m=0,$ thus the coin is recurrent by Theorem 6.13.
+$\Diamond$
+[…]
+Acknowledgements This work was supported by Universidade Federal do Rio Grande do Sul (UFRGS).
 […]

### v4

@@
-Recurrence Criteria for Reducible Homogeneous Open Quantum Walks on the Line
+Recurrence Criteria for Reducible Homogeneous Open Quantum Walks on the Line and on the Grid
 Abstract
@@
 The first criterion applies to coins whose auxiliary map has a unique invariant state, resulting in the first recurrence criterion for Lazy OQWs.
-The second one applies to Lazy OQWs of dimension 2, where we provide a complete characterization of the recurrence for this low-dimensional case.
-The third one presents a general criterion for finite-dimensional coins in the non-lazy case, which generalizes many of the previously known results for OQWs on $\mathbb{Z}$.
+The second applies to Lazy OQWs of dimension 2, where we provide a complete characterization of the recurrence for this low-dimensional case.
+The third one presents a general criterion for finite-dimensional coins in the non-lazy case, which generalizes several of the previously known results for OQWs on $\mathbb{Z}$.
 Also, we present a general recurrence criterion for OQWs on $\mathbb{Z}^{2}$ via the open quantum jump chain, obtained from a recurrence criterion for CTOQWs on $\mathbb{Z}^{2}$.
@@
 Originally proposed by [2], OQWs describe open quantum systems in which repeated measurements of the particle’s position, together with the particle’s state space, give rise to a Markov chain.
-In this context, we first consider homogeneous Open Quantum Walks (HOQWs) defined on the set of integers, that is, the graph of this type of walk has the set of sites $V=\mathbb{Z}$, where three linear operators $L,B,R$ correspond to transitions between sites in the quantum walk, remaining constant, with each step moving only to the nearest neighbors $(R,L)$ or staying at the same site $(B)$.
-Recurrence is an important probabilistic property of any walk, often characterized by the expected number of returns to the initial site, bringing a central concern in the quantum context, and can be directly related to classical recurrence criteria for walks in $\mathbb{Z}$.
+OQWs have proved to be a versatile framework for describing discrete-time quantum dynamics driven by environmental interaction.
+Beyond their structural simplicity, they provide an effective language for dissipative implementations of logical operations and state engineering.
+In particular, it has been shown that single-qubit gates and the CNOT gate can be realized as OQWs on fully connected graphs, and that arbitrary single-qubit states as well as Bell states can be prepared through dissipative mechanisms [3].
+From a probabilistic viewpoint, central limit theorems (CLTs) have been established for microscopically derived Lazy OQWs [16], as well as for the CTOQWs [8].
+At the same time, OQWs admit a rich potential-theoretic and asymptotic analysis.
+Fundamental quantities such as passage times, expected number of visits, recurrence and transience have been investigated [4].
+Concerning the implementation of dissipative quantum computing protocols based on the OQW formalism, the approach [26] enables faster convergence to the desired steady state and increases the probability of detecting the computational outcome compared to the canonical approach.
+More recently a quantum PageRank algorithm based on discrete-time OQWs with Weyl operators as Kraus operators was introduced, demonstrating faster convergence properties than several existing proposals and extending ranking methods to complex networks [14].
+Recurrence properties of CTOQWs were investigated in [22], where connections with discrete-time OQWs were established and recurrence criteria were derived.
+Complementarily, [21] analyzed CTOQWs in one dimension through a matrix framework, characterizing site recurrence and related statistics via matrix-valued orthogonal polynomials and Lindblad generators modeling quantum birth-death processes.
+Finally, [24] examined hitting-time statistics under quantum channels, providing practical tools to compute first-visit expectations and variances through monitored quantum evolutions and induced absorbing quantum Markov chains.
+In this work, we first consider homogeneous Open Quantum Walks (HOQWs) defined on the set of integers, that is, the graph of this type of walk has the set of sites $V=\mathbb{Z}$, where three linear operators $L,B,R$ correspond to transitions between sites in the quantum walk, remaining constant, with each step moving only to the nearest neighbors $(R,L)$ or staying at the same site $(B)$.
+Recurrence is an important probabilistic property of any walk, often characterized by the expected number of returns to the initial site.
+This is a central concern in the quantum context, and can be directly related to classical recurrence criteria for walks on $\mathbb{Z}$.
 For example, consider the classical homogeneous walk on $\mathbb{Z}$ with graph as shown in Figure 1.
+[…]
+Figure 1: Classical Homogeneous Lazy Random Walk on $\mathbb{Z}$.
 It is well known that this lazy random walk with $r,l\geq 0,\;b>0,\;l+b+r=1,$ is recurrent if and only if the transition parameters $r$ (rate of jumping to the right) and $l$ (rate of jumping to the left) are equal.
 Also, if $b=0,$ then the walk is non-lazy and it is recurrent if and only if $r=1/2=l$ [25].
-[…]
-Figure 1: Classical Homogeneous Lazy Random Walk on $\mathbb{Z}$.
 In the context of discrete-time Markov chains, a state is called recurrent if, once the chain enters that state, the probability of returning to it is 1.
 This is equivalent to saying that the average number of returns to that state is infinite, indicating that the chain will revisit the state indefinitely [25].
-With this description, it is natural to seek criteria to determine under what conditions an HOQW is recurrent, given $L,B,R.$ This problem has been addressed in the literature with the development of specific criteria, which are essential to understanding the dynamics and long-term behavior of these systems.
+The definition of recurrence employed on this work follows the terminology of [4].
+As shown therein, this notion is better suited to the quantum setting than the classical “probability of return equals 1” criterion, since for OQWs the two are not equivalent.
+Moreover, for irreducible semifinite OQWs, the average number of returns satisfies a universality property: it is either finite for all vertices and all initial internal states, or infinite for all vertices and all initial internal states.
+The same does not hold for the probability of return, which may equal 1 for some internal states and be strictly less than 1 for others, even when the walk starts at the same site.
+The mean return definition also connects naturally to the limit theorems used throughout this work and allows recurrence to be characterized in terms of a drift parameter $m$.
+With this description, it is natural to seek criteria to determine under what conditions an HOQW is recurrent.
+This problem has been addressed in the literature with the development of specific criteria, which are essential to understanding the dynamics and long-term behavior of these systems.
 See, for instance, [13, 18, 19].
-These works define recurrence in terms of probability of return, where a state is said to be recurrent if the probability of the walk returning to it is equal to 1.
-However, it has been shown that for OQWs, the two definitions of recurrence based on returning with probability 1 are not equivalent to the one based on the average number of returns being infinite [4].
-This property highlights a distinct characteristic of OQWs compared to classical Markov chains.
-Once this difference is understood, we discuss the recurrence criteria for HOQWs, adopting the definition proposed in [4].
-Specifically, we analyze the recurrence of HOQWs based on the average number of returns.
-Furthermore, recurrence criteria concerning the mean number of returns for higher-order HOQWs have been presented in [15].
-The authors of that work extend the analysis of recurrence to a more general class of open quantum walks, such as irreducible coins.
+In this context, recurrence criteria for higher-order HOQWs have been established in [17], where the authors derive such criteria for non-lazy HOQWs with two-dimensional coins as well as for arbitrary finite dimensions in the case of irreducible coins.
 In this article, we present more general criteria for the recurrence of lazy and non-Lazy HOQWs, based on the transition operators $L,B,R$ when acting on finite-dimensional vector spaces, that is, we allow the HOQW to be reducible.
-Furthermore, we present the first recurrence criterion for lazy HOQWs, using results obtained from recent studies on Central Limit Theorems for this kind of HOQWs.
-A first Central Limit Theorem (CLT) for Open Quantum Walks (OQWs) was established in [1].
+Furthermore, we present the first recurrence criterion for Lazy HOQWs, using results obtained from recent studies on CLTs for this kind of HOQWs.
+A first CLT for OQWs was established in [1].
 In this work, we rely on more recent developments, including the CLT for Lazy HOQWs presented in [16], as well as the general CLT for HOQWs on lattices with finite internal degrees of freedom given in [11].
 By making use of the ideas above and generalizing results from the classical version to recurrence, we also obtain a recurrence criterion for the HOQW on the grid, which is the case where the set of vertices is $\mathbb{Z}^{2}$; however, this is achieved through continuous-time OQWs (CTOQWs).
-For this purpose, we use recent results by [20], where the introduction of the open quantum jump chain allowed for the association between recurrence in discrete-time and continuous-time versions of OQWs.
+For this purpose, we use recent results of [22], where the introduction of the open quantum jump chain allowed for the association between recurrence in discrete-time and continuous-time versions of OQWs.
 This work is structured as follows.
-In Section 2, we define the open quantum walks induced by a coin and their transition probabilities.
-Section 3 is devoted to presenting a recurrence criterion for a class of Lazy HOQWs, and one restricted to Lazy HOQWs of dimension 2.
-In Section 4, we focus on the case where the coin is non-lazy, and we present a general recurrence criterion for finite-dimensional coins.
-Section 5 illustrates a first pack of comprehensive examples.
-Section 6 discusses OQWs on the grid, and Section 7 brings examples of HOQWs on those walks.
+In Section 2, we define the OQWs induced by a coin and their transition probabilities.
+Section 3 is devoted to presenting a recurrence criterion for Lazy HOQWs under ergodicity.
+Section 4 discusses recurrence under ergodicity for discrete-time and continuous-time OQWs on the grid.
+General recurrence criteria for HOQWs on the line and on the grid are detailed on Section 5.
+Section 6 illustrates a pack of comprehensive examples.
+While a unified criterion encompassing both laziness and reducibility is theoretically desirable, the primary goal of this paper is to elucidate the recurrence properties of lazy walks in the ergodic regime.
+The effect of reducibility is addressed exclusively in the non-lazy regime.
+A combined treatment would require a more involved technical framework that could obscure these distinct contributions.
+For this reason, we have chosen to present them separately, leaving the general lazy reducible case as a natural direction for future research.
+For the reader’s convenience, we summarize in Table 1 the main recurrence criteria for discrete-time HOQWs established throughout this paper, indicating for each result the graph ($\mathbb{Z}$ or $\mathbb{Z}^{2}$), the dimension of the internal space, whether the walk is lazy or non-lazy, and whether it applies to the ergodic or general (reducible) case.
+For continuous-time HOQWs on the grid, the main result is Theorem 4.13, which provides a recurrence criterion under the assumption that the auxiliary map has at most one minimal enclosure.
+Criterion Graph Dimension Lazy/Non-Lazy Type
+Theorem 3.8 $\mathbb{Z}$ Finite $d$ Lazy Ergodic
+Theorem 5.2 $\mathbb{Z}$ $d=2$ Lazy General
+Theorem 5.6 $\mathbb{Z}$ Finite $d$ Non-Lazy General
+Theorem 4.15 $\mathbb{Z}^{2}$ Finite $d$ Non-Lazy Ergodic
+Theorem 5.9 $\mathbb{Z}^{2}$ Finite $d$ Non-Lazy General
+Table 1: Summary of discrete-time recurrence criteria.
 2 General Settings
@@
 The label $i\in\mathbb{Z}$ represents the position of the particle and, when the particle is located at the vertex $i\in\mathbb{Z}$, its internal state is encoded in the space $\mathfrak{h}_{i},$ as it will be detailed below.
+2.1 Lazy Open Quantum Walks
 We consider the set of density operators over a finite Hilbert space $\mathcal{W},$
@@
 $\mathcal{S}(\mathcal{H}\otimes\mathcal{K})=\left\{\tau\in\mathcal{I}(\mathcal{H}\otimes\mathcal{K}):\textstyle\tau=\displaystyle\sum_{j\in\mathbb{Z}}\rho(j)\otimes\lvert j\rangle\langle j\rvert,\quad\rho(j)\geq 0,\quad\sum_{j\in\mathbb{Z}}\mathrm{Tr}\left(\rho(j)\right)=1\right\},$
-the set of density operators acting on $\mathcal{H}\otimes\mathcal{K}.$ We are interested in HOQWs of the form
+the set of density operators acting on $\mathcal{H}\otimes\mathcal{K}.$ Also, given a vector space $T$, we denote the set of bounded operators acting on $T$ by $\mathcal{B}(T)$.
+We are interested in HOQWs of the form
 $\Phi(\tau)=\sum_{i\in\mathbb{Z}}\Big(L\rho(i+1)L^{*}+B\rho(i)B^{*}+R\rho(i-1)R^{*}\Big)\otimes|i\rangle\langle i|,\;\;\;\tau\in\mathcal{S}(\mathcal{H}\otimes\mathcal{K}),$ (1)
@@
 Therefore, we can assume that $\mathfrak{h}_{i}=\mathfrak{h}\;\forall i\in\mathbb{Z},$ and if $\mathfrak{h}=\mathbb{C}^{d}$ for some $d$, the dimension of the coin is said to be $d$.
-In this article, we consider only the case $d<\infty.$
+In this article, we consider only the finite-dimensional case $d<\infty.$
 If $B\neq 0,$ the quantum walk exhibits a lazy behavior, meaning that in the context of quantum walks, there is a non-zero probability for the walker to stay in the same vertex after a step.
@@
 […]
-Figure 2: Lazy Quantum Walk on $\mathbb{Z}$.
+Figure 2: Lazy HOQW on $\mathbb{Z}$.
 The quantum trajectory of this kind of HOQW on the line, induced by a coin $(L,B,R)$ and starting from a state $\tau$ of the form $\tau=\sum_{i\in\mathbb{Z}}\tau_{i}\otimes|i\rangle\langle i|,$ is any path generated by the Markov chain $(X_{n},\tau_{n})_{n\geq 0}$, where $X_{n}$ denotes the position of the particle at time $n$ and $\tau_{n}$ its internal degree.
@@
     $\frac{R_{i}\rho R_{i}^{*}}{\operatorname{Tr}(R_{i}\rho R_{i}^{*})}\otimes\lvert i+1\rangle\langle i+1\rvert$.
+While irreducibility has been the focus of previous studies on the asymptotic behavior of HOQWs, this work extends the analysis to the reducible case.
+We now recall the distinction between these two notions.
+We consider a coin $(L,B,R)$ and its auxiliary map $\mathcal{L}:\mathcal{B}(\mathfrak{h})\rightarrow\mathcal{B}(\mathfrak{h})$ given by
+$\mathcal{L}(\rho)=L\rho L^{*}+B\rho B^{*}+R\rho R^{*},\quad\rho\in\mathcal{B}(\mathfrak{h}).$
+Note that $\mathcal{L}$ is a quantum channel, thus it has at least one invariant state, that is, there exists some density operator $\rho_{\infty}\in\mathcal{B}(\mathfrak{h})$ satisfying $\mathcal{L}(\rho_{\infty})=\rho_{\infty}.$ When the invariant state of $\mathcal{L}$ is unique, we will say that the map $\mathcal{L}$ is ergodic.
+As in classical Markov chains, an irreducible HOQW cannot be decomposed into dynamically independent subsystems, ensuring well-defined ergodic properties [12].
+Proposition 2.1.
+[12, Proposition 4.1] Let $\Phi$ be an HOQW defined by transition operators $L_{s}$, and $\mathcal{L}$ its auxiliary map.
+  1. 1.
+    $\mathcal{L}$ is irreducible if and only if the operators $\{L_{s},s\in S\}$ have no nontrivial closed invariant subspace in common.
+  2. 2.
+    $\Phi$ is irreducible if and only if all possible concatenations of transitions starting and ending at the same site have no nontrivial closed invariant subspace in common.
+A direct consequence is the following:
+Corollary 2.2.
+[12, Corollary 4.2] If $\Phi$ is irreducible, then $\mathcal{L}$ is irreducible.
+This corollary implies that if the auxiliary map $\mathcal{L}$ is reducible, then the HOQW $\Phi$ is also reducible.
+As noted in [12, Remark 3.6], an irreducible map admits at most one invariant state, and if such a state exists it must be faithful.
+Conversely, if a map has a unique faithful invariant state, then it is irreducible.
+This connection between algebraic structure and spectral properties provides a practical way to verify irreducibility in concrete examples.
+For instance, let $\mathfrak{h}=\mathbb{C}^{2}$, $S=\{-1,+1\}$, and denote $L_{-1}=L$, $L_{+1}=R,$ where
+$L=\frac{1}{\sqrt{20}}\begin{bmatrix}4&-1\\ 0&2\end{bmatrix},\qquad R=\frac{1}{\sqrt{20}}\begin{bmatrix}2&2\\ 0&\sqrt{11}\end{bmatrix}$.
+The auxiliary map $\mathcal{L}$ has a unique invariant state $\rho_{\infty}=\begin{bmatrix}1&0\\ 0&0\end{bmatrix}$, which is not faithful.
+Therefore, $\mathcal{L}$ is reducible, and consequently $\Phi$ is also reducible.
 3 Recurrence Criteria for Lazy Open Quantum Walks
@@
 The positive semi-definiteness criterion relaxes the strictness of positive definiteness, allowing for a broader range of quantum state representations, while still maintaining the one-to-one correspondence between quantum states and density operators.
-An important subset of the set of non-faithful density operators is composed of the pure density operators.
+An important subset of non-faithful density operators is composed of the pure density operators.
 A pure density is a density $\rho$ that is a projection of rank 1, i.e., there exists some vector $\lvert u\rangle\in\mathbb{C}^{d}$ such that $\rho=\lvert u\rangle\langle u\rvert$.
-Let us consider an HOQW starting the walk on vertex $\lvert i\rangle$ with initial density operator $\rho.$ For evolution on time $n=0,1,2,\ldots,$ we define
+Let us consider an HOQW starting the walk at vertex $\lvert i\rangle$ with initial density operator $\rho.$ For evolution at time $n=0,1,2,\ldots,$ we define
 […]
@@
     $\rho$-recurrent if
-    $\sum_{n=0}^{\infty}p_{ii;\rho}(n)=\infty.$ (2)
+    ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{E}_{i,\rho}(n_{i}):=}\sum_{n=0}^{\infty}p_{ii;\rho}(n)=\infty,\quad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{ where }p_{ii;\rho}(0)=1.}$ (2)
     Otherwise, $\lvert i\rangle$ is said to be $\rho$-transient;
@@
 Concerning all vertices, an OQW is called: recurrent if every vertex is recurrent; transient if every vertex is transient.
-Note that equation (2) computes the mean number of returns to site $\lvert i\rangle.$
-The following Proposition describes $\rho$-recurrence in a concentrated vertex $i\in\mathbb{Z}$ when assumed that $\lvert i\rangle$ is recurrent for some density and $\mathfrak{h}$ has a finite internal degree of freedom.
+Note that equation (2) computes the mean number of returns to site $\lvert i\rangle.$ The choice of this definition, based on the mean number of returns, rather than the probability of return equals 11 ($\mathbb{P}_{i,\rho}(t_{i}<\infty)<1$, where $t_{i}$ is the returning time to $i$) definition, requires a careful justification.
+As demonstrated in [4], for OQWs these two notions are not equivalent, unlike the classical case.
+This non-equivalence makes a conceptual choice necessary regarding which definition to adopt.
+The definition adopted in this work, based on the mean number of returns $\mathbb{E}_{i,\rho}(n_{i})$, follows the terminology proposed in [4] for semifinite irreducible OQWs.
+As established in [4, Corollary 3.10], for irreducible semifinite OQWs the value of $\mathbb{E}_{i,\rho}(n_{i})$ is universal: it is either finite for every $i\in V$ and every $\rho\in\mathcal{D}(\mathfrak{h}_{i})$, or infinite for every $i$ and every $\rho$.
+This universality does not hold for the return probability $\mathbb{P}_{i,\rho}(t_{i}<\infty)$, which may assume different values for different internal states even in irreducible OQWs (as illustrated in [4, Example 5.2]).
+Therefore, the definition based on the mean number of visits is more suitable for a robust and consistent classification of asymptotic behavior, connecting naturally to the limit theorems used throughout this work and allowing the characterization of recurrence via the drift parameter $m$, as developed in the following sections.
+It is worth recalling that in the particular case of irreducible HOQWs, Jacq and Lardizabal [17] demonstrated that these two notions of recurrence do coincide.
+However, their proof relies heavily on irreducibility assumptions, particularly on the uniqueness and faithfulness of the invariant state, which forces the equivalence between the divergence of the mean number of returns and the certainty of return.
+For reducible HOQWs, the relationship between these two definitions remains an open problem.
+The existence of multiple invariant states and the possibility of nontrivial transient components within the internal space $\mathfrak{h}$ may lead to situations where a vertex is $\rho$-recurrent under one definition but not the other, depending on the initial internal state.
+This question, while beyond the scope of the present work, represents an interesting direction for future investigation.
+Having clarified the choice of recurrence definition and its nuances, we now proceed to the following proposition, which describes $\rho$-recurrence in a concentrated vertex $i\in\mathbb{Z}$ when assumed that $\lvert i\rangle$ is recurrent for some density and $\mathfrak{h}$ has a finite internal degree of freedom.
+This result establishes the behavior of $\rho$-recurrence when recurrence occurs for at least one initial state.
+In the following sections, we will build upon this to analyze the structure of the internal space $\mathfrak{h}$ and understand how it decomposes in order to develop recurrence criteria for reducible OQWs.
 An immediate consequence will also be given for the case where $\lvert i\rangle$ is $\sigma$-transient for some non-faithful density.
@@
 $f(\tau)=\inf\{c>0\mid\rho>c\tau\}$
-is continuous and thus attains its minimum and maximum. In particular, there exists $c>0$ such that
+is upper semicontinuous and thus attains its minimum and maximum. In particular, there exists $c>0$ such that
 $\rho>c\tau,\quad\forall\rho\in\mathcal{D}(\mathfrak{h}).$
@@
 […]
-where at least one of the members of the sum must be infinity. This completes the proof.
+where at least one term in the sum must be infinite. This completes the proof.
 $\square$
-By a contra-position argument, we obtain the following direct consequence of Theorem 3.1.
+By a contraposition argument, we obtain the following direct consequence of Theorem 3.1.
 Corollary 3.2.
@@
 3.2 Enclosures and Invariant States
-Firstly, a subspace $\mathcal{Y}$ of $\mathfrak{h}$ is called an enclosure for a quantum channel $\mathcal{L}$ if for every $\sigma\in\mathcal{D}\mathfrak{h}),$ we have
+A subspace $\mathcal{Y}$ of $\mathfrak{h}$ is called an enclosure for a quantum channel $\mathcal{L}$ if for every $\sigma\in\mathcal{D}(\mathfrak{h}),$ we have
 $\mathrm{supp}(\sigma)\subset\mathcal{Y}\Rightarrow\mathrm{supp}\left(\mathcal{L}(\sigma)\right)\subset\mathcal{Y}.$
 If an enclosure contains no other non-trivial enclosures, then it is called a minimal enclosure.
-A minimal enclosure with finite dimension is the support of a unique invariant state of the channel, that is, a density operator $\sigma$ such that $\mathcal{L}(\sigma)=\sigma$.
-For more on enclosures of HOQWs, we refer the reader to [10, 11, 4].
+A minimal enclosure with finite dimension is the support of a unique invariant state of the channel, that is, a density operator $\sigma$ such that $\mathcal{L}(\sigma)=\sigma$ [10, Section 2.2].
+For more on enclosures of HOQWs, we refer the reader to [10, 11, 4, 15].
 Following this property and using the homogeneity of HOQWs induced by a coin, we split $\mathfrak{h}$ into two subsets $\mathcal{Y}$ and $\mathcal{X}$ in such a way that the walk is transient for $\mathcal{D}(\mathcal{Y})$ and recurrent everywhere else, with $\mathcal{Y}$ being an enclosure of the quantum channel with Kraus operators $L,B,R.$
@@
 where $J_{\pi}$ is a concatenation of $n$ operators, each one belonging to $\{L,B,R\},$ proving the claim, once this is valid for any $n=1,2,3,\ldots$
-Moreover, by equation (4) we have that $\cup_{n}\;\text{supp}(\rho_{n})$ is a subspace of $\mathfrak{h}.$ Joining all the possible initial states in which $(L,B,R)$ is transient, we obtain a decomposition (3), where $\mathcal{Y}$ or $\mathcal{X}$ may be $\{0\}$, and by definition $\mathcal{Y}$ is an enclosure.
+[…]
 Let us suppose that $(L,B,R)$ is neither recurrent nor transient.
@@
 $\square$
-In our subsequent discussion, we consider a coin $(L,B,R)$ and its auxiliary map $\mathcal{L}:\mathcal{B}(\mathfrak{h})\rightarrow\mathcal{B}(\mathfrak{h})$ given by
-$\mathcal{L}(\rho)=L\rho L^{*}+B\rho B^{*}+R\rho R^{*},\quad\rho\in\mathcal{B}(\mathfrak{h}).$
-Note that $\mathcal{L}$ is a quantum channel, thus it has at least one invariant state, that is, there exists some density operator $\rho_{\infty}\in\mathcal{B}(\mathfrak{h})$ satisfying $\mathcal{L}(\rho_{\infty})=\rho_{\infty}.$ When the invariant state of $\mathcal{L}$ is unique, we will say that the map $\mathcal{L}$ is ergodic.
-We will use the following result obtained straightforwardly from [14].
+3.3 Recurrence under Ergodicity
+Following the CLT approach for lazy OQWs given in [16], we consider the Markov chain $(\rho_{n},\Delta X_{n})_{n\geq 0}$ associated with the quantum trajectory description of the walk.
+Let $m\in\mathbb{R}^{d}$ denote the drift vector of the walk, given by the expected value of the position increment under the invariant state $\rho_{\infty}$, namely
+$m=\sum_{k}\mathrm{Tr}(A_{k}\rho_{\infty}A_{k}^{*})\,e_{k},$
+where $(e_{k})_{k}$ are the possible displacement vectors of the walk, $A_{k}$ are the transition operators.
+For a fixed vector $l\in\mathbb{R}^{d}$, the quantity $(X_{n}-nm)\cdot l$ admits a Doob decomposition into the sum of a martingale part and a predictable process.
+The martingale term is given by
+$M_{n}=\sum_{j=2}^{n}\left[f(\rho_{j},\Delta X_{j})-Pf(\rho_{j-1},\Delta X_{j-1})\right],$ (5)
+where $f:\mathcal{D}(\mathfrak{h})\times\mathbb{R}^{d}\to\mathbb{R}$ is defined by
+$f(\rho,x)=\mathrm{Tr}(\rho L_{l})+x\cdot l,$
+and $L_{l}=L\cdot l$ denotes the operator associated with the vector $l\in\mathbb{R}^{d}$ acting on the internal Hilbert space $\mathfrak{h}$.
+Here $\rho_{j}$ denotes the internal state of the walker at time $j$, $\Delta X_{j}=X_{j}-X_{j-1}$ represents the increment of the position process, and $P$ is the transition operator of the Markov chain $(\rho_{n},\Delta X_{n})$.
 Theorem 3.6.
@@
 Proof.
-This result can be obtained by using a Doob decomposition as described in [14, Equation 27] and [1, Theorem 5.2] for a Martingale $M_{n}$ satisfying
+This result can be obtained by using the Doob decomposition as described in Equation (5) and [1, Theorem 5.2] for the Martingale $M_{n}$ satisfying
 $X_{n}-n.m=X_{0}+M_{n}+\beta_{n},$
-where $M_{n}/n\rightarrow 0$ and $\beta_{n}\rightarrow 0$ when $n\rightarrow 0.$
+[…]
 $\square$
@@
 Proof.
-The proof is analogous to [15, Theorem 7] and we will proceed by remarking on how it can be generalized.
-Firstly, the maximum appearing in [15, Lemma 6] can be taken in $\mathcal{D}=\mathcal{D}(\mathcal{Y}).$ Indeed, if the initial state is $\rho\in\mathcal{D}(\mathcal{Y}),$ then $\rho_{n}\in\mathcal{D}(\mathcal{Y})$ for every $n,$ thereby the maximum taken in the proof of [15, Lemma 6] can be reduced to $\mathcal{D}(\mathcal{Y}).$ Therefore, by the same reason, the density operator considered in [15, Theorem 7] belongs to $\mathcal{D}(\mathcal{Y}),$ proving the Theorem for the non-lazy case.
+The proof is analogous to [17, Theorem 7] and we will proceed by remarking on how it can be generalized.
+Firstly, the maximum appearing in [17, Lemma 6] can be taken in $\mathcal{D}=\mathcal{D}(\mathcal{Y}).$ Indeed, if the initial state is $\rho\in\mathcal{D}(\mathcal{Y}),$ then the density operator $\rho_{n}$ defined in equation (4) lies on $\mathcal{D}(\mathcal{Y})$ for every $n,$ otherwise $\mathcal{Y}$ would not be an enclosure, by definition, thereby the maximum taken in the proof of [17, Lemma 6] can be reduced to $\mathcal{D}(\mathcal{Y}).$ Therefore, by the same reason, the density operator considered in [17, Theorem 7] belongs to $\mathcal{D}(\mathcal{Y}),$ proving the Theorem for the non-lazy case.
 Moreover, the lazy case holds as an extension of this proof, since all the results work for HOQWs on $\mathbb{Z}$.
@@
 The following theorem presents the first recurrence criterion for non-irreducible HOQWs of arbitrary finite dimension.
-This result extends [15, Corollary 8 and Proposition 8] for reducible CTOQWs and can also be applied to Lazy HOQWs.
+This result extends [17, Corollary 8 and Proposition 8] for reducible CTOQWs and can also be applied to Lazy HOQWs.
 Theorem 3.8 (Recurrence Criteria).
-Let us consider an HOQW on $\mathbb{Z}$ induced by a coin $(L,B,R)$ of dimension $d$ such that the auxiliary map $\mathcal{L}$ is ergodic.
-Then $\mathcal{L}$ has a unique invariant state $\rho_{\infty}$ and
+Let us consider an HOQW on $\mathbb{Z}$ induced by a coin $(L,B,R)$ of dimension $d$ such that the auxiliary map $\mathcal{L}$ is ergodic (hence $\mathcal{L}$ has a unique invariant state $\rho_{\infty}$).
+Under this assumption, the mixed case described in Proposition 3.5 cannot occur, and we have the following dichotomy:
   • $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)=\textmd{Tr}\left(R\rho_{\infty}R^{*}\right)\Rightarrow$ the walk is recurrent;
@@
 Moreover, the unique invariant state $\rho_{\infty}$ of $\mathcal{L}$ must belong to $\mathcal{D}(\mathcal{Y}).$
-If $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)=\textmd{Tr}\left(R\rho_{\infty}R^{*}\right)$ then $X_{n}/n\rightarrow 0$ by Theorem 3.6, thus $(L,B,R)$ is $\rho$-recurrent for some $\rho\in\mathcal{D}(\mathcal{Y})$ by Theorem 3.7, which is a contradiction.
+If $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)=\textmd{Tr}\left(R\rho_{\infty}R^{*}\right),$ then $X_{n}/n\rightarrow 0$ by Theorem 3.6, thus $(L,B,R)$ is $\rho$-recurrent for some $\rho\in\mathcal{D}(\mathcal{Y})$ by Theorem 3.7, which is a contradiction.
 On the other hand, if $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)\neq\textmd{Tr}\left(R\rho_{\infty}R^{*}\right),$ then $X_{n}/n\rightarrow m\neq 0.$ Therefore $\mathbb{P}_{0,\rho}(X_{n}=0\mbox{ i.o.})=0,$ showing that the mean number of returns to the initial site is finite for any $\rho$, thus the walk is transient, which is also a contradiction.
@@
 Therefore, if $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)=\textmd{Tr}\left(R\rho_{\infty}R^{*}\right),$ then the walk is $\rho$-recurrent for some density and thus it is recurrent.
-If $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)\neq\textmd{Tr}\left(R\rho_{\infty}R^{*}\right),$ then the walks is transient for every density.
+If $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)\neq\textmd{Tr}\left(R\rho_{\infty}R^{*}\right),$ then the walk is transient for every density.
 $\square$
+Note that the mixed case from Proposition 3.5 does not occur on Theorem 3.8, as ergodicity ensures a strict dichotomy.
 We remark that the operator $B$ does influence in the recurrence of $(L,B,R).$ Indeed, even it does not appear on the bullet points that describe the recurrence criterion, the invariant state $\rho_{\infty}$ may change when we pick distinct $B$’s.
 See example 6.3 below.
-3.3 General Recurrence Criteria for HOQWs induced by a coin $(L,B,R)$ of dimension 2
+In addition to, a distinction must be drawn between lazy and non-lazy walks regarding their recurrence behavior.
+In the non-lazy case ($B=0$), the walker is forced to jump at every step, and recurrence is determined solely by the balance between left and right transitions [17].
+In the lazy case ($B\neq 0$), the presence of a nonzero “stay” probability introduces additional freedom: the walker may remain at the same site, which affects the asymptotic drift and the structure of invariant states.
+Consequently, recurrence criteria for lazy walks involve the full triple $(L,B,R)$, while non-lazy criteria depend only on $(L,R)$.
+This distinction is reflected throughout Theorem 3.8, which provides a criterion for lazy walks under ergodicity.
+It is worth comparing our results with those of [17].
+In that work, the authors established recurrence criteria for irreducible OQWs in the non-lazy setting, with recurrence occurring precisely when $\mathrm{Tr}(L\rho_{\infty}L^{*})=\mathrm{Tr}(R\rho_{\infty}R^{*})=1/2$.
+Our Theorem 3.8 extends this analysis to lazy walks under the assumption that the auxiliary map $\mathcal{L}$ is ergodic.
+In this context, recurrence is determined by the condition $\mathrm{Tr}(L\rho_{\infty}L^{*})=\mathrm{Tr}(R\rho_{\infty}R^{*})$, without requiring this common value to be $1/2$.
+The value $1/2$ in [17] is a consequence of the irreducibility assumption and the particular normalization used there.
+By taking $B=0$ and assuming irreducibility in our framework, we recover the criterion of [17] as a special case.
+4 2D Open Quantum Walks
+We start this section with the continuous-time OQW.
+4.1 Continuous-Time Open Quantum Walks
+In order to define the CTOQWs, we recall that an operator semigroup $\mathcal{T}$ on a Hilbert space $\mathcal{K}$ is a family of bounded linear operators $(T_{t})$ acting on $\mathcal{K}$, $t\geq 0,$ such that
+$T_{t}T_{s}=T_{t+s},\;\forall s,t\in\mathbb{R}_{+}\;\mbox{ and }\;T_{0}=I_{\mathcal{K}}.$
+If $t\mapsto T_{t}$ is continuous for the operator norm of $\mathcal{K}$, then $\mathcal{T}$ is said to be uniformly continuous, which is equivalent to the following assertion [7]: There exists a bounded linear operator $L$ on $\mathcal{K}$ such that $T_{t}=e^{tL},\;\forall t\geq 0.$ In this case,
+$L=\lim_{t\rightarrow\infty}\frac{1}{t}(T_{t}-I_{\mathcal{K}})$
+and the operator $L$ is called the generator of $\mathcal{T}.$
+[…]
+Definition 4.1.
+Let $V$ be a finite or countable infinite set and $\mathcal{H}$ be a Hilbert space of the form (2).
+A Continuous-time Open Quantum Walk in $V$ is an uniformly continuous QMS on $\mathcal{I}_{1}(\mathcal{H})$ with Lindblad operator of the form
+$\begin{split}\mathcal{\mathcal{L}}:\mathcal{I}_{1}(\mathcal{H})&\rightarrow\mathcal{I}_{1}(\mathcal{H})\\ \rho&\mapsto-i[\textbf{H},\rho]+\sum_{i,j\in V}\left(S_{i}^{j}\rho S_{i}^{j^{*}}-\frac{1}{2}\{S_{i}^{j*}S_{i}^{j},\rho\}\right),\end{split}$ (6)
+where $A,B]\equiv AB-BA$ and $\{A,B\}\equiv AB+BA.$
+To rearrange Equation (6), we can write $S_{i}^{j}=R_{i}^{j}\otimes\lvert j\rangle\langle i\rvert$ for bounded operators $R_{i}^{j}:\mathfrak{h}_{i}\to\mathfrak{h}_{j},$ the bounded operators H and $S_{i}^{j}$ on $\mathcal{H}$ are of the form $\textbf{H}=\sum_{i\in V}H_{i}\otimes\lvert i\rangle\langle i\rvert$, $H_{i}$ is self-adjoint on $\mathfrak{h}_{i},$ and $\sum_{i,j\in V}S_{i}^{j*}S_{i}^{j}$ converges in the strong sense.
+Letting $\rho=\sum_{i\in V}\rho(i)\otimes\lvert i\rangle\langle i\rvert\in\mathcal{S}(\mathcal{H}\otimes\mathcal{K}),$ denoting
+$e^{t\mathcal{L}}(\rho)=\mathcal{T}_{t}(\rho)=\sum_{i\in V}\rho_{t}(i)\otimes\lvert i\rangle\langle i\rvert,\quad\forall t\geq 0,$
+we have
+$\frac{d}{dt}\rho_{t}(i)=-i[H_{i},\rho_{t}(i)]+\sum_{j\in V}\left(R_{j}^{i}\rho_{t}(j)R_{j}^{i^{*}}-\frac{1}{2}\{R_{i}^{j*}R_{i}^{j},\rho_{t}(i)\}\right).$
+Equation (6) can be written as given in [5, Equation 18.7]:
+$\mathcal{L}(\rho)=\sum_{i\in V}\left(G_{i}\rho(i)+\rho(i)G_{i}^{*}+\sum_{j\in V}R_{j}^{i}\rho(j)R_{j}^{i*}\right)\otimes\lvert i\rangle\langle i\rvert,\qquad G_{i}=-iH_{i}-\frac{1}{2}\sum_{j\in V}R_{i}^{j*}R_{i}^{j}.$ (7)
+In this paper we deal with the quantum trajectories of CTOQWs, which are more complex than the trajectories of discrete-time HOQWs.
+Thus, we recall that for a given finite or countably infinite set of vertices $V,$ a CTOQW is a stochastic process evolving on a Hilbert space of the form (2), the label $i\in V$ represents the position of the walker and when the walker is located at $i\in V,$ its internal state is encoded in $\mathfrak{h}_{i},$ that is, $\mathfrak{h}_{i}$ describes the internal degrees of freedom of the walker when it is at site $i\in V.$
+Let $i\in V$ and $\rho\in\mathcal{D}(\mathfrak{h}_{i}).$ Starting the walk at site $\lvert i\rangle$ with initial density operator $\rho,$ the quantum measurement of the “position” gives rise to a probability distribution $p_{0}$ on $V,$ such that
+$p_{0}(j)=\mathbb{P}(\mbox{the quantum particle is in site}\;\lvert j\rangle)=\mathrm{Tr}(\rho(j)),$
+and for evolution at time $t\geq 0,$
+[…]
+where
+$e^{t\mathcal{L}}(\rho)=\sum_{k\in V}\rho_{t}(i)\otimes\lvert k\rangle\langle k\rvert$.
+4.2 Quantum Trajectories
+Let $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$ be a probability space where independent Poisson point processes $N^{ij},i,j\in V,\;i\neq j$ ($N^{ii}=0$ by convention) on $\mathbb{R}^{2}$ are defined.
+The jump from site $i$ to site $j$ on the graph $V$ will be governed by these Poisson point processes [5]:
+Definition 4.2.
+Consider a CTOQW with generator of the form (7) and an initial density operator $\mu=\sum_{i\in V}\rho(i)\otimes\lvert i\rangle\langle i\rvert\in\mathcal{S}(\mathcal{H}\otimes\mathcal{K}).$ The quantum trajectory describing the indirect measurement of the position of the CTOQW is the Markov chain represented by the density operators $(\mu_{t})_{t\geq 0}$ such that $\mu_{0}=\rho_{0}\otimes\lvert X_{0}\rangle\langle X_{0}\rvert$, where $X_{0}$ and $\rho_{0}$ are random variables with distribution
+$\mathbb{P}\left((X_{0},\rho)=\left(i,\frac{\rho(i)}{\mathrm{Tr}(\rho(i))}\right)\right)=\mathrm{Tr}\left(\rho(i)\right)\mbox{ for all }i\in V,$
+and $\mu_{t}=:\rho_{t}\otimes\lvert X_{t}\rangle\langle X_{t}\rvert$ satisfies the stochastic differential equation
+$\begin{split}\mu_{t}=&\mu_{0}+\int_{0}^{t}M(\mu_{s^{-}})ds\\ +&\sum_{ij}\int_{0}^{t}\int_{\mathbb{R}}\left(\frac{S_{i}^{j}\mu_{s^{-}}S_{i}^{j*}}{\mathrm{Tr}(S_{i}^{j*}\mu_{s^{-}}S_{i}^{j*})}-\mu_{s^{-}}\right)1_{0<y<\mathrm{Tr}(S_{i}^{j}\mu_{s^{-}}S_{i}^{j*})}N^{ij}(dy,ds)\end{split}$ (9)
+for all $t\geq 0,$ where
+$M(u)=\mathcal{L}(u)-\sum_{ij}\left(S_{i}^{j}\mu S_{i}^{j*}-\mu\mathrm{Tr}(S_{i}^{j}\mu S_{i}^{j*})\right).$
+Hence, for a fixed $\mu=\sum_{i}\rho(i)\otimes\lvert i\rangle\langle i\rvert\in\mathcal{S}(\mathcal{H}\otimes\mathcal{K}),$
+$M(\mu)=\sum_{i}\left(G_{i}\rho(i)+\rho(i)G_{i}^{*}-\rho(i)\mathrm{Tr}\left(G_{i}\rho(i)+\rho(i)G_{i}^{*}\right)\right)\otimes\lvert i\rangle\langle i\rvert$.
+For simplicity, suppose $X_{0}=i_{0}$ for some $i_{0}\in V$ and $\rho_{0}\in V(\mathfrak{h}_{i_{0}}).$ For all $t\geq 0,$ consider the solution of Equation (9)
+$\eta_{t}=\rho_{0}+\int_{0}^{t}\left(G_{i_{0}}\eta_{s}+\eta_{s}G_{i_{0}}^{*}-\eta_{s}\mathrm{Tr}\left(G_{i_{0}}\eta_{s}+\eta_{s}G_{i_{0}}^{*}\right)\right)ds,$
+which is a density operator on $\mathfrak{h}_{i_{0}}.$ For $j\neq i_{0},$ define
+$T_{1}^{j}=\mathrm{inf}\left\{t\geq 0;N^{i_{0},j}\left({u,y|0\leq u\leq t,0\leq y\leq\mathrm{Tr}(R_{i_{0}}^{j}\eta_{u}R_{i_{0}}^{j*})}\right)\geq 1\right\}.$
+Since the random variables $T_{1}^{j}$ are mutually independent and nonatomic, we can define $T_{1}=\inf_{j\neq i_{0}}\{T_{1}^{j}\}$ once there exists a unique $j\in V$ such that $T_{1}^{j}=T_{1}.$ The random variable $T_{1}$ is said to be the first jump time of the CTOQW conditional on $X_{0}=i_{0}.$
+The first jump time to site $\lvert j\rangle$ is then denoted by $T_{1}^{j}$ and has distribution
+$\mathbb{P}(T_{1}^{j}>\varepsilon)=e^{-\int_{0}^{\varepsilon}\mathrm{Tr}(R_{i_{0}}^{j}\eta_{u}R_{i_{0}}^{j*})du},$
+thus
+$\mathbb{P}(T_{1}\leq\varepsilon)\leq\varepsilon\sum_{j\neq i_{0}}\|R_{i_{0}}^{j*}R_{i_{0}}^{j}\|.$
+The strong convergence of $\sum_{ij}S_{i}^{j*}S_{i}^{j}$ implies that $\mathbb{P}(T_{1}>0)=1.$ Thereby, on $0,T_{1}],$ we can define the solution $(X_{t},\rho_{t})_{t\geq 0}$ as
+$(X_{t},\rho_{t})$ $=$ $(i_{0},\eta_{t})\mbox{ for }t\in[0,T_{1})\mbox{ and }$
+$(X_{T_{1}},\rho_{T_{1}})$ $=$ $\left(j,\frac{R_{i}^{j}\eta_{T_{1}-}R_{i}^{j*}}{\mathrm{Tr}(R_{i}^{j}\eta_{T_{1}-}R_{i}^{j*})}\right)\;\mbox{ if }T_{1}=T_{1}^{j}.$
+Now we solve
+$\eta_{t}=\rho_{T_{1}}+\int_{0}^{t}\left(G_{j}\eta_{s}+\eta_{s}G_{j}^{*}-\eta_{s}\mathrm{Tr}\left(G_{j}\eta_{s}+\eta_{s}G_{j}^{*}\right)\right)ds$
+and then obtain the second jump time $T_{2}.$ And so on we obtain an increasing sequence of jumps $(T_{n})_{n}$ with $\lim_{n\rightarrow\infty}T_{n}=\infty$ almost surely (see section 18.2.3 of [5] for more details).
+This means that the walk does not explode, thus the walker has a finite number of jumps in a finite interval.
+For details concerning explosions of classical Markov chains, see [25, Section 2.2].
+4.3 Recurrence of CTOQWs
+Moving forward, we deal with the following class of 2-dimensional CTOQWs, that is, we will consider the nearest-neighbor quantum random walk on $V=\mathbb{Z}^{2}.$
+Definition 4.3.
+Consider a CTOQW on $V=\mathbb{Z}^{2}$. If there exist $A_{1},A_{2},A_{3},A_{4},H\in\mathcal{B}(\mathbb{C}^{n})$ such that
+$H_{k}=H\;\forall k\in\mathbb{Z}^{2},\quad A_{1}=R_{(i,j)}^{(i+1,j)},\;A_{2}=R_{(i,j)}^{(i,j+1)},\;A_{3}=R_{(i,j)}^{(i-1,j)},\;A_{4}=R_{(i,j)}^{(i,j-1)},$
+and $R_{x}^{y}=0_{n}$ for the remaining $(x,y)\in\mathbb{Z}^{2},$ then the CTOQW is said to be induced by a coin $(A,H)$ of dimension $n.$
+For this kind of CTOQW we can let $\mathfrak{h}_{i}=\mathfrak{h}$ for all $i\in\mathbb{Z}^{2}.$ If $\dim(\mathfrak{h})=n,$ then $A_{1},A_{2},A_{3},A_{4}$ and $H$ can be represented by square matrices of order $n.$
+The auxiliary map of a coin $(A,H)$ is defined as
+$\mathbb{L}(\rho)=G\rho+\rho G^{*}+\sum_{j=1}^{4}A_{j}\rho A_{j}^{*},\qquad G=-iH-\frac{1}{2}\sum_{k=1}^{4}A_{k}^{*}A_{k}.$
+The map $\mathbb{L}$ has at least one stationary state $\rho_{\infty},$ that is, $\mathbb{L}(\rho_{\infty})=0$ [6]. When the stationary state is unique, we define
+$m=\begin{bmatrix}m_{1}\\ m_{2}\end{bmatrix},\quad m_{1}=\mathrm{Tr}\left(A_{1}\rho_{\infty}A_{1}^{*}\right)-\mathrm{Tr}\left(A_{3}\rho_{\infty}A_{3}^{*}\right),\quad m_{2}=\mathrm{Tr}\left(A_{2}\rho_{\infty}A_{2}^{*}\right)-\mathrm{Tr}\left(A_{4}\rho_{\infty}A_{4}^{*}\right).$
+To demonstrate the recurrence criteria of this work for CTOQWs, we recall the following limit theorems for such walks.
+Theorem 4.4.
+[8, Proposition 3.0.9] Consider a coin $(A,H)$, where its auxiliary map has a stationary state $\rho_{\infty}.$ Then
+$\frac{X_{t}-tm}{\sqrt{t}}\stackrel{{\scriptstyle t\rightarrow\infty}}{{\longrightarrow}}N(0,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Sigma})$
+for some $\Sigma\in\mathbb{M}_{2}(\mathbb{C}).$
+Proposition 4.5 (Law of Large Numbers for CTOQWs ).
+[23, Theorem 15] If the auxiliary map of a coin $(A,H)$ has a unique stationary state, then
+$\frac{X_{t}}{t}\rightarrow m\quad\mbox{a.s}.$
+Let us recall the classical definition of recurrence of CTOQWs.
+For this, $p_{ji;\rho}(t)$ will denote the probability of being at site $j$ at time $t,$ given that we started at site $i,$ with initial density $\rho$ concentrated at $i$:
+[…]
+Therefore, the walk starts with a density operator $\rho$ concentrated at some vertex $\lvert i\rangle,$ takes the evolution up to time $t>0$ through the quantum Markov semigroup generated by $\mathcal{L},$ producing a new density operator
+$\rho_{t}=\sum_{k}\rho_{t}(k)\otimes\lvert k\rangle\langle k\rvert=e^{t\mathcal{L}}(\rho\otimes\lvert i\rangle\langle i\rvert),\quad\mathrm{Tr}\left(\sum_{k}\rho_{t}(k)\right)=1.$
+Now we project $\rho_{t}$ onto the subspace generated by vertex $\lvert j\rangle$ giving
+$e^{t\mathcal{L}}(\rho\otimes\lvert i\rangle\langle i\rvert)(I\otimes\lvert j\rangle\langle j\rvert)=\rho_{t}(j)\otimes\lvert j\rangle\langle j\rvert$,
+which represents the data concentrated at vertex $\lvert j\rangle$ at time $t.$
+Let $i\in V,\rho\in\mathcal{D}(\mathfrak{h}_{i}).$ We say that vertex $i$ is
+  • $\rho$-recurrent if
+    $\int_{0}^{\infty}p_{ii;\rho}(t)dt=\infty.$
+    Otherwise, $i$ is said to be $\rho$-transient;
+  • recurrent, if $i$ is $\rho$-recurrent for all $\rho\in\mathcal{D}(\mathfrak{h}_{i});$
+  • transient, if $i$ is $\rho$-transient for all $\rho\in\mathcal{D}(\mathfrak{h}_{i}).$
+By definition, $i$ is $\rho$-recurrent when the mean number of returns to $i$ is infinite.
+Definition 4.6.
+A CTOQW is said to be:
+  • recurrent if every vertex is recurrent;
+  • transient if every vertex is transient.
+Within this setting, recurrence can be naturally associated with quantum trajectories.
+This connection relies on the homogeneity of the coin, which guarantees the validity of the result stated in the next proposition for a CTOQW induced by a coin.
+Proposition 4.7.
+A CTOQW induced by a coin is $\rho$-recurrent if and only if it is $\rho_{T_{n}}$-recurrent for any possible jump time $T_{n}.$
+Proof.
+Note that if we start the walk with $(X_{0},\rho_{0})=(i,\rho),$ then at some time $t>0,$ we have $(X_{t},\rho_{t})=(j,\rho_{T}).$ If we refresh the walk at this instant, then the walk is still recurrent with respect to the density $\rho_{t},$ by the homogeneity in space.
+The other hand follows the same idea, assuming by contraposition that the walk is $\rho$-transient.
+$\square$
+Proposition 4.8.
+If $(A,H)$ is $\tau$-recurrent for some density $\tau\in\mathcal{D}(\mathfrak{h})$ and $\mathfrak{h}$ has at most one non-trivial invariant subspace for $\Phi:\mathfrak{h}\rightarrow\mathfrak{h}$ given by
+$\Phi(.)=\sum_{k}A_{k}(.)A_{k}^{*},$
+then $(A,H)$ is recurrent.
+In particular, if $(A,H)$ is $\tau$-recurrent for some density $\tau\in\mathcal{D}(\mathfrak{h})$ and $\mathfrak{h}$ has at most one minimal enclosure for $\Phi,$ then $(A,H)$ is recurrent.
+Proof.
+Let us start supposing that $(A,H)$ is recurrent with respect to some density operator $\tau$ with support on $\mathcal{V}.$ If $(A,H)$ starts the walk with initial density operator $\rho\in\mathcal{D}(\mathfrak{h}),$ then $\rho_{T_{n}}$ is of the form
+$\rho_{T_{n}}=\frac{A_{k}\eta A_{k}^{*}}{\mathrm{Tr}\left(A_{k}\eta A_{k}^{*}\right)}\text{ for some }k\in\{1,2,3,4\},$ (10)
+[…]
+$\sigma:=\rho_{T_{n}}-\frac{\tau}{a}\geq 0.$
+Therefore,
+$\frac{1}{\mathrm{Tr}(\sigma)}\int_{0}^{\infty}p_{ii;\rho_{T_{n}}}(t)\;\text{d}t=\int_{0}^{\infty}p_{ii;\frac{\sigma}{\mathrm{Tr}(\sigma)}}(t)\;\text{d}t+\frac{1}{a\mathrm{Tr}(\sigma)}\int_{0}^{\infty}p_{ii;\tau}(t)\;\text{d}t=\infty.$
+By Proposition 4.7, $(A,H)$ is $\rho$-recurrent.
+The particular case holds, since $\Phi$ has only one non-trivial invariant subspace for $A_{k},\;k=1,2,3,4,$ when there is only one minimal enclosure.
+Indeed, the invariant subspaces of $\Phi$ correspond to enclosures, since the projections in the fixed point set are exactly the projections onto enclosures (see [15, Corollary 2.2.6]).
+Moreover, space $\mathcal{R}$ admits a decomposition in minimal enclosures, as in Equation (16), where every non-trivial enclosure contains a minimal enclosure.
+Hence, if $\mathfrak{h}$ has at most one minimal enclosure for $\Phi$, then $\Phi$ has at most one non-trivial invariant subspace, and the Proposition applies.
+$\square$
+4.4 Recurrence Criterion for CTOQWs on the Grid
+This section is devoted to exhibit a recurrence criterion for finite dimensional coins $(A,H),$ where the auxiliary map has at most one common invariant subspace.
+The result uses a CLT for CTOQWs and the following Lemma.
+Lemma 4.9.
+Let us consider a homogeneous CTOQW on $\mathbb{Z}^{d}.$ Let $\varepsilon>0$ and $m\geq 1$ be an integer. Then, for every $\rho\in\mathcal{D}(\mathfrak{h})$,
+$\int_{0}^{\infty}\mathbb{P}_{0,\rho}(|X_{t}|<m\varepsilon)dt\leq(2m)^{d}\cdot\max_{\sigma\in\mathcal{D}}(\mathfrak{h})\int_{0}^{\infty}\mathbb{P}_{0,\sigma}(|X_{t}|<\varepsilon)dt.$
+Proof. Firstly, note that
+$\int_{0}^{\infty}\mathbb{P}_{0,\rho}\left(|X_{t}|<m.\varepsilon\right)dt\leq\int_{0}^{\infty}\sum_{k}\mathbb{P}_{0,\rho}\left(X_{t}\in k.\varepsilon+[0,\varepsilon)^{d}\right)dt,$ (11)
+where the sum runs over $k$ in the set $\{-m,-m+1,\ldots,m-2,m-1\}^{d}$.
+Let
+$T_{k}=\inf\left\{z\geq 0:X_{z}\in k.\varepsilon+\left[0,\varepsilon\right)^{d}\right\}\;\text{ and }\;\zeta_{\rho}^{k,\varepsilon}=\int_{0}^{\infty}\mathbb{P}_{0,\rho}\left(X_{t}\in k.\varepsilon+[0,\varepsilon)^{d}\right)dt.$
+. By the law of total probability and Fubini’s theorem, we can rewrite $\zeta_{\rho}^{k,\varepsilon}$ as
+$\begin{split}\zeta_{\rho}^{k,\varepsilon}=&\int_{0}^{\infty}\int_{0}^{t}\mathbb{P}_{0,\rho}\left(X_{t}\in k.\varepsilon+[0,\varepsilon)^{d},\;T_{k}=y\right)dy\;dt\\ =&\int_{0}^{\infty}\int_{0}^{\infty}\mathbb{P}_{0,\rho}(|X_{t}-X_{y}|<\varepsilon,T_{k}=y)dy\;dt\\ =&\int_{0}^{\infty}\int_{y}^{\infty}\mathbb{P}_{0,\rho}(|X_{t}-X_{y}|<\varepsilon,T_{k}=y)dt\;dz\\ =&\int_{0}^{\infty}\mathbb{P}_{0,\rho}(T_{k}=z)\int_{y}^{\infty}\mathbb{P}_{0,\rho}(|X_{t}-X_{y}|<\varepsilon)dt\;dy,\end{split}$ (12)
+[…]
+Denote $\mathbb{P}_{0,\rho}^{k,y}=\mathbb{P}_{0,\rho}\left(T_{k}=y\right).$ An application of Fubini’s theorem gives
+$\begin{split}\zeta_{\varepsilon,k}^{\rho}=&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}\int_{y}^{\infty}\mathbb{P}_{0,\rho}\left(|X_{t}-X_{y}|<\varepsilon\right)dtdy\\ =&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}\int_{y}^{\infty}\sum_{\sigma\in\mathcal{D}(\mathfrak{h}),j\in\mathbb{Z}}\mathbb{P}_{0,\rho}\left(|X_{t}-X_{y}|<\varepsilon,(X_{y},\rho_{y})=(j,\sigma)\right)dtdy\\ =&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}\sum_{\sigma\in\mathcal{D}(\mathfrak{h}),j\in\mathbb{Z}}\int_{y}^{\infty}\mathbb{P}_{0,\rho}\left(|X_{t}-X_{y}|<\varepsilon|(X_{y},\rho_{y})=(j,\sigma)\right)\mathbb{P}_{0,\rho}\left((X_{y},\rho_{y})=(j,\sigma)\right)dtdy.\end{split}$
+Due to the spatial homogeneity of the walk and the change of variables s=ty,s=t-y, we have
+$\begin{split}\zeta^{\rho}_{\varepsilon,k}=&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}\sum_{\sigma\in\mathcal{D}(\mathfrak{h}),j\in\mathbb{Z}}\int_{y}^{\infty}\mathbb{P}_{j,\sigma}\left(|X_{t-y}-X_{0}|<\varepsilon\right)dt\mathbb{P}_{0,\rho}\left((X_{y},\rho_{y})=(j,\sigma)\right)dy\\ =&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}\sum_{\sigma\in\mathcal{D}(\mathfrak{h}),j\in\mathbb{Z}}\int_{0}^{\infty}\mathbb{P}_{j,\sigma}\left(|X_{s}-j|<\varepsilon\right)ds\mathbb{P}_{0,\rho}\left((X_{y},\rho_{y})=(j,\sigma)\right)dy\\ =&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}\sum_{\sigma\in\mathcal{D}}(\mathfrak{h})\int_{0}^{\infty}\mathbb{P}_{0,\sigma}\left(|X_{s}|<\varepsilon\right)ds\sum_{j\in\mathbb{Z}}\mathbb{P}_{0,\rho}\left((X_{y},\rho_{y})=(j,\sigma)\right)dy.\end{split}$
+The inner integral depends only on the internal state $\sigma$ and, since $\mathcal{D}(\mathfrak{h})$ is compact and the integral is continuous in $\sigma$, it attains a finite maximum.
+Replacing it by this maximum - which is independent of $\sigma$ and $j$ - yields the estimate:
+$\begin{split}\zeta^{\rho}_{\varepsilon,k}\leq&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}\sum_{\sigma\in\mathcal{D}(\mathfrak{h})}\max_{\tau\in\mathcal{D}(\mathfrak{h})}\int_{0}^{\infty}\mathbb{P}_{0,\tau}\left(|X_{s}|<\varepsilon\right)ds\sum_{j\in\mathbb{Z}}\mathbb{P}_{0,\rho}\left((X_{y},\rho_{y})=(j,\sigma)\right)dy\\ =&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}\max_{\tau\in\mathcal{D}(\mathfrak{h})}\int_{0}^{\infty}\mathbb{P}_{0,\tau}\left(|X_{s}|<\varepsilon\right)ds\sum_{\sigma\in\mathcal{D}(\mathfrak{h}),j\in\mathbb{Z}}\mathbb{P}_{0,\rho}\left((X_{y},\rho_{y})=(j,\sigma)\right)dy\\ =&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}dy\max_{\tau\in\mathcal{D}(\mathfrak{h})}\int_{0}^{\infty}\mathbb{P}_{0,\tau}\left(|X_{s}|<\varepsilon\right)ds\\ =&\max_{\tau\in\mathcal{D}(\mathfrak{h})}\int_{0}^{\infty}\mathbb{P}_{0,\tau}\left(|X_{s}|<\varepsilon\right)ds.\end{split}$
+Therefore, Equation (11) and the inequality above give
+$\int_{0}^{\infty}\mathbb{P}_{0,\rho}(|X_{t}|<m\varepsilon)dt\leq\sum_{k}\zeta_{\rho}^{k,\varepsilon}\leq\sum_{k}\max_{\tau\in\mathcal{D}(\mathfrak{h})}\int_{0}^{\infty}\mathbb{P}_{0,\tau}(|X_{s}|<\varepsilon)ds=(2m)^{d}\cdot\max_{\tau\in\mathcal{D}(\mathfrak{h})}\int_{0}^{\infty}\mathbb{P}_{0,\tau}(|X_{s}|<\varepsilon)ds.$
+Theorem 4.10.
+If $X_{t}$ denotes a CTOQW on $\mathbb{Z}^{2}$ and $X_{t}/t^{1/2}\Rightarrow$ a Gaussian measure, then the CTOQW is recurrent for some density.
+Proof.
+Let us denote $u(t,m)=\mathbb{P}(|X_{t}|<m).$ By Lemma 4.9,
+$\max_{\tau\in D}\int_{0}^{\infty}\mathbb{P}(|X_{t}|<1)dt=\max_{\tau\in D}\int_{0}^{\infty}u(t,1)dt\geq(4m^{2})^{-1}\int_{0}^{\infty}u(t,m)dt.$
+Assume that $m/\sqrt{t}\rightarrow c,$ then
+$u(t,m)\rightarrow\int_{[-c,c]^{2}}f(x)dx,$ (13)
+where $f(x)$ represents the density of the limiting normal distribution. We denote the right-hand side of Equation (13) by $g(c)$ and let $t=\theta m^{2}$ to obtain
+$u(\theta m^{2},m)\rightarrow g(\theta^{-1/2}).$
+Now we use the change of variables
+$m^{-2}\int_{0}^{\infty}u(t,m)dt=\int_{0}^{\infty}u(\theta m^{2},m)d\theta,$
+let $m\rightarrow\infty,$ and use Fatou’s lemma to get
+$\liminf_{m\rightarrow\infty}(4m^{2})^{-1}\int_{0}^{\infty}u(t,m)dt\geq 4^{-1}\int_{0}^{\infty}g(\theta^{-1/2})d\theta.$ (14)
+Since the normal density is positive and continuous at $0,$
+$g(c)=\int_{[-c,c]^{2}}n(x)dx\sim n(0)(2c)^{2}$
+as $c\rightarrow 0.$ So $g(\theta^{-1/2})\sim 4n(0)/\theta$ as $\theta\rightarrow\infty,$ the integral in (14) diverges.
+As a result, $\max_{\tau\in D}\int_{0}^{\infty}\mathbb{P}(|X_{t}|<1)dt=\infty,$ leading us to conclude that the walk is $\tau$-recurrent.
+$\square$
+Note that the maximum in Lemma 4.9 can be assumed to be taken over the jump times $(T_{n})_{n}.$ Therefore, by Proposition 4.8 and Theorem 4.10 we obtain the following result.
+Theorem 4.11.
+If $X_{t}$ denotes a CTOQW induced by a coin $(A,H),$ there is at most one minimal enclosure for $\Phi$ and $X_{t}/t^{1/2}\Rightarrow$ a Gaussian measure, then the CTOQW is recurrent.
+We already have a sufficient condition to recurrence given by the Theorem 4.11. For transience, we also have a sufficient condition:
+Theorem 4.12.
+Let us consider a CTOQW induced by a coin $(A,H),$ having at most one minimal enclosure, and let $\rho_{\infty}$ be the unique stationary state of the auxiliary map.
+If $m\neq 0,$ then the CTOQW is transient.
+Proof.
+Let us start the CTOQW at some vertex $\lvert i\rangle\in\mathbb{Z}^{2}$ with initial density operator $\rho.$ By the Law of Large Numbers, Proposition 4.5,
+$\frac{X_{t}}{t}\rightarrow m\neq 0^{2}.$
+This means that at least one coordinate of $X_{t}$ is unbounded, from where we conclude that
+$\mathbb{P}\left(X_{t}=X_{0}\;\text{i.o.}\right)=0.$
+Hence, the CTOQW is transient, since this is valid for any $\rho.$
+$\square$
+Finally, we can present a recurrence criteria for CTOQWs on $\mathbb{Z}^{2}$ induced by a coin $(A,H).$
+Theorem 4.13 (Recurrence criteria for CTOQWs on $\mathbb{Z}^{2}$).
+Let us consider a CTOQW induced by a coin $(A,H),$ having at most one minimal enclosure, and let $\rho_{\infty}$ be the unique stationary state of the auxiliary map. Then
+  • $m=0\;\Rightarrow\;$ the CTOQW is recurrent;
+  • $m\neq 0\;\Rightarrow\;$ the CTOQW is transient.
+Proof.
+For $m\neq 0,$ the result is given in Theorem 4.12. It remains to prove that $m=0$ implies that the CTOQW is recurrent. If $m=0,$ then by Theorem 4.4
+$\frac{X_{t}}{\sqrt{t}}=\frac{X_{t}-t.m}{\sqrt{t}}\Rightarrow\text{ a non-degenerate normal distribution,}$
+thus we use Theorem 4.10 to obtain a density operator $\tau$ such that the CTOQW is $\tau$-recurrent. Therefore, $(A,H)$ is recurrent by Theorem 4.11.
+$\square$
+4.5 Recurrence Criterion for HOQWs on the Grid
+We begin with an easily verifiable criterion based on the quantum jump chains associated with CTOQWs, as presented in [22]. Let $\Pi$ denote a discrete HOQW given by
+$\Pi(\rho)=\sum_{i,j\in V}M_{j}^{i}\rho(j)M_{j}^{i*}\otimes\lvert i\rangle\langle i\rvert$.
+This HOQW is the open quantum jump chain of the CTOQW with Lindblad generator
+$\mathbb{L}=\Psi-I,\;\;\Psi(\rho)=\sum_{i,j\in V}B_{j}^{i}\rho(j)B_{j}^{i*},\;B_{j}^{i}=M_{j}^{i}\otimes\lvert i\rangle\langle j\rvert$, (15)
+where $I$ is the identity operator acting on $\mathcal{H}$ [22].
+Theorem 4.14.
+[22, Theorem 5.6] A vertex $i\in V$ is $\rho$-recurrent for the HOQW $\Pi$ if and only if it is $\rho$-recurrent for the CTOQW generated by $\mathbb{L}$ given by Equation (15).
+Theorem 4.14 relies on the open quantum jump chain construction introduced in [22].
+Given a CTOQW with Lindblad generator $\mathbb{L}$, one defines a discrete-time OQW $\Phi$ via $\Phi=\Psi-I$, where $\Psi(\rho)=\sum_{i,j}B_{j}^{i}\rho(j)B_{j}^{i*}$ and $B_{j}^{i}=M_{j}^{i}\otimes|i\rangle\langle j|$ are the jump operators.
+This construction yields a discrete-time quantum walk that records the sequence of jump events and the post-jump quantum states.
+The preservation of recurrence follows from the fact that the jump operators encode the same transitions as the continuous-time process.
+This equivalence holds because the jump times in the continuous-time process are almost surely finite and the sequence of visited sites and internal states coincides with that of the discrete-time walk.
+For a detailed exposition of this construction and the proof of recurrence equivalence, we refer the reader to [22, Section 5].
+We consider an HOQW with a set of vertices $V=\mathbb{Z}^{2}.$ Analogously to the continuous-time version, if there exists $D_{1},D_{2},D_{3},D_{4}\in\mathcal{B}(\mathbb{C}^{n})$ such that
+$D_{1}=M_{(i,j)}^{(i+1,j)},\;D_{2}=M_{(i,j)}^{(i,j+1)},\;D_{3}=M_{(i,j)}^{(i-1,j)},\;D_{4}=M_{(i,j)}^{(i,j-1)},$
+and $M_{x}^{y}=0_{n}$ for the remaining $(x,y)\in\mathbb{Z}^{2},$ then the HOQW $\Pi$ is said to be induced by a coin $(D)$ of dimension $n.$
+The HOQW induced by a coin $(D)$ is the homogeneous, nearest-neighbor HOQW on the grid, with the right, up, left and down transitions given by $D_{1},D_{2},D_{3},D_{4},$ respectively.
+Therefore, the quantum trajectory of those walks, starting from a state $\tau$ of the form $\tau=\sum_{i\in\mathbb{Z}}\rho(i)\otimes|i\rangle\langle i|,$ is any path generated by the Markov chain $(X_{n},\tau_{n})_{n\geq 0}$, where $X_{n}$ denotes the position of the particle at time $n$ and $\tau_{n}$ its internal degree.
+The transition probabilities are given by
+$\begin{split}\mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left((i+1,j),\frac{D_{1}\sigma D_{1}^{*}}{\mathrm{Tr}(D_{1}\sigma D_{1}^{*})}\right)\;\Bigg|\;(X_{n},\rho_{n})=((i,j),\sigma)\,\right)&=\mathrm{Tr}(D_{1}\sigma D_{1}^{*}),\\ \mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left((i,j+1),\frac{D_{2}\sigma D_{2}^{*}}{\mathrm{Tr}(D_{2}\sigma D_{2}^{*})}\right)\;\Bigg|\;(X_{n},\rho_{n})=((i,j),\sigma)\,\right)&=\mathrm{Tr}(D_{2}\sigma D_{2}^{*}),\\ \mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left((i-1,j),\frac{D_{3}\sigma D_{3}^{*}}{\mathrm{Tr}(D_{3}\sigma D_{3}^{*})}\right)\;\Bigg|\;(X_{n},\rho_{n})=((i,j),\sigma)\,\right)&=\mathrm{Tr}(D_{3}\sigma D_{3}^{*}),\\ \mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left((i,j-1),\frac{D_{4}\sigma D_{4}^{*}}{\mathrm{Tr}(D_{4}\sigma D_{4}^{*})}\right)\;\Bigg|\;(X_{n},\rho_{n})=((i,j),\sigma)\,\right)&=\mathrm{Tr}(D_{4}\sigma D_{4}^{*}),\end{split}$
+for every $(i,j)\in\mathbb{Z}^{2}$, $\sigma\in\mathcal{D}(\mathfrak{h})$, and initial law
+$\mathbb{P}\left(\,(X_{0},\rho_{0})=\left((i,j),\frac{\rho(i,j)}{\mathrm{Tr}\,\rho(i,j)}\right)\;\right)=\mathrm{Tr}(\rho(i,j)),$
+and all other transition probabilities are null. The graph of this kind of HOQW is represented in Figure 3.
+$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}\vdots\vdots\vdots\vdots\ldots\vdots\vdots\vdots\vdots$
+Figure 3: Homogeneous Open Quantum Walk on $\mathbb{Z}^{2}$.
+Now we can give a recurrence criteria for HOQWs on $\mathbb{Z}^{2}.$ To this, note that if $\rho$ is an invariant state for the auxiliary map $\Phi$ of $(D),$ defined as
+$\Phi(.)=\sum_{j=1}^{4}D_{j}(.)D_{j}^{*},$
+then it is a stationary state for its open quantum jump chain $\mathbb{L}.$ Consequently, if $\Phi$ has at most one minimal enclosure, then $\Phi$ has a unique invariant state $\rho_{\infty}$ and we can let
+$m=\begin{bmatrix}m_{1}\\ m_{2}\end{bmatrix},\quad m_{1}=\mathrm{Tr}\left(D_{1}\rho_{\infty}D_{1}^{*}\right)-\mathrm{Tr}\left(D_{3}\rho_{\infty}D_{3}^{*}\right),\quad m_{2}=\mathrm{Tr}\left(D_{2}\rho_{\infty}D_{2}^{*}\right)-\mathrm{Tr}\left(D_{4}\rho_{\infty}D_{4}^{*}\right),$
+to obtain, as a consequence of Theorems 4.13 and 4.14, the following criterion, which follows the same structure as Theorem 3.8.
+Theorem 4.15 (Recurrence criterion for HOQWs on $\mathbb{Z}^{2}$).
+Let us consider an HOQW induced by a coin $(D),$ where the map
+$\Phi(.)=\sum_{j=1}^{4}D_{j}(.)D_{j}^{*}$
+has at most one minimal enclosure, and let $\rho_{\infty}$ be the unique invariant state of $\Phi.$ Then
+  • $m=\begin{bmatrix}0\\ 0\end{bmatrix}\;\Rightarrow\;$ the HOQW is recurrent;
+  • $m\neq\begin{bmatrix}0\\ 0\end{bmatrix}\;\Rightarrow\;$ the HOQW is transient.
+The results obtained so far provide recurrence criteria for coins whose associated auxiliary map is ergodic.
+In the next section, we extend these results to more general settings in which the auxiliary map may admit multiple invariant states.
+To achieve this, we partition the state space into suitable invariant components, allowing a refined analysis while preserving the role of invariant states in the construction of recurrence criteria.
+We begin with the case of lazy coins in dimension two, where the low dimensionality enables a complete characterization in terms of eigenvalues and eigenvectors.
+For higher dimensions, we turn our attention to non-lazy coins.
+5 General Recurrence Criteria for HOQWs and CTOQWs
+5.1 General Recurrence Criterion for HOQWs Induced by a Coin $(L,B,R)$ of Dimension 2
 The following theorem is due to [9].
@@
 $\mathcal{L}$ has a minimal invariant subspace, that is, a subspace $S\neq\{0\}$ such that $S\subset S^{\prime}$ for every invariant subspace $S^{\prime}\neq\{0\}.$
-Building upon the established preceding proposition, we will advance to the following theorem, delineating a comprehensive criterion for 2-dimensional coins, extending [15, Theorem 17] to encompass the Lazy case.
+Building upon the established preceding theorem, we will advance to the following result, delineating a comprehensive criterion for 2-dimensional coins, extending [17, Theorem 17] to encompass the Lazy HOQW case.
 This criterion will be structured into two distinct parts: the first adhering to the logic of the previous theorem, and the second articulated in terms of the eigenvalues of the operators of the coin, independent of the invariant densities of the auxiliary map.
@@
 If the elements of the coin do not share at least two orthogonal eigenvectors, then the auxiliary map is ergodic, thus item (1) is a particular case of Theorem 3.8.
-Now we suppose that $L,B,R$ share two linearly independent eigenvectors, then we can apply Theorem 3.9 to obtain $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ common orthonormal eigenvectors of $L,B,R.$ We assume that $|l_{1}|=|r_{1}|$ and $|l_{2}|=|r_{2}|,$ then for $\rho=\lvert u\rangle\langle u\rvert$, where $\lvert u\rangle$ is some common eigenvector of $L,B,R,$ we have $p_{j,j+1;\rho}(n)=p_{j,j-1;\rho}(n)=w$ for all $j\in\mathbb{Z}$ and all $n\geq 0$ for some constant $w$ and therefore, in this case, we have a classical homogeneous walk which is recurrent.
-If $(L,B,R)$ was not transient, then we would be in third situation of proposition 3.5, with
+Now we suppose that $L,B,R$ share two unit linearly independent eigenvectors $\lvert u_{1}\rangle,\lvert u_{2}\rangle.\;$ It is straightforward that $\rho_{1}=\lvert u_{1}\rangle\langle u_{1}\rvert$ and $\rho_{2}=\lvert u_{2}\rangle\langle u_{2}\rvert$ are distinct invariant states of the auxiliary map $\mathcal{L}$, thus $\mathcal{L}$ can not be ergodic.
+By Theorem 5.1, it has two invariant subspaces $S_{1}\neq\{0\}$ and $S_{2}\neq\{0\}$ such that $S_{1}\cap S_{2}=\{0\}.$ Since the coin has dimension 2, two distinct pure invariant states must correspond to orthogonal rank-1 projections.
+Indeed, if $\rho_{1}$ and $\rho_{2}$ are both invariant and $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ are linearly independent, the only way for $\mathcal{L}$ to have both as fixed points is if they are orthogonal.
+Otherwise, any operator that fixes both $\rho_{1}$ and $\rho_{2}$ would fix all operators, forcing $\mathcal{L}$ to be the trivial identity map.
+We assume that $|l_{1}|=|r_{1}|$ and $|l_{2}|=|r_{2}|,$ then for $\rho=\lvert u\rangle\langle u\rvert$, where $\lvert u\rangle$ is some common eigenvector of $L,B,R,$ we have $p_{j,j+1;\rho}(n)=p_{j,j-1;\rho}(n)=w$ for all $j\in\mathbb{Z}$ and all $n\geq 0$ for some constant $w$ and therefore, in this case, we have a classical homogeneous walk which is recurrent.
+If $(L,B,R)$ was not recurrent, then we would be in third situation of proposition 3.5, with
 $\mathfrak{h}=\mathcal{Y}\oplus\mathcal{X},\quad\text{where}\quad\text{dim}(\mathcal{Y})\geq 1,\;\;\text{ and }\;\;\text{dim}(\mathcal{X})=2,$
@@
 $\square$
-4 A General Recurrence Criteria for Open Quantum Walks
-Following the approach described in [10, Section 4], as well as the references therein, we can obtain for a coin $(L,R)$ a decomposition
+5.2 General Recurrence Criterion for HOQWs Induced by Finite Dimensional Coins on the Line
+Following the approach of [11, Section 4] and the references therein, we can obtain for a coin $(L,R)$ a decomposition
 […]
-Since $\mathfrak{h}$ is finite-dimensional, any minimal enclosure is included in $\mathfrak{R},$ and is the support of a unique extremal invariant state.
+Since $\mathfrak{h}$ is finite-dimensional, any minimal enclosure is contained in $\mathfrak{R},$ and is the support of a unique extremal invariant state.
 Moreover, we have a unique decomposition of $\mathfrak{R}$ of the form
-$\mathfrak{R}=\bigoplus_{\alpha\in A}\mathfrak{X}_{\alpha},$
+$\mathfrak{R}=\bigoplus_{\alpha\in A}\mathfrak{X}_{\alpha},$ (16)
 where $\{\mathfrak{X}_{\alpha}\}_{\alpha}\in A$ is a finite set of mutually orthogonal enclosures and every $\mathfrak{X}_{\alpha}$ is minimal in the set of enclosures satisfying
 […]
-Further, every $\mathfrak{X}_{\alpha}$ either is a minimal enclosure or assumes a decomposition (not unique) as the sum of mutually orthogonal isomorphic minimal enclosures $\{\mathcal{Y}_{\alpha,\beta},\;\beta\in I_{\alpha}\}:$
+Further, every $\mathfrak{X}_{\alpha}$ is either a minimal enclosure or admits a decomposition (not unique) as the sum of mutually orthogonal isomorphic minimal enclosures $\{\mathcal{Y}_{\alpha,\beta},\;\beta\in I_{\alpha}\}:$
 $\mathfrak{X}_{\alpha}=\bigoplus_{\beta\in I_{\alpha}}\mathcal{Y}_{\alpha,\beta}.$
-Given a minimal enclosure $\mathcal{Y}_{\beta},$ let $\tau_{\beta}$ be the unique invariant state with support on $\mathcal{Y}_{\beta}.$ Introducing the parameter
+Given a minimal enclosure $\mathcal{Y}_{\beta},$ let $\tau_{\beta}$ be the unique invariant state with support on $\mathcal{Y}_{\beta}.$ We introduce the parameter
 $m_{\beta}=\mathrm{Tr}\left(L\tau_{\beta}L^{*}\right)-\mathrm{Tr}\left(R\tau_{\beta}R^{*}\right),$
-we have by [10, Lemma 4.2] that $m_{\beta}$ is independent of the particular minimal enclosure $\mathcal{Y}_{\beta}$ in $\mathfrak{X}_{\alpha}.$ Therefore, we can combine Proposition 3.5, Theorem 3.8 and the discussion above to state the property
+where $\tau_{\beta}$ is the unique invariant state supported on the minimal enclosure $\mathcal{Y}_{\beta}$.
+Lemma 5.3.
+[11, Lemma 4.2] The parameter $m_{\beta}$ is independent of the particular minimal enclosure $\mathcal{Y}_{\beta}$ in $\mathfrak{X}_{\alpha}$.
+Consequently, $m_{\beta}$ is a characteristic parameter of the component $\mathfrak{X}_{\alpha}$ itself, and we may denote it by $m_{\alpha}$ when referring to the whole component.
+Therefore, we can combine Proposition 3.5, Theorem 3.8 and the discussion above to state the property
 Proposition 5.4.
@@
 where $A$ is a finite set, $\mathcal{Y}_{\alpha}$ is a subspace of $\mathfrak{h}$ being the support of exactly one invariant state $\tau_{\alpha}$ for $\mathcal{L},$ $\mathcal{X}$ contains no enclosure, and
-  • $\textmd{Tr}\left(L\tau_{\alpha}L^{*}\right)=\dfrac{1}{2}\Rightarrow$ the walk is $\tau_{\alpha}$-recurrent for all $\tau\in\mathcal{D}(\mathcal{Y}_{\alpha})$;
-  • $\textmd{Tr}\left(L\tau_{\alpha}L^{*}\right)\neq\dfrac{1}{2}\Rightarrow$ the walk is $\tau_{\alpha}$-transient for all $\tau\in\mathcal{D}(\mathcal{Y}_{\alpha}).$
-Once the decomposition on Equation (5) is obtained, we can say that $\mathcal{Y}_{\alpha}$ is either recurrent or transient.
-The following result is due to [10] and will be crucial to our last criterion.
+  • $\textmd{Tr}\left(L\tau_{\alpha}L^{*}\right)=\dfrac{1}{2}\Rightarrow$ the walk is $\tau$-recurrent for all $\tau\in\mathcal{D}(\mathcal{Y}_{\alpha})$;
+  • $\textmd{Tr}\left(L\tau_{\alpha}L^{*}\right)\neq\dfrac{1}{2}\Rightarrow$ the walk is $\tau$-transient for all $\tau\in\mathcal{D}(\mathcal{Y}_{\alpha}).$
+Once the decomposition in Equation (17) is obtained, we can say that $\mathcal{Y}_{\alpha}$ is either recurrent or transient.
+The following result is due to [11] and will be crucial to our next criterion.
 To state it, consider a coin $(L,R),$ and let $\mathcal{Y}$ be an enclosure for $\mathcal{L}.$ Denoting $p_{\mathcal{Y}}$ the orthogonal projection onto $\mathcal{Y},$ we can define the associated absorption operator (see [10]):
@@
 Corollary 5.5.
-Consider a coin $(L,R)$ with the decomposition of $\mathfrak{h}$ given on Equation (5), and let $a_{\alpha}(\rho)=\mathbb{E}_{\rho}\left[\mathrm{Tr}\left(A(\mathcal{Y}_{\alpha})\rho_{0}\right)\right].$ Then we have the convergence in law
+[11, Corollary 4.5] Consider a coin $(L,R)$ with the decomposition of $\mathfrak{h}$ given in Equation (17), and let $a_{\alpha}(\rho)=\mathbb{E}_{\rho}\left[\mathrm{Tr}\left(A(\mathcal{Y}_{\alpha})\rho_{0}\right)\right].$ Then we have the convergence in law
 $\frac{X_{n}-X_{0}}{n}\stackrel{{\scriptstyle n\rightarrow\infty}}{{\longrightarrow}}\sum_{\alpha\in A}a_{\alpha}(\rho)\delta_{m_{\alpha}},$ (18)
 where $\delta_{m_{\alpha}}$ is the Dirac measure concentrated in $m_{\alpha}.$
-We remark that the convergence in law brought by corollary 4.2 can be upgraded to a convergence a.s., as demonstrated in [13, Theorem 3.5.2].
+We remark that the convergence in law brought by Corollary 5.5 can be upgraded to a convergence a.s., as demonstrated in [15, Theorem 3.5.2].
 Let us start an HOQW induced by a coin $(L,R).$ There are two possibilities concerning recurrence when $\text{supp}(\rho_{0})\not\perp\mathcal{X}:$
@@
 $(3)\quad\mathrm{Tr}(L\tau_{\alpha}L^{*})=\dfrac{1}{2}\;$ for every $\alpha\Rightarrow$ the walk is recurrent.
-Remark $4.$
+Remark 5.7.
 […]
-5 Examples - 1D
+5.3 A Generalized Recurrence Criterion for HOQWs on the Grid
+We have already presented some recurrence criteria for discrete-time and continuous-time OQWs on the set of vertices $\mathbb{Z}^{2}$, constrained by a minimal enclosure.
+These criteria are interesting tools for understanding the long-term behavior of OQWs.
+Building on these consequences, we aim to derive a stronger recurrence criterion for discrete-time HOQWs, allowing for a more comprehensive analysis of their statistics.
+At this point we will allow the auxiliary map of a coin $(D)$ to have more than one invariant state, thus we consider the decomposition of $\mathfrak{h}$ given on Section 5.2.
+Therefore, given a minimal enclosure $\mathcal{Y}_{\alpha},$ let $\tau_{\alpha}$ be the unique invariant state with support on $\mathcal{Y}_{\alpha}.$ Consider the parameter
+$\tilde{m}_{\alpha}=\begin{bmatrix}m_{1,\alpha}\\ m_{2,\alpha}\end{bmatrix},\quad m_{1,\alpha}=\mathrm{Tr}\left(D_{1}\tau_{\alpha}D_{1}^{*}\right)-\mathrm{Tr}\left(D_{3}\tau_{\alpha}D_{3}^{*}\right),\quad m_{2,\alpha}=\mathrm{Tr}\left(D_{2}\tau_{\alpha}D_{2}^{*}\right)-\mathrm{Tr}\left(D_{4}\tau_{\alpha}D_{4}^{*}\right).$
+As stated in Lemma 5.3, the value $\tilde{m}_{\alpha}$ is independent of the particular minimal enclosure $\mathcal{Y}_{\alpha}$ in $\mathfrak{X}_{\alpha}.$ Therefore, we can combine Theorem 4.15 and results concerning CTOQWs above to state the next property, which is a continuous-time version on the grid of Proposition 5.4.
+Proposition 5.8.
+Consider a coin $(D).$ We have a decomposition
+$\mathfrak{h}=\left(\bigoplus_{\alpha\in A}\mathcal{Y}_{\alpha}\right)\oplus\mathcal{X},$ (19)
+where $A$ is a finite set, $\mathcal{Y}_{\alpha}$ is a subspace of $\mathfrak{h}$ being the support of exactly one invariant state $\tau_{\alpha}$ for $\Phi,$ $\mathcal{X}$ contains no enclosure, and
+  • $\tilde{m}_{\alpha}=0\Rightarrow$ the walk is $\tau_{\alpha}$-recurrent for all $\tau\in\mathcal{D}(\mathcal{Y}_{\alpha})$;
+  • $\tilde{m}_{\alpha}\neq 0\Rightarrow$ the walk is $\tau_{\alpha}$-transient for all $\tau\in\mathcal{D}(\mathcal{Y}_{\alpha}).$
+Proof.
+Fixed a subspace $\mathcal{Y}_{\alpha}$ which is the support of exactly one invariant state $\tau_{\alpha}\in\mathcal{D}(\mathfrak{h}),$ we assume that $\tilde{m}_{\alpha}=0.$ Then $(D)$ is $\rho$-recurrent for every $\rho$ by Theorem 4.11.
+If $\tilde{m}_{\alpha}\neq 0,$ then $(D)$ is transient by Theorem 4.12.
+$\square$
+With the decomposition in Equation (19) established, we can say that $\mathcal{Y}_{\alpha}$ is either recurrent or transient.
+It remains to understand what happens when a quantum state has support outside of $\oplus_{\alpha}\mathcal{Y}_{\alpha}.$
+Analogously to the one dimensional case, our recurrence criterion is:
+Theorem 5.9 (A Generalized Criterion for coins $(D)$).
+Consider a coin $(D)$ of dimension $d<\infty.$ Consider the decomposition of $\mathfrak{h}$ as given in Equation (19), and $\tau_{1},\ldots,\tau_{w}$ the respectively invariant states of $\Phi$ in $\mathcal{Y}_{\alpha},\;\alpha=1,\ldots,w.$
+$(1)\quad\tilde{m}_{\alpha}\neq 0\;\forall j\in\{1,\ldots w\}\Rightarrow(D)\mbox{ is transient;}$
+[…]
+[…]
+$\begin{cases}\rho\text{-transient for }\rho\in\mathcal{D}_{1}:=\mathcal{D}(\oplus_{k\in J^{C}}\mathcal{Y}_{k})\\ \rho\text{-recurrent for }\rho\in\mathcal{D}_{1}^{C}.\end{cases}$
+[…]
+$\begin{cases}\rho\text{-transient for }\rho\in\mathcal{D}_{2}:=\mathcal{D}(\oplus_{k\in J^{C}}\mathcal{Y}_{k}\oplus\mathcal{X})\\ \rho\text{-recurrent for }\mathcal{D}_{2}^{C}\end{cases}$
+$(3)\quad\tilde{m}_{\alpha}=0\;$ for every $\alpha\Rightarrow$ the walk is recurrent.
+For quick reference, Table 2 gathers the notation for the density sets and operator spaces employed in the paper.
+Notation Description
+$\mathfrak{h}$ Internal space
+$\mathcal{D}(\mathfrak{h})$ Set of density operators on the internal space $\mathfrak{h}$
+$\mathcal{D}_{i}(\mathfrak{h})$ Set of density operators on the internal space $\mathfrak{h}_{i}$
+$\mathcal{S}(\mathcal{H}\otimes\mathcal{K})$ Set of density operators on $\mathcal{H}\otimes\mathcal{K}$ of diagonal form with respect to position
+$\mathcal{D}(\mathcal{Y})$ Set of density operators supported in a subspace $\mathcal{Y}\subset\mathfrak{h}$
+$\mathcal{D}_{1}$, $\mathcal{D}_{2}$ Sets of transient densities in Theorems 5.6 and 5.9
+$\mathcal{I}_{1}(\mathcal{H})$ Trace-class operators on $\mathcal{H}$
+$\mathcal{B}(\mathcal{W})$ Bounded linear operators on a Hilbert space $\mathcal{W}$
+Table 2: Notation for operator sets and density spaces.
+Having established rigorous recurrence criteria for HOQWs and CTOQWs on the line and on the grid, which classify the behavior based on the decomposition of the internal space $\mathfrak{h}$ and the drift parameters $\tilde{m}_{\alpha}$, we now illustrate these concepts with explicit numerical matrices.
+The following examples are also designed to demonstrate how the decomposition into enclosures are performed in practice.
+6 Examples
+6.1 1D
 Example 6.1.
@@
     by Item (1) of Theorem 5.2.
-    It is worth noting that if we assume $m=0,$ then the walk is transient by Item (2) of Theorem 3.10.
-More explicitly, in this case, for the item (2.2), $|l_{1}|=0,\;|l_{2}|=n,\;|r_{1}|=n,\;|r_{2}|=0.$
+    Note that for $m=0$, the operators $L,B,R$ have the common and linearly independent eigenvectors $1,0]^{T}$ and $0,1]^{T}$.
+According to item (2.2) of Theorem 5.2, with $|l_{1}|=0$, $|l_{2}|=n$, $|r_{1}|=n$, $|r_{2}|=0$, the walk is transient.
     $\Diamond$
@@
 $L=\frac{1}{8}\begin{bmatrix}2(1+\sqrt{2})&0&2(1-\sqrt{2})\\ 0&\sqrt{31}&0\\ 2(1-\sqrt{2})&0&2(1+\sqrt{2})\end{bmatrix},\quad B=\frac{1}{16}\begin{bmatrix}\sqrt{30}&2i&\sqrt{30}\\ 2&0&2\\ \sqrt{30}&-2i&\sqrt{30}\end{bmatrix},\quad R=\frac{1}{8}\begin{bmatrix}2(1-\sqrt{2})&0&2(1+\sqrt{2})\\ 0&\sqrt{31}&0\\ 2(1+\sqrt{2})&0&2(1-\sqrt{2})\end{bmatrix}$.
-There is a unique invariant state $\rho_{\infty}$ given by
+There is a unique invariant state $\rho_{\infty}$ for the auxiliary map given by
 $\rho_{\infty}=\frac{1}{2}\begin{bmatrix}1&0&-1\\ 0&0&0\\ -1&0&1\end{bmatrix}$,
@@
 $L=\frac{1}{7}\begin{bmatrix}1&0\\ 0&2\end{bmatrix},\qquad B=\frac{1}{7}\begin{bmatrix}-1&1\\ 2&0\end{bmatrix},\qquad R=\frac{1}{7}\begin{bmatrix}1&1\\ 0&1\end{bmatrix}$.
-Then
-$\rho_{\infty}=\frac{1}{3}\begin{bmatrix}1&0\\ 0&2\end{bmatrix}$,
-$\mathrm{Tr}\left(L\rho_{\infty}L^{*}\right)=3/7$ and $\mathrm{Tr}\left(R\rho_{\infty}R^{*}\right)=5/21.$ The coin $(L,B,R)$ is transient. However, if we take
+Again, there is a unique invariant state $\rho_{\infty}$ for the auxiliary map $\mathcal{L}$ given by
+$\rho_{\infty}=\frac{1}{3}\begin{bmatrix}1&0\\ 0&2\end{bmatrix}$.
+In this case, $\mathrm{Tr}\left(L\rho_{\infty}L^{*}\right)=3/7$ and $\mathrm{Tr}\left(R\rho_{\infty}R^{*}\right)=5/21,$ thus the coin $(L,B,R)$ is transient by item (1) of Theorem 5.2.
+However, if we take
 $B^{\prime}=\frac{1}{\sqrt{35}}\begin{bmatrix}5&-1\\ 0&2\end{bmatrix}$,
-then the invariant state is
+then the unique invariant state for the auxiliary map $\mathcal{L}^{\prime}$ is
 $\rho^{\prime}_{\infty}=\begin{bmatrix}1&0\\ 0&0\end{bmatrix}$,
-and $\mathrm{Tr}\left(L\rho_{\infty}L^{*}\right)=1/\sqrt{7}=\mathrm{Tr}\left(R\rho_{\infty}R^{*}\right).$ Thus the coin $(L,B^{\prime},R)$ is recurrent.
+and $\mathrm{Tr}\left(L\rho_{\infty}L^{*}\right)=1/\sqrt{7}=\mathrm{Tr}\left(R\rho_{\infty}R^{*}\right).$ Thus the coin $(L,B^{\prime},R)$ is recurrent, also by item (1) of Theorem 5.2
 $\Diamond$
-We remark that the choices of the coins above were taken in such a way that $L,B^{\prime}$ and $R$ had a common invariant eigenvector, namely $1,0]^{T},$ which was not an eigenvector of $B.$ This is due to the hard job of, given non-trivial $R$ and $L,$ to find $B^{\prime}$ in such a way that $(L,B^{\prime},R)$ is recurrent.
-Attempt were conduct using mathematic software with many $L$’s and $R$’s having simple entries, but even though the entries of $B$’s that made $(L,B,R)$ to be transient were very long to write in this paper.
+We remark that this choice was possible because $L$ and $R$ share exactly one common eigenvector, namely $\lvert v\rangle=[1,0]^{T}$, and satisfy $L\lvert v\rangle=R\lvert v\rangle$.
+Under these conditions, $\mathrm{Tr}(L\rho L^{*})=\mathrm{Tr}(R\rho R^{*})$ holds for $\rho=\lvert v\rangle\langle v\rvert$.
+Hence, recurrence could be achieved by finding a matrix $B^{\prime}$ such that:
+  1. 1.
+    $\lvert v\rangle$ is also an eigenvector of $B^{\prime},$ thus $\rho$ becomes the unique invariant state for $\mathcal{L}^{\prime}$;
+  2. 2.
+    the normalization condition $L^{*}L+B^{\prime*}B^{\prime}+R^{*}R=I$ is satisfied.
+The explicit matrix $B^{\prime}$ presented in the example was obtained by solving these constraints.
+It is worth noting that finding such $B^{\prime}$ for more general $L$ and $R$ is considerably more difficult.
+Several attempts were made using mathematical software (Maple) with various choices of $L$ and $R$ having simple entries; however, the resulting matrices $B^{\prime}$ that yielded recurrence often had very lengthy expressions, making them impractical to display in this paper.
+Another natural continuation of this work would be to investigate under what conditions on $L$ and $R$ there exists a matrix $B^{\prime}$ such that $(L,B^{\prime},R)$ is recurrent, and to develop methods for explicitly constructing such $B^{\prime}$.
+Given fixed $L,R$, one would like to determine whether the set of admissible $B$ (satisfying $L^{*}L+B^{*}B+R^{*}R=I$) contains recurrent elements, and to describe them explicitly.
+As seen in this work, this problem is closely tied to the invariant state $\rho_{\infty}$ of $\mathcal{L}$ and the condition $\mathrm{Tr}(L\rho_{\infty}L^{*})=\mathrm{Tr}(R\rho_{\infty}R^{*})$.
 Example 6.4.
-Let us consider the coin given in [10, Section 6.2]:
+Let us consider the coin given in [11, Section 6.2]:
 $R=\begin{bmatrix}\sqrt{\frac{3}{8}}&0&0&0\\ -\sqrt{\frac{p_{1}}{2}}&\frac{1}{\sqrt{2}}&0&0\\ -\sqrt{\frac{p_{2}}{2}}&0&\frac{1}{\sqrt{2}}&0\\ \sqrt{\frac{2p_{3}}{3}}&0&0&\frac{1}{\sqrt{3}}\end{bmatrix},\qquad L=\begin{bmatrix}\frac{1}{2\sqrt{2}}&0&0&0\\ \sqrt{\frac{p_{1}}{2}}&\frac{1}{\sqrt{2}}&0&0\\ \sqrt{\frac{p_{2}}{2}}&0&\frac{1}{\sqrt{2}}&0\\ -\sqrt{\frac{p_{3}}{3}}&0&0&\frac{\sqrt{2}}{\sqrt{3}}\end{bmatrix}$,
@@
 By item (3) of Theorem 5.6, the walk is recurrent.
-6 2D Open Quantum Walks
-We start this section with the continuous-time OQW.
-6.1 Continuous-Time Open Quantum Walks
-In order to define the CTOQWs, we recall that an operator semigroup $\mathcal{T}$ on a Hilbert space $\mathcal{K}$ is a family of bounded linear operators $(T_{t})$ acting on $\mathcal{K}$, $t\geq 0,$ such that
-$T_{t}T_{s}=T_{t+s},\;\forall s,t\in\mathbb{R}^{+}\;\mbox{ and }\;T_{0}=I_{\mathcal{K}}.$
-If $t\mapsto T_{t}$ is continuous for the operator norm of $\mathcal{K}$, then $\mathcal{T}$ is said to be uniformly continuous, which is equivalent to the following assertion [7]: There exists a bounded linear operator $L$ on $\mathcal{K}$ such that $T_{t}=e^{tL},\;\forall t\geq 0.$ In this case,
-$L=\lim_{t\rightarrow\infty}\frac{1}{t}(T_{t}-I_{\mathcal{K}})$
-and the operator $L$ is called the generator of $\mathcal{T}.$
-[…]
-Definition 6.1.
-Let $V$ be a finite or countable infinite set and $\mathcal{H}$ be a Hilbert space of the form (2).
-A Continuous-time Open Quantum Walk in $V$ is an uniformly continuous QMS on $\mathcal{I}_{1}(\mathcal{H})$ with Lindblad operator of the form
-$\begin{split}\mathcal{\mathcal{L}}:\mathcal{I}_{1}(\mathcal{H})&\rightarrow\mathcal{I}_{1}(\mathcal{H})\\ \rho&\mapsto-i[\textbf{H},\rho]+\sum_{i,j\in V}\left(S_{i}^{j}\rho S_{i}^{j^{*}}-\frac{1}{2}\{S_{i}^{j*}S_{i}^{j},\rho\}\right),\end{split}$ (7)
-where $A,B]\equiv AB-BA$ and $\{A,B\}\equiv AB+BA.$
-To rearrange Equation 7 We can write $S_{i}^{j}=R_{i}^{j}\otimes\lvert j\rangle\langle i\rvert$ for bounded operators $R_{i}^{j}\in\mathcal{B}(\mathfrak{h}_{i},\mathfrak{h}_{j}),$ the bounded operators H and $S_{i}^{j}$ on $\mathcal{H}$ are of the form $\textbf{H}=\sum_{i\in V}H_{i}\otimes\lvert i\rangle\langle i\rvert$, $H_{i}$ is self-adjoint on $\mathfrak{h}_{i},$ and $\sum_{i,j\in V}S_{i}^{j*}S_{i}^{j}$ converges in the strong sense.
-Letting $\rho=\sum_{i\in V}\rho(i)\otimes\lvert i\rangle\langle i\rvert\in\mathcal{D},$ denoting
-$e^{t\mathcal{L}}(\rho)=\mathcal{T}_{t}(\rho)=\sum_{i\in V}\rho_{t}(i)\otimes\lvert i\rangle\langle i\rvert,\forall t\geq 0,$
-we have
-$\frac{d}{dt}\rho_{t}(i)=-i[H_{i},\rho_{t}(i)]+\sum_{j\in V}\left(R_{j}^{i}\rho_{t}(j)R_{j}^{i^{*}}-\frac{1}{2}\{R_{i}^{j*}R_{i}^{j},\rho_{t}(i)\}\right).$
-Equation (7) can be written as given in [3, Equation 18.7]:
-$\mathcal{L}(\rho)=\sum_{i\in V}\left(G_{i}\rho(i)+\rho(i)G_{i}^{*}+\sum_{j\in V}R_{j}^{i}\rho(j)R_{j}^{i*}\right)\otimes\lvert i\rangle\langle i\rvert,\qquad G_{i}=-iH_{i}-\frac{1}{2}\sum_{j\in V}R_{i}^{j*}R_{i}^{j}.$ (8)
-In this paper we deal with the quantum trajectories of CTOQWs, which are more complex than the trajectories od discrete-time HOQWs.
-Thus, we recall that for a givenfinite or countably infinite set of vertices $V,$ a CTOQW is a stochastic process evolving on a Hilbert space of the form (2), the label $i\in V$ represents the position of the walker and when the walker is located at $i\in V,$ its internal state is encoded in $\mathfrak{h}_{i},$ that is, $\mathfrak{h}_{i}$ describes the internal degrees of freedom of the walker when it is at site $i\in V.$
-Given $i\in V,$ we denote the set of density operators concentrated at vertex $i$ by
-$\mathcal{D}_{i}=\{\rho\in\mathcal{I}(\mathfrak{h}_{i}),\;\rho\geq 0,\;\mathrm{Tr}(\rho)=1\}.$
-Thus, for $\rho\in\mathcal{D}_{i},$ we have $\rho\otimes\lvert i\rangle\langle i\rvert\in\mathcal{D}.$
-Starting the walk on site $\lvert i\rangle$ with initial density operator $\rho\in\mathcal{D}_{i},$ the quantum measurement of the “position” gives rise to a probability distribution $p_{0}$ on $V,$ such that
-$p_{0}(j)=\mathbb{P}(\mbox{the quantum particle is in site}\;\lvert j\rangle)=\mathrm{Tr}(\rho(j)),$
-and for evolution on time $t\geq 0,$
-[…]
-where
-$e^{t\mathcal{L}}(\rho)=\sum_{k\in V}\rho_{t}(i)\otimes\lvert k\rangle\langle k\rvert$.
-6.2 Quantum Trajectories
-Let $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$ be a probability space where independent Poisson point processes $N^{ij},i,j\in V,\;i\neq j$ ($N^{ii}=0$ by convention) on $\mathbb{R}^{2}$ are defined.
-The jump from site $i$ to site $j$ on the graph $V$ will be governed by these Poisson point processes [3]:
-Definition 6.2.
-Consider a CTOQW with generator of the form (8) and an initial density operator $\mu=\sum_{i\in V}\rho(i)\otimes\lvert i\rangle\langle i\rvert\in\mathcal{D}.$ The quantum trajectory describing the indirect measurement of the position of the CTOQW is the Markov chain represented by the density operators $(\mu_{t})_{t\geq 0}$ such that $\mu_{0}=\rho_{0}\otimes\lvert X_{0}\rangle\langle X_{0}\rvert$, where $X_{0}$ and $\rho_{0}$ are random variables with distribution
-$\mathbb{P}\left((X_{0},\rho)=\left(i,\frac{\rho(i)}{\mathrm{Tr}(\rho(i))}\right)\right)=\mathrm{Tr}\left(\rho(i)\right)\mbox{ for all }i\in V,$
-and $\mu_{t}=:\rho_{t}\otimes\lvert X_{t}\rangle\langle X_{t}\rvert$ satisfies the stochastic differential equation
-$\begin{split}\mu_{t}=&\mu_{0}+\int_{0}^{t}M(\mu_{s^{-}})ds\\ +&\sum_{ij}\int_{0}^{t}\int_{\mathbb{R}}\left(\frac{S_{i}^{j}\mu_{s^{-}}S_{i}^{j*}}{\mathrm{Tr}(SS_{i}^{j*}\mu_{s^{-}}S_{i}^{j*})}-\mu_{s^{-}}\right)1_{0<y<\mathrm{Tr}(S_{i}^{j}\mu_{s^{-}}S_{i}^{j*})}N^{ij}(dy,ds)\end{split}$ (10)
-for all $t\geq 0,$ where
-$M(u)=\mathcal{L}(u)-\sum_{ij}\left(S_{i}^{j}\mu S_{i}^{j*}-\mu\mathrm{Tr}(S_{i}^{j}\mu S_{i}^{j*})\right).$
-Hence, for a fixed $\mu=\sum_{i}\rho(i)\otimes\lvert i\rangle\langle i\rvert\in\mathcal{D},$
-$M(\mu)=\sum_{i}\left(G_{i}\rho(i)+\rho(i)G_{i}^{*}-\rho(i)\mathrm{Tr}\left(G_{i}\rho(i)+\rho(i)G_{i}^{*}\right)\right)\otimes\lvert i\rangle\langle i\rvert$.
-For simplicity, suppose $X_{0}=i_{0}$ for some $i_{0}\in V$ and $\rho_{0}\in V(\mathfrak{h}_{i_{0}}).$ For all $t\geq 0,$ consider the solution of Equation (10)
-$\eta_{t}=\rho_{0}+\int_{0}^{t}\left(G_{i_{0}}\eta_{s}+\eta_{s}G_{i_{0}}^{*}-\eta_{s}\mathrm{Tr}\left(G_{i_{0}}\eta_{s}+\eta_{s}G_{i_{0}}^{*}\right)\right)ds,$
-which is a density operator on $\mathfrak{h}_{i_{0}}.$ For $j\neq i_{0},$ define
-$T_{1}^{j}=\mathrm{inf}\{t\geq 0;N^{i_{0},j}\left({u,y|0\leq u\leq t,0\leq y\leq\mathrm{Tr}(R_{i_{0}}^{j}\eta_{u}R_{i_{0}}^{j*})}\right)\geq 1\}.$
-Since the random variables $T_{1}^{j}$ are mutually independent and nonatomic, we can define $T_{1}=\inf_{j\neq i_{0}}\{T_{1}^{j}\}$ once there exists a unique $j\in V$ such that $T_{1}^{j}=T_{1}.$ The random variable $T_{1}$ is said to be the first jump time of the CTOQW conditional on $X_{0}=i_{0}.$
-The first jump time to site $\lvert j\rangle$ is then denoted by $T_{1}^{j}$ and has distribution
-$\mathbb{P}(T_{1}^{j}>\varepsilon)=e^{-\int_{0}^{\varepsilon}\mathrm{Tr}(R_{i_{0}}^{j}\eta_{u}R_{i_{0}}^{j*})du},$
-thus
-$\mathbb{P}(T_{1}\leq\varepsilon)\leq\varepsilon\sum_{j\neq i_{0}}\|R_{i_{0}}^{j*}R_{i_{0}}^{j}\|.$
-The strongly convergence of $\sum_{ij}S_{i}^{j*}S_{i}^{j}$ implies that $\mathbb{P}(T_{1}>0)=1.$ Thereby, on $0,T_{1}],$ we can define the solution $(X_{t},\rho_{t})_{t\geq 0}$ as
-$(X_{t},\rho_{t})$ $=$ $(i_{0},\eta_{t})\mbox{ for }t\in[0,T_{1})\mbox{ and }$
-$(X_{T_{1}},\rho_{T_{1}})$ $=$ $\left(j,\frac{R_{i}^{j}\eta_{T_{1}-}R_{i}^{j*}}{\mathrm{Tr}(R_{i}^{j}\eta_{T_{1}-}R_{i}^{j*})}\right)\;\mbox{ if }T_{1}=T_{1}^{j}.$
-Now we solve
-$\eta_{t}=\rho_{T_{1}}+\int_{0}^{t}\left(G_{j}\eta_{s}+\eta_{s}G_{j}^{*}-\eta_{s}\mathrm{Tr}\left(G_{j}\eta_{s}+\eta_{s}G_{j}^{*}\right)\right)ds$
-and then obtain the second jump time $T_{2}.$ So on we obtain an increasing sequence of jumps $(T_{n})_{n}$ with $\lim_{n\rightarrow\infty}T_{n}=\infty$ almost surely (see section 18.2.3 of [3] for more details).
-This means that the walk does not explode, thus the walker has a finite number of jumps in a finite interval.
-For details concerning explosions of classical Markov chains, see [21, Section 2.2].
-6.3 Recurrence of CTOQWs
-Movin forward, we deal with the following class of 2-dimensional CTOQWs, that is, we will consider the nearest-neighbor quantum random walk on $V=\mathbb{Z}^{2}.$
-Definition 6.3.
-Consider a CTOQW on $V=\mathbb{Z}^{2}$. If there exists $A_{1},A_{2},A_{3},A_{4},H\in\mathcal{B}(\mathbb{C}^{n})$ such that
-$H_{k}=H\;\forall k\in\mathbb{Z}^{2},\quad A_{1}=R_{(i,j)}^{(i+1,j)},\;A_{2}=R_{(i,j)}^{(i,j+1)},\;A_{3}=R_{(i,j)}^{(i-1,j+1)},\;A_{4}=R_{(i,j)}^{(i,j-1)},$
-and $R_{x}^{y}=0_{n}$ for the remaining $x,y\in\mathbb{Z}^{2},$ then the CTOQW is said to be induced by a coin $(A,H)$ of dimension $n.$
-For this kind of CTOQW we can let $\mathfrak{h}_{i}=\mathfrak{h}$ for all $i\in\mathbb{Z}^{2}.$ If $\dim(\mathfrak{h})=n,$ then $A_{1},A_{2},A_{3},A_{4}$ and $H$ can be represented by square matrices of order $n.$
-The auxiliary map of a coin $(A,H)$ is defined as
-$\mathbb{L}(\rho)=G\rho+\rho G^{*}+\sum_{j=1}^{4}A_{j}\rho A_{j}^{*},\qquad G=-iH-\frac{1}{2}\sum_{k=1}^{4}A_{k}^{*}A_{k}.$
-The map $\mathbb{L}$ has at least one stationary state $\rho_{\infty},$ that is, $\mathbb{L}(\rho_{\infty})=0$ [5]. When the stationary state is unique, we define
-$m=\begin{bmatrix}m_{1}\\ m_{2}\end{bmatrix},\quad m_{1}=\mathrm{Tr}\left(A_{1}\rho_{\infty}A_{1}^{*}\right)-\mathrm{Tr}\left(A_{3}\rho_{\infty}A_{3}^{*}\right),\quad m_{2}=\mathrm{Tr}\left(A_{2}\rho_{\infty}A_{2}^{*}\right)-\mathrm{Tr}\left(A_{4}\rho_{\infty}A_{4}^{*}\right).$
-To demonstrate the recurrence criteria of this work for CTOQWs, we recall the following statistical results for those random walks.
-Theorem 6.4 ([6]).
-Consider a coin $(A,H)$, where its auxiliary map has a unique invariant state $\rho_{\infty}.$ Then
-$\frac{X_{t}-tm}{\sqrt{t}}\stackrel{{\scriptstyle t\rightarrow\infty}}{{\longrightarrow}}N(0,\sigma^{2})$ (11)
-for some $\sigma>0.$
-Proposition 6.5 (Law of Large Numbers for CTOQWs ).
-[19, Theorem 15] If the auxiliary map of a coin $(D)$ has a unique stationary state, then
-$\frac{X_{t}}{t}\rightarrow m\quad\mbox{a.s}.$
-Let us recall the classical definition of recurrence of CTOQWs.
-For this, $p_{ji;\rho}(t)$ will denote the probability of being at site $j$ at time $t,$ given that we started at site $i,$ with initial density $\rho$ concentrated at $i$:
-[…]
-Therefore, the walk starts with a density operator $\rho$ concentrated at some vertex $\lvert i\rangle,$ takes the evolution up to time $t>0$ through the quantum Markov semigroup generated by $\mathcal{L},$ producing a new density operator
-$\rho_{t}=\sum_{k}\rho_{t}(k)\otimes\lvert k\rangle\langle k\rvert=e^{t\mathcal{L}}(\rho\otimes\lvert i\rangle\langle i\rvert),\quad\mathrm{Tr}\left(\sum_{k}\rho_{t}(k)\right)=1.$
-Now we project $\rho_{t}$ onto the subspace generated by vertex $\lvert j\rangle$ giving
-$e^{t\mathcal{L}}(\rho\otimes\lvert i\rangle\langle i\rvert)(I\otimes\lvert j\rangle\langle j\rvert)=\rho_{t}(j)\otimes\lvert j\rangle\langle j\rvert$,
-which represents the data concentrated at vertex $\lvert j\rangle$ at time $t.$
-Let $i\in V,\rho\in\mathcal{D}_{i}.$ We say that vertex $i$ is
-  • $\rho$-recurrent if
-    $\int_{0}^{\infty}p_{ii;\rho}(t)dt=\infty.$
-    Otherwise, $i$ is said to be $\rho$-transient
-  • recurrent, if $i$ is $\rho$-recurrent for all $\rho\in\mathcal{D}_{i};$
-  • transient, if $i$ is $\rho$-transient for all $\rho\in\mathcal{D}_{i}.$
-By definition, $i$ is $\rho$-recurrent when the mean number of returns to $i$ is infinite.
-Definition 6.6.
-A CTOQW is said to be:
-  • recurrent if every vertex is recurrent;
-  • transient if every vertex is transient.
-Proposition 6.7.
-A CTOQW induced by a coin is $\rho$-recurrent if and only if it is $\rho_{T_{n}}$-recurrent for any possible jump time $T_{n}.$
-Proof.
-Note that if we start the walk with $(X_{0},\rho_{0})=(i,\rho),$ then at some time $t>0,$ we have $(X_{t},\rho_{t})=(j,\rho_{T}).$ If we refresh the walk at this instant, then the walk is still recurrent with respect to the density $\rho_{t},$ by the homogeneity in space.
-The other hand follows the same idea, assuming by contraposition that the walk is $\rho$-transient.
-$\square$
-Proposition 6.8.
-If $(A,H)$ is $\tau$-recurrent for some density $\tau\in\mathcal{D}$ and $\mathfrak{h}$ has at most one non-trivial invariant subspace for $\Phi:\mathfrak{h}\rightarrow\mathfrak{h}$ given by
-$\Phi(.)=\sum_{k}A_{k}(.)A_{k}^{*},$
-then $(A,H)$ is recurrent.
-Proof.
-Let us start supposing that $(A,H)$ is recurrent with respect to some density operator $\tau$ with support on $\mathcal{V}.$ Let $\rho\in\mathcal{D}$ be another density operator.
-Any $\rho_{T_{n}}$ is of the form
-$\rho_{T_{n}}=\frac{A_{k}\eta A_{k}^{*}}{\mathrm{Tr}\left(A_{k}\eta A_{k}^{*}\right)}\text{ for some }k\in\{1,2,3,4\},$ (12)
-[…]
-$\sigma:=\rho_{T_{n}}-\frac{\tau}{a}\geq 0.$
-Therefore,
-$\frac{1}{\mathrm{Tr}(\sigma)}\int_{0}^{\infty}p_{ii;\rho_{T_{n}}}(t)\;\text{d}t=\int_{0}^{\infty}p_{ii;\frac{\sigma}{\mathrm{Tr}(\sigma)}}(t)\;\text{d}t+\frac{1}{a\mathrm{Tr}(\sigma)}\int_{0}^{\infty}p_{ii;\tau}(t)\;\text{d}t=\infty.$
-By Proposition 6.7, $(A,H)$ is $\rho$-recurrent.
-$\square$
-6.4 Recurrence Criterion
-This section is devoted to exhibit a recurrence criterion for finite dimensional coins $(A,H),$ where the auxiliary map has at most one common invariant subspace.
-The result uses a Central Limit Theorem for CTOQWs and the following Lemma.
-Lemma 6.9.
-Let us consider a homogeneous CTOQW on $\mathbb{Z}^{d}.$ Let $\varepsilon>0$ and $m\geq 1$ be an integer. Then, for every $\rho\in\mathcal{D}$,
-$\int_{0}^{\infty}\mathbb{P}_{0,\rho}(|X_{t}|<m\varepsilon)dt\leq(2m)^{d}\cdot\max_{\sigma\in\mathcal{D}}\int_{0}^{\infty}\mathbb{P}_{0,\sigma}(|X_{t}|<\varepsilon)dt.$
-Proof. Firstly, note that
-$\int_{0}^{\infty}\mathbb{P}_{0,\rho}\left(|X_{t}|<m.\varepsilon\right)dt\leq\int_{0}^{\infty}\sum_{k}\mathbb{P}_{0,\rho}\left(X_{t}\in k.\varepsilon+[0,\varepsilon)^{d}\right)dt,$ (13)
-where the sum over $k$ is on the interval $\{-m,-m+1,\ldots,m-2,m-1\}^{d}$.
-Let
-$T_{k}=\inf\left\{z\geq 0:X_{z}\in k.\varepsilon+\left[0,\varepsilon\right)^{d}\right\}\;\text{ and }\;\zeta_{\rho}^{k,\varepsilon}=\int_{0}^{\infty}\mathbb{P}_{0,\rho}\left(X_{t}\in k.\varepsilon+[0,\varepsilon)^{d}\right)dt.$
-. By the Law of Total Probability, and Fubini’s Theorem, we can rewrite $\zeta_{\rho}^{k,\varepsilon}$ as
-$\begin{split}\zeta_{\rho}^{k,\varepsilon}=&\int_{0}^{\infty}\int_{0}^{t}\mathbb{P}_{0,\rho}\left(X_{t}\in k.\varepsilon+[0,\varepsilon)^{d},\;T_{k}=y\right)dy\;dt\\ =&\int_{0}^{\infty}\int_{0}^{\infty}\mathbb{P}_{0,\rho}(|X_{t}-X_{y}|<\varepsilon,T_{k}=y)dy\;dt\\ =&\int_{0}^{\infty}\int_{y}^{\infty}\mathbb{P}_{0,\rho}(|X_{t}-X_{y}|<\varepsilon,T_{k}=y)dt\;dz\\ =&\int_{0}^{\infty}\mathbb{P}_{0,\rho}(T_{k}=z)\int_{y}^{\infty}\mathbb{P}_{0,\rho}(|X_{t}-X_{y}|<\varepsilon)dt\;dy,\end{split}$ (14)
-[…]
-Denote $\mathbb{P}_{0,\rho}^{k,y}=\mathbb{P}_{0,\rho}\left(T_{k}=y\right).$ An application of Fubini’s Theorem gives
-$\begin{split}\zeta_{\varepsilon,k}^{\rho}=&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}\int_{y}^{\infty}\mathbb{P}_{0,\rho}\left(|X_{t}-X_{y}|<\varepsilon\right)dtdy\\ =&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}\int_{y}^{\infty}\sum_{\sigma\in\mathcal{D},j\in\mathbb{Z}}\mathbb{P}_{0,\rho}\left(|X_{t}-X_{y}|<\varepsilon,(X_{y},\rho_{y})=(j,\sigma)\right)dtdy\\ =&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}\sum_{\sigma\in\mathcal{D},j\in\mathbb{Z}}\int_{y}^{\infty}\mathbb{P}_{0,\rho}\left(|X_{t}-X_{y}|<\varepsilon|(X_{y},\rho_{y})=(j,\sigma)\right)\mathbb{P}_{0,\rho}\left((X_{y},\rho_{y})=(j,\sigma)\right)dtdy\\ \zeta_{\varepsilon,k}^{\rho}=&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}\sum_{\sigma\in\mathcal{D},j\in\mathbb{Z}}\int_{y}^{\infty}\mathbb{P}_{j,\sigma}\left(|X_{t-y}-X_{0}|<\varepsilon\right)dt\mathbb{P}_{0,\rho}\left((X_{y},\rho_{y})=(j,\sigma)\right)dy\\ =&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}\sum_{\sigma\in\mathcal{D},j\in\mathbb{Z}}\int_{0}^{\infty}\mathbb{P}_{j,\sigma}\left(|X_{s}-j|<\varepsilon\right)ds\mathbb{P}_{0,\rho}\left((X_{y},\rho_{y})=(j,\sigma)\right)dy\\ =&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}\sum_{\sigma\in\mathcal{D}}\int_{0}^{\infty}\mathbb{P}_{0,\sigma}\left(|X_{s}|<\varepsilon\right)ds\sum_{j\in\mathbb{Z}}\mathbb{P}_{0,\rho}\left((X_{y},\rho_{y})=(j,\sigma)\right)dy\\ \leq&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}\max_{\tau\in\mathcal{D}}\int_{0}^{\infty}\mathbb{P}_{0,\tau}\left(|X_{s}|<\varepsilon\right)ds\sum_{\sigma\in\mathcal{D},j\in\mathbb{Z}}\mathbb{P}_{0,\rho}\left((X_{y},\rho_{y})=(j,\sigma)\right)dy\\ =&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}dy\max_{\tau\in\mathcal{D}}\int_{0}^{\infty}\mathbb{P}_{0,\tau}\left(|X_{s}|<\varepsilon\right)ds\\ =&\max_{\tau\in\mathcal{D}}\int_{0}^{\infty}\mathbb{P}_{0,\tau}\left(|X_{s}|<\varepsilon\right)ds.\end{split}$
-Therefore, Equation (13) and the inequality above give
-$\int_{0}^{\infty}\mathbb{P}_{0,\rho}(|X_{t}|<m\varepsilon)dt\leq\sum_{k}\zeta_{\rho}^{k,\varepsilon}\leq\sum_{k}\max_{\tau\in\mathcal{D}}\int_{0}^{\infty}\mathbb{P}_{0,\tau}(|X_{s}|<\varepsilon)ds=(2m)^{d}\cdot\max_{\tau\in\mathcal{D}}\int_{0}^{\infty}\mathbb{P}_{0,\tau}(|X_{s}|<\varepsilon)ds.$
-Theorem 6.10.
-If $X_{t}$ denotes a CTOQW on $\mathbb{Z}^{2}$ and $X_{t}/t^{1/2}\Rightarrow$ a Gaussian measure, then the CTOQW is recurrent for some density.
-Proof.
-Let us denote $u(t,m)=\mathbb{P}(|X_{t}|<m).$ By Lemma 6.9,
-$\max_{\tau\in D}\int_{0}^{\infty}\mathbb{P}(|X_{t}|<1)dt=\max_{\tau\in D}\int_{0}^{\infty}u(t,1)dt\geq(4m^{2})^{-1}\int_{0}^{\infty}u(t,m)dt.$
-Assume that $m/\sqrt{t}\rightarrow c,$ then
-$u(t,m)\rightarrow\int_{[-c,c]^{2}}f(x)dx,$ (15)
-where $f(x)$ represents the density of the limiting normal distribution. We denote the right-hand side of Equation (15) by $g(c)$ and let $t=\theta m^{2}$ to obtain
-$u(\theta m^{2},m)\rightarrow g(\theta^{-1/2}).$
-Now we use the chenge of variables
-$m^{-2}\int_{0}^{\infty}u(t,m)dt=\int_{0}^{\infty}u(\theta m^{2},m)d\theta,$
-let $m\rightarrow\infty,$ and use Fatou’s lemma to get
-$\liminf_{m\rightarrow\infty}(4m^{2})^{-1}\int_{0}^{\infty}u(t,m)dt\geq 4^{-1}\int_{0}^{\infty}g(\theta^{-1/2})d\theta.$ (16)
-Since the normal density is positive and continuous at $0,$
-$g(c)=\int_{[-c,c]^{2}}n(x)dx\sim n(0)(2c)^{2}$
-as $c\rightarrow 0.$ So $g(\theta^{-1/2})\sim 4n(0)/\theta$ as $\theta\rightarrow\infty,$ the integral in Equation (16) diverges.
-As a result, $\max_{\tau\in D}\int_{0}^{\infty}\mathbb{P}(|X_{t}|<1)dt=\infty,$ leading us to conclude that the walk is $\tau$-recurrent.
-$\square$
-Note that the maximum in Lemma 6.9 can be assumed to be taken over the jump times $(T_{n})_{n}.$ Therefore, by Proposition 6.7 and Theorem 6.10 we obtain the following result.
-Theorem 6.11.
-If $X_{t}$ denotes a CTOQW induced by a coin $(A,H),$ there is at most one minimal enclosure for $\Phi$ and $X_{t}/t^{1/2}\Rightarrow$ a Gaussian measure, then the CTOQW is recurrent.
-We already have a sufficient condition to recurrence given by the Theorem 6.11. For transience, we also have a sufficient condition:
-Theorem 6.12.
-Let us consider a CTOQW induced by a coin $(A,H),$ having at most one minimal enclosure, and let $\rho_{\infty}$ be the unique stationary state of the auxiliary map.
-If $m\neq 0,$ then the CTOQW is transient.
-Proof.
-Let us start the CTOQW at some vertex $\lvert i\rangle\in\mathbb{Z}^{d}$ with initial density operator $\rho.$ By the Law of Large Numbers, Proposition 6.5,
-$\frac{X_{t}}{t}\rightarrow m\neq 0^{d}.$
-This means that at least one coordinate of $X_{t}$ is unbounded, from where we conclude that
-$\mathbb{P}\left(X_{t}=X_{0}\;\text{i.o.}\right)=0.$
-Hence, the CTOQW is transient.
-$\square$
-Finally, we can present a recurrence criteria for CTOQWs on $\mathbb{Z}^{2}$ induced by a coin $(A,H).$
-Theorem 6.13 (Recurrence criteria for CTOQWs on $\mathbb{Z}^{2}$).
-Let us consider a CTOQW induced by a coin $(A,H),$ having at most one minimal enclosure, and let $\rho_{\infty}$ be the unique stationary state of the auxiliary map. Then
-  • $m=0\;\Rightarrow\;$ the CTOQW is recurrent;
-  • $m\neq 0\;\Rightarrow\;$ the CTOQW is transient.
-Proof.
-For $m\neq 0,$ the result is given in Theorem 6.12. It remains to proof that $m=0$ implies that the CTOQW is recurrent. Firstly, if $m=0,$ then by Theorem 6.4
-$\frac{X_{t}}{\sqrt{t}}=\frac{X_{t}-t.m}{\sqrt{t}}\Rightarrow\text{ a non-degenerate normal distribution,}$
-thus we use Theorem 6.10 to obtain a density operator $\tau$ such that the CTOQW is $\tau$-recurrent. Therefore, $(A,H)$ is recurrent by Theorem 6.11.
-$\square$
-6.5 Recurrence criteria for HOQWs
-6.5.1 Criterion for auxiliary map having at most one minimal enclosure
-To present a complete criterion for the recurrence of finite-dimensional coins $(A)$, we will start this section with an easily verifiable criterion based on the quantum jump chains associated with CTOQWs, as presented in [20].
-The complete recurrence criterion will be developed from a decomposition of the Hilbert space $\mathfrak{h}$ into subspaces that support invariant states.
-The central idea is identifying and characterizing these subspaces, thus we will be able to determine the recurrence of quantum walks by analyzing the evolution within each invariant component.
-Furthermore, these invariant states will play an indispensable role in formulating the criterion.
-Let $\Pi$ denote a discrete HOQW given by
-$\Pi(\rho)=\sum_{i,j\in V}M_{j}^{i}\rho(j)M_{j}^{i*}\otimes\lvert i\rangle\langle i\rvert$.
-This HOQW is the open quantum jump chain of the CTOQW with Lindblad generator
-$\mathbb{L}=\Psi-I,\;\;\Psi(\rho)=\sum_{i,j\in V}B_{j}^{i}\rho(j)B_{j}^{i*},\;B_{j}^{i}=M_{j}^{i}\otimes\lvert i\rangle\langle j\rvert$, (17)
-where $I$ is the identity operator acting on $\mathcal{H}$ [20].
-Theorem 6.14.
-[20, Theorem 5.6] A vertex $i\in V$ is $\rho$-recurrent for the HOQW $\Pi$ if and only if it is $\rho$-recurrent for the CTOQW generated by $\mathbb{L}$ given by Equation (17).
-We consider an HOQW with a set of vertices $V=\mathbb{Z}^{2}.$ Analogously to the continuous-time version, if there exists $D_{1},D_{2},D_{3},D_{4}\in\mathcal{B}(\mathbb{C}^{n})$ such that
-$D_{1}=M_{(i,j)}^{(i+1,j)},\;D_{2}=M_{(i,j)}^{(i,j+1)},\;D_{3}=M_{(i,j)}^{(i-1,j+1)},\;D_{4}=M_{(i,j)}^{(i,j-1)},$
-and $M_{x}^{y}=0_{n}$ for the remaining $x,y\in\mathbb{Z}^{2},$ then the HOQW $\Pi$ is said to be induced by a coin $(D)$ of dimension $n.$
-The HOQW induced by a coin $(D)$ is the homogeneous, nearest-neighbor HOQW on the grid, with the right, up, left and down transitions given by $D_{1},D_{2},D_{3},D_{4},$ respectively.
-Therefore, the quantum trajectory of those walks, starting from a state $\tau$ of the form $\tau=\sum_{i\in\mathbb{Z}}\rho(i)\otimes|i\rangle\langle i|,$ is any path generated by the Markov chain $(X_{n},\tau_{n})_{n\geq 0}$, where $X_{n}$ denotes the position of the particle at time $n$ and $\tau_{n}$ its internal degree.
-The transition probabilities are given by
-$\begin{split}\mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left((i+1,j),\frac{D_{1}\sigma D_{1}^{*}}{\mathrm{Tr}(D_{1}\sigma D_{1}^{*})}\right)\;\Bigg|\;(X_{n},\rho_{n})=((i,j),\sigma)\,\right)&=\mathrm{Tr}(D_{1}\sigma D_{1}^{*}),\\ \mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left((i,j+1),\frac{D_{2}\sigma D_{2}^{*}}{\mathrm{Tr}(D_{2}\sigma D_{2}^{*})}\right)\;\Bigg|\;(X_{n},\rho_{n})=((i,j),\sigma)\,\right)&=\mathrm{Tr}(D_{2}\sigma D_{2}^{*}),\\ \mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left((i-1,j),\frac{D_{3}\sigma D_{3}^{*}}{\mathrm{Tr}(D_{3}\sigma D_{3}^{*})}\right)\;\Bigg|\;(X_{n},\rho_{n})=((i,j),\sigma)\,\right)&=\mathrm{Tr}(D_{3}\sigma D_{3}^{*}),\\ \mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left((i,j-1),\frac{D_{4}\sigma D_{4}^{*}}{\mathrm{Tr}(D_{4}\sigma D_{4}^{*})}\right)\;\Bigg|\;(X_{n},\rho_{n})=((i,j),\sigma)\,\right)&=\mathrm{Tr}(D_{4}\sigma D_{4}^{*}),\end{split}$
-for every $(i,j)\in\mathbb{Z}^{2}$, $\sigma\in\mathcal{D}$, and initial law
-$\mathbb{P}\left(\,(X_{0},\rho_{0})=\left((i,j),\frac{\rho(i,j)}{\mathrm{Tr}\,\rho(i,j)}\right)\;\right)=\mathrm{Tr}(\rho(i,j)),$
-and all other transition probabilities are null. The graph of this kind of HOQW is represented in Figure 3.
-$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}\vdots\vdots\vdots\vdots\ldots\vdots\vdots\vdots\vdots$
-Figure 3: Homogeneous Open Quantum walk on $\mathbb{Z}^{2}$.
-Now we can give a recurrence criteria for HOQWs on $\mathbb{Z}^{2}.$ To this, note that if $\rho$ is an invariant state for the auxiliary map $\Phi$ of $(D),$ defined as
-$\Phi(.)=\sum_{j=1}^{4}D_{j}(.)D_{j}^{*},$
-then it is a stationary state for its open quantum jump chain $\mathbb{L}.$ Consequently, if $\Phi$ has at most one minimal enclosure, then $\Phi$ has a unique invariant state $\rho_{\infty}$ and we can let
-$m=\begin{bmatrix}m_{1}\\ m_{2}\end{bmatrix},\quad m_{1}=\mathrm{Tr}\left(D_{1}\rho_{\infty}D_{1}^{*}\right)-\mathrm{Tr}\left(D_{3}\rho_{\infty}D_{3}^{*}\right),\quad m_{2}=\mathrm{Tr}\left(D_{2}\rho_{\infty}D_{2}^{*}\right)-\mathrm{Tr}\left(D_{4}\rho_{\infty}D_{4}^{*}\right)$
-to obtain, as a consequence of Theorems 6.13 and 6.14, the following criterion.
-Theorem 6.15 (Recurrence criterion for HOQWs on $\mathbb{Z}^{2}$).
-Let us consider an HOQW induced by a coin $(D),$ where the map
-$\Phi(.)=\sum_{j=1}^{4}D_{j}(.)D_{j}^{*}$
-has at most one minimal enclosure, and let $\rho_{\infty}$ be the unique invariant state of $\Phi.$ Then
-  • $m=0\;\Rightarrow\;$ the HOQW is recurrent;
-  • $m\neq 0\;\Rightarrow\;$ the HOQW is transient.
-6.5.2 A generalized recurrence criteria for HOQWs on $\mathbb{Z}^{2}$
-We have already presented several recurrence criteria for discrete-time and continuous-time Open Quantum Walks on the set of vertices $\mathbb{Z}^{2}$, constrained by a minimal enclosure.
-These criteria are interesting tools for understanding the long-term behavior of open quantum walks.
-Building on these consequences, we aim to derive a stronger recurrence criterion for discrete-time HOQWs, allowing for a more comprehensive analysis of their statistics.
-At this point we will allow the auxiliary map of a coin $(D)$ to have more than one invariant state, thus we consider the decomposition of $\mathfrak{h}$ given on Section 4.
-Therefore, given a minimal enclosure $\mathcal{Y}_{\alpha},$ let $\tau_{\alpha}$ be the unique invariant state with support on $\mathcal{Y}_{\alpha}.$ Introducing the parameter
-$\tilde{m}_{\alpha}=\begin{bmatrix}m_{1,\alpha}\\ m_{2,\alpha}\end{bmatrix},\quad m_{1,\alpha}=\mathrm{Tr}\left(D_{1}\tau_{\alpha}D_{1}^{*}\right)-\mathrm{Tr}\left(D_{3}\tau_{\alpha}D_{3}^{*}\right),\quad m_{2,\alpha}=\mathrm{Tr}\left(D_{2}\tau_{\alpha}D_{2}^{*}\right)-\mathrm{Tr}\left(D_{4}\tau_{\alpha}D_{4}^{*}\right),$
-we have by [10, Lema 4.2] that $\tilde{m}_{\alpha}$ is independent of the particular minimal enclosure $\mathcal{Y}_{\alpha}$ in $\mathfrak{X}_{\alpha}.$
-Therefore, we can combine Theorem 6.15 and results concerning CTOQWs above to state the next property, which is a continuous-time version on the grid of Proposition 4.1.
-Proposition 6.16.
-Consider a coin $(D).$ We have a decomposition
-$\mathfrak{h}=\left(\bigoplus_{\alpha\in A}\mathcal{Y}_{\alpha}\right)\oplus\mathcal{X},$ (18)
-where $A$ is a finite set, $\mathcal{Y}_{\alpha}$ is a subspace of $\mathfrak{h}$ being the support of exactly one invariant state $\tau_{\alpha}$ for $\Phi,$ $\mathcal{X}$ contains no enclosure, and
-  • $\tilde{m}_{\alpha}=0\Rightarrow$ the walk is $\tau_{\alpha}$-recurrent for all $\tau\in\mathcal{S}(\mathcal{Y}_{\alpha})$;
-  • $\tilde{m}_{\alpha}\neq 0\Rightarrow$ the walk is $\tau_{\alpha}$-transient for all $\tau\in\mathcal{S}(\mathcal{Y}_{\alpha}).$
-Proof.
-Fixed a subspace $\mathcal{Y}_{\alpha}$ which is the support of exactly one invariant state $\tau_{\alpha}\in\mathcal{D},$ we assume that $\tilde{m}_{\alpha}=0.$ Then $(D)$ is $\rho$-recurrent for every $\rho$ by Theorem 6.11.
-If $\tilde{m}_{\alpha}\neq 0,$ then $(D)$ is transient by Theorem 6.12.
-$\square$
-With the decomposition on Equation (18) established, we can say that $\mathcal{Y}_{\alpha}$ is either recurrent or transient.
-It remains to understand what happens when a quantum state has support outside of $\oplus_{\alpha}\mathcal{Y}_{\alpha}.$
-Analogously to the one dimensional case, our recurrence criterion is:
-Theorem 6.17 (A Generalized Criterion for coins $(D)$).
-Consider a coin $(D)$ of dimension $d<\infty.$ Consider the decomposition of $\mathfrak{h}$ as given in Equation (18), and $\tau_{1},\ldots,\tau_{w}$ the respectively invariant states of $\Phi$ in $\mathcal{Y}_{\alpha},\;\alpha=1,\ldots,w.$
-$(1)\quad\tilde{m}_{\alpha}\neq 0\;\forall j\in\{1,\ldots w\}\Rightarrow(D)\mbox{ is transient;}$
-[…]
-[…]
-$\begin{cases}\rho\text{-transient for }\rho\in\mathcal{D}_{1}:=\mathcal{D}(\oplus_{k\in J^{C}}\mathcal{Y}_{k})\\ \rho\text{-recurrent for }\rho\in\mathcal{D}_{1}^{C}.\end{cases}$
-[…]
-$\begin{cases}\rho\text{-transient for }\rho\in\mathcal{D}_{2}:=\mathcal{D}(\oplus_{k\in J^{C}}\mathcal{Y}_{k}\oplus\mathcal{X})\\ \rho\text{-recurrent for }\mathcal{D}_{2}^{C}\end{cases}$
-$(3)\quad\tilde{m}_{\alpha}=0\;$ for every $\alpha\Rightarrow$ the walk is recurrent.
-7 Examples
+6.2 2D
 Example 6.7.
@@
 $A_{1}=\begin{bmatrix}3&-1\\ 0&0\end{bmatrix},\quad A_{2}=\begin{bmatrix}1&-2\\ 2i&0\end{bmatrix},\quad A_{3}=\begin{bmatrix}1&1\\ -2&2\end{bmatrix},\quad A_{4}=\begin{bmatrix}-2i&i\\ 0&2\end{bmatrix},\quad H=\begin{bmatrix}-1&h\\ \overline{h}&2\end{bmatrix},\;h\in\mathbb{C}.$
-In this case,
-$m=\frac{2\;\text{Re}(h)-10\;\text{Im}(h)+19}{\alpha}\;\begin{bmatrix}4\\ 1\end{bmatrix}$,
-where $\alpha$ depends on $h,$ but it is never null. Therefore, by Theorem 6.13, $(A,H)$ is recurrent if and only if
+Operator $\mathbb{L}$ has a unique stationary state $\rho_{\infty},$ where
+$\rho_{\infty}=\begin{bmatrix}\rho_{1}&\rho_{2}\\ \rho_{2}&1-\rho_{1}\end{bmatrix},\quad\rho_{1}=\dfrac{19(\text{Re}\,h)^{2}-8(\text{Re}\,h)(\text{Im}\,h)+15(\text{Im}\,h)^{2}-54\,\text{Re}\,h+100\,\text{Im}\,h+830}{\alpha},$
+$\rho_{2}=\dfrac{20i(\text{Re}\,h)(\text{Im}\,h)-39i(\text{Re}\,h)-37i(\text{Im}\,h)+20(\text{Re}\,h)^{2}-45(\text{Re}\,h)+15(\text{Im}\,h)+1165+309i}{2\alpha},$
+$\alpha=38(\text{Re}\,h)^{2}-16(\text{Re}\,h)(\text{Im}\,h)+30(\text{Im}\,h)^{2}-78\,\text{Re}\,h+50\,\text{Im}\,h+1945.$ In this case,
+$m=\frac{2\;\text{Re}\;h-10\;\text{Im}\;h+19}{\alpha}\;\begin{bmatrix}4\\ 1\end{bmatrix}$,
+Therefore, by Theorem 4.13, $(A,H)$ is recurrent if and only if
 $\text{Re}(h)=5\;\text{Im}(h)-\frac{19}{2}.$
@@
 $H_{1}=\begin{bmatrix}0&0&0\\ 0&0&0\\ 0&0&0\end{bmatrix},\qquad H_{2}=\begin{bmatrix}0&1&0\\ 1&0&0\\ 0&0&0\end{bmatrix}$.
-Its straightforward that $(A,H_{1})$ is transient, since we cain join the results of Example 7.2 and Theorem 6.14. However, it is very interesting to see that $(A,H_{2})$ is recurrent. Indeed,
+It is straightforward that $(A,H_{1})$ is transient, since we can join the results of Example 6.8 and Theorem 4.14. However, it is very interesting to see that $(A,H_{2})$ is recurrent. Indeed,
 $\rho_{\infty}=\frac{1}{2}\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&0\end{bmatrix}$
@@
 $\Diamond$
+In Example 6.3, the operator $H_{2}$ plays a role analogous to that of $B^{\prime}$ in Example 6.9.
+Given the transition operators $A_{k}$ (with $k=1,2,3,4$), the choice of $H_{2}$ alters the stationary state $\rho_{\infty}$ of the auxiliary map, yielding $m=0$ and consequently recurrence for the CTOQW.
+This illustrates how, in both discrete and continuous-time settings, modifying the “non-jump” component (whether $B$ in the lazy HOQW case or $H$ in the CTOQW case) can tune the drift parameters and induce recurrence, even when the transition operators alone would lead to transience.
 […]
 Acknowledgements This work was supported by Universidade Federal do Rio Grande do Sul (UFRGS).
+The author also thanks the reviewers for their comments and suggestions, which significantly improved the quality of the manuscript.
 […]